### v1

# RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model

Abstract
Human Action Recognition (HAR) encompasses the task of monitoring human activities across various domains, including but not limited to medical, educational, entertainment, visual surveillance, video retrieval, and the identification of anomalous activities. Over the past decade, the field of HAR has witnessed substantial progress by leveraging Convolutional Neural Networks (CNNs) to effectively extract and comprehend intricate information, thereby enhancing the overall performance of HAR systems. Recently, the domain of computer vision has witnessed the emergence of Vision Transformers (ViTs) as a potent solution. The efficacy of transformer architecture has been validated beyond the confines of image analysis, extending their applicability to diverse video-related tasks. Notably, within this landscape, the research community has shown keen interest in HAR, acknowledging its manifold utility and widespread adoption across various domains. This article aims to present an encompassing survey that focuses on CNNs and the evolution of Recurrent Neural Networks (RNNs) to ViTs given their importance in the domain of HAR. By conducting a thorough examination of existing literature and exploring emerging trends, this study undertakes a critical analysis and synthesis of the accumulated knowledge in this field. Additionally, it investigates the ongoing efforts to develop hybrid approaches. Following this direction, this article presents a novel hybrid model that seeks to integrate the inherent strengths of CNNs and ViTs.
keywords:
Human action recognition; convolutional neural networks; recurrent neural networks; vision transformers; deep learning; video classification
1 Introduction
Human Action Recognition (HAR) focuses on the classification of the specific actions exhibited within a given video. On the other hand, action detection and segmentation focus on the precise localization or extraction of individual instances of actions from video content [1]. The capacity of deep learning models to effectively capture the spatial and temporal complexities inherent in video representations plays a vital role in the recognition and understanding of actions.
Over the preceding decade, a considerable amount of research has been dedicated to the thorough investigation of action recognition, resulting in an extensive collection of review articles and survey papers addressing the topic [2, 3, 4]. However, it is worth noting that a predominant focus of these scholarly works has been placed on the examination and evaluation of Convolutional Neural Networks (CNNs) and traditional machine learning models within the realm of action recognition.
The advent of Transformer[5] architecture has sparked a paradigm shift in deep learning, leading researchers to swiftly adopt and adapt these architectures to enhance the accuracy and efficiency of various computer vision tasks including action recognition. Consequently, there is a compelling need for a recent review that comprehensively examines the state-of-the-art research including Vision Transformers (ViTs) [6], a special Transformer for computer vision tasks, and Hybrid models in addition to CNNs. Such a review would serve as a crucial guiding resource to shape the future research directions with Transformer and CNN-Transformer hybrid architectures beside CNNs which previously were seen as unique and influential models for action recognition.
By employing a multi-head self-attention layer, the Transformer model computes sequence representations by effectively aligning words within the sequence with other words in the same sequence [1]. This approach outperforms traditional convolutional and recursive operations in terms of representation quality while utilizing fewer computational resources. As a consequence, the Transformer architecture diverges from conventional convolutional and recursive methods, favoring a more focused utilization of multiple processing nodes. The incorporation of multi-head attention allows the Transformer model to collectively learn a range of representations from diverse perspectives through the collaboration of multiple attention layers. Inspired by Transformers, several Natural Language Processing (NLP) tasks have achieved remarkable performance, reaching human-level capabilities, as exemplified by models such as GPT [7] and BERT [8].
The remarkable achievements of Transformers in handling sequential data, particularly in the domain of NLP, have prompted the exploration and advancement of ViTs [6]. These novel models have demonstrated comparable or even superior performance compared to CNNs in the context of image recognition tasks, especially when operating on vast datasets such as Imagenet [9, 10, 11]. This observation signifies a noteworthy shift in the field, wherein ViTs possess the potential to supplant the established dominance of CNNs in computer vision, mirroring the displacement witnessed in the case of Recurrent Neural Networks (RNNs) [1]. The achievements of Transformer models have engendered considerable scholarly interest within the computer vision research community, prompting rigorous exploration of their efficacy in pure computer vision tasks.
The natural progression in the advancement of ViTs has led to the logical exploration of video recognition tasks. Unlike image recognition, video recognition focuses on the complex challenge of identifying and understanding events within video sequences, including the recognition of human actions. The main contributions of this paper is as follows.
  • We present a thorough review of the CNNs, RNNs, and ViTs. This review examines the evolution from traditional methods to the latest advancements in neural network architectures.
  • We present an extensive examination of existing literature related to action recognition.
  • We propose a novel hybrid model integrating the strengths of CNNs and ViTs. In addition, we provide a detailed performance comparison the proposed hybrid model against existing models. The analysis highlights the model’s efficacy in handling complex HAR tasks with improved accuracy and efficiency.
  • We also discuss emerging trends and the future direction of HAR technologies, emphasizing the importance of hybrid models in enhancing the interpretability and robustness of action recognition systems.
These contributions enrich the understanding of the current state and future prospects of HAR, proposing innovative approaches and highlighting the importance of integrating different neural network architectures to advance the field.
The paper is structured as follows: Section 2 delves into the background, covering foundational concepts and technologies crucial to HAR, including CNNs, RNNs, and ViTs. Section 3 reviews related work, highlighting the chronological evolution of HAR deep learning technologies. The novel hybrid model combining CNNs and ViTs is introduced in Section 4, followed by Section 5, which details the experimental setup and evaluation of the model. Section 6 discusses the challenges and their implications for future directions. Finally, Section 7 concludes the paper.
2 Background
2.1 From Vanilla RNN to Attention-based Transformers
The fundamental aim of the Transformers lies in effectively addressing time series data tasks, placing a distinct emphasis on NLP as one of its primary objectives. To address these tasks, RNNs have emerged as a prominent deep learning technique. However, the development of the Transformers did not occur in isolation; rather, it was the result of a series of iterative optimisations that were progressively integrated into the existing RNN structures. This section aims to elucidate the underlying RNN techniques and the subsequent advancements that were incorporated into them, ultimately leading to the final formulation of the Transformer architecture.
Notation. Here, we establish common notations for RNN architectures including Vanilla RNNs, Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) to streamline discussions in subsequent sections. In these architectures, each iteration involves a cell that sequentially processes an input embedding $\boldsymbol{x}_{t}\in\mathbb{R}^{n_{x}}$ and retains information from the previous sequence through the hidden state $\boldsymbol{h}_{t-1}\in\mathbb{R}^{n_{h}}$ using weight matrices $\boldsymbol{W}\in\mathbb{R}^{n_{h}\times n_{h}}$ and $\boldsymbol{U}\in\mathbb{R}^{n_{h}\times n_{x}}$. The $\boldsymbol{W}$-like matrices encompasses all weights related to hidden-to-hidden connections, while $\boldsymbol{U}$-like matrices encompasses all weight matrices related to input-to-hidden connections. Additionally, bias terms are represented by $\boldsymbol{b}$-like vectors. Each cell produces a new hidden state $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ as its output.
Vanilla RNNs [12, 13] lack the presence of a cell state, relying solely on the hidden states as the primary means of memory retention within the RNN framework. The hidden state $\boldsymbol{h}_{t}$ is subsequently updated and propagated to the subsequent cell, or alternatively, depending on the specific task at hand, it can be employed to generate a prediction. Figure 1(a) illustrates the internal mechanisms of an RNN and a mathematical description of it given as
$\boldsymbol{h}_{t}=\tanh(\boldsymbol{W}\boldsymbol{h}_{t-1}+\boldsymbol{U}\boldsymbol{x}_{t}+\boldsymbol{b}),$ (1)
where tanh is the activation function.
Vanilla RNNs effectively incorporate short-term dependencies of temporal order and past inputs in a meaningful manner. However, they are characterised by certain limitations. Firstly, due to their intrinsic sequential nature, RNNs pose challenges in parallelizing computations [14]. Consequently, this limitation can impose restrictions on the overall speed and scalability of the network. Secondly, when processing lengthy sequences, the issue of exploding or vanishing gradients may arise, thereby impeding the stable training of the network [15].
[…]
(a) Vanilla
[…]
(b) LSTM
[…]
Figure 1: GRU
Figure 2: Various types of RNN cells.
LSTM Hochreiter and Schmidhuber (1997) introduced the Long Short-Term Memory (LSTM) cell as a solution to address the issue of long-term dependencies and to mitigate the challenge of interdependencies among successive steps [16]. LSTM architecture incorporates a distinct component known as the cell state $\boldsymbol{c}_{t}\in\mathbb{R}^{n_{h}}$, illustrated in Figure 1(b). Analogous to a freeway, this cell state facilitates the smooth flow of information, ensuring that it can readily traverse without undergoing significant alterations.
In 2000, Gers, Schmidhuber, and Cummins [17] made modifications to the initial LSTM architecture by incorporating a forget gate within the cell structure. The mathematical expressions describing this modified LSTM cell are derived from its inner connections. Hence, the LSTM cell can be formally represented based on the depicted interconnections as follows:
  • Forget Gate decides what information should be thrown away or kept from the cell state with the equation
    $\boldsymbol{f}_{t}=\sigma(\boldsymbol{W}_{f}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{f}\boldsymbol{x}_{t}+\boldsymbol{b}_{f}),$ (2)
    where $\boldsymbol{f}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the forget gate and $\sigma$ is sigmoid activation function.
  • Input Gate determines which new information is added to the cell state with two activation functions defined as
    $\boldsymbol{i}_{t}=\sigma(\boldsymbol{W}_{i}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{i}\boldsymbol{x}_{t}+\boldsymbol{b}_{i}),$ (3)
    where $\boldsymbol{i}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the sigmoid activation function and
    $\boldsymbol{\tilde{c}}_{t}=\tanh(\boldsymbol{W}_{\tilde{c}}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{\tilde{c}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{c}})$ (4)
    where $\boldsymbol{\tilde{c}}_{t}\in\mathbb{R}^{n_{h}}$ is known as candidate value. After obtaining $\boldsymbol{i}_{t}$ and $\boldsymbol{\tilde{c}}_{t}$, we can update the cell state with
    $\boldsymbol{c}_{t}=\boldsymbol{f}_{t}\odot\boldsymbol{c}_{t-1}+\boldsymbol{i}_{t}\odot\boldsymbol{\tilde{c}}_{t}$ (5)
    where $\boldsymbol{c}_{t-1}\in\mathbb{R}^{n_{h}}$ is the previous cell state.
  • Output Gate determines the next hidden state based on the cell state and output gate’s activity
    $\boldsymbol{o_{t}}=\sigma(\boldsymbol{W}_{o}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{o}\boldsymbol{x}_{t}+\boldsymbol{b}_{o})$ (6)
    where $\boldsymbol{o}_{t}\in\mathbb{R}^{n_{h}}$ is the output of output gate. Finally the updated hidden state,
    $\boldsymbol{h}_{t}=\tanh(\boldsymbol{c}_{t})\odot\boldsymbol{o}_{t}$ (7)
    is fed to the next iteration.
To enable selective information retention, LSTM employs three distinct gates. The first gate, known as the forget gate, examines the previous hidden state $\boldsymbol{h}_{t-1}$ and the current input $\boldsymbol{x}_{t}$. It generates a vector $\boldsymbol{f}_{t}$ containing values between 00 and $1$, determining the portion of information to discard from the previous cell state $\boldsymbol{c}_{t-1}$. The second gate, referred to as the input gate, follows a similar process to the forget gate. However, instead of discarding information, it utilizes the output $\boldsymbol{i}_{t}$ to determine the new information to be stored in the cell state based on a candidate cell state $\boldsymbol{\tilde{c}}_{t}$. Lastly, the output gate employs the output $\boldsymbol{o}_{t}$ to filter the updated cell state $\boldsymbol{c}_{t}$, thereby transforming it into the new hidden state $\boldsymbol{h}_{t}$. The LSTM cell exhibits superior performance in retaining both long-term and short-term memory compared to the Vanilla RNN cell. However, this advantage comes at the expense of increased complexity.
GRU The LSTM cell surpasses the learning capability of the conventional recurrent cell, yet the additional number of parameters escalates the computational load. Consequently, to address this concern, Chung et al. [18] introduced the gated recurrent unit (GRU), see Figure 2. GRU demonstrates comparable performance to LSTM while offering a more computationally efficient design with fewer weights. This is achieved by merging the cell state and the hidden state into “reset state” resulting in a simplified architecture. Furthermore, GRU combines the forget and input gates into an “update gate”, contributing to a more streamlined computational process. For further elaboration, GRU cell incorporates two essential gates. The first gate is the reset gate, which examines the previous hidden state $\boldsymbol{h}_{t-1}$ and the current input $\boldsymbol{x}_{t}$. It generates a vector $\boldsymbol{r}_{t}$ containing values between 00 and $1$, determining the extent to which past information in $\boldsymbol{h}_{t-1}$ should be disregarded. The second gate is the update gate, which governs the selection of information to either retain or discard when updating the new hidden state $\boldsymbol{h}_{t}$, based on the value of $\boldsymbol{r}_{t}$.
Based on the depicted information in Figure 2, the mathematical expressions governing the behavior of the GRU cell can be expressed as follows.
  • Update Gate decides how much of the past information needs to be passed along with
    $\boldsymbol{z}_{t}=\sigma(\boldsymbol{W}_{z}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{z}\boldsymbol{x}_{t}+\boldsymbol{b}_{z})$ (8)
    where $\boldsymbol{z}_{t}\in\mathbb{R}^{n_{h}}$ is the output of update gate. A candidate activation for the subsequent step
    $\boldsymbol{\tilde{h}}_{t}=\tanh(\boldsymbol{W}_{\tilde{h}}(\boldsymbol{r}_{t}\odot\boldsymbol{h}_{t-1})+\boldsymbol{U}_{\tilde{h}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{h}})$ (9)
    where $\boldsymbol{\tilde{h}}_{t}\in\mathbb{R}^{n_{h}}$ and
    $\boldsymbol{r}_{t}=\sigma(\boldsymbol{W}_{r}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{r}\boldsymbol{x}_{t}+\boldsymbol{b}_{r})$ (10)
    where $\boldsymbol{r}_{t}\in\mathbb{R}^{n_{h}}$ is the output of reset gate.
  • Final Activation is a blend of the previous hidden state and the candidate activation, weighted by the update gate. This mechanism allows the GRU to effectively retain or replace old information with new information.
    $\boldsymbol{h}_{t}=\boldsymbol{z}_{t}\odot\boldsymbol{\tilde{h}}_{t}+(1-\boldsymbol{z}_{t})\odot\boldsymbol{h}_{t-1}$ (11)
    where $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ is the updated hidden state.
Types of RNNs RNNs were created with an internal memory mechanism that allows them to store and use information from previous outputs. This unique trait enables RNNs to retain important contextual information over time, enabling reasoned decision-making based on past results. There are four types of popular RNN variants that each serve different purposes across a variety of applications, see Figure 3.
The One-to-one is considered the simplest form of RNNs, where a single input corresponds to a single output. It operates with fixed input and output sizes, functioning similarly to a standard neural network.
[…]
Figure 3: Types of RNN structures based on input-output pairs.
One-to-many represents a specific category of RNNs that is characterized by its ability to produce multiple outputs based on a single input provided to the model. This type of RNN is particularly useful in applications like Image Captioning, where a fixed input size results in a series of data outputs.
The Many-to-one RNNs merge a sequence of inputs into a single output through a series of hidden layers that learn relevant features. An illustrative instance of this RNN type is Sentiment Analysis, where the model analyzes a sequence of text inputs and produces a single output indicating the sentiment expressed in the text.
The Many-to-many RNNs are employed to generate a sequence of output data from a sequence of input units. It can be categorized into two subcategories: Equal Size and Unequal Size. In the Equal Size subcategory, the input and output layers have the same size, see Many-to-many architecture in Figure 3. Several research efforts have emerged to tackle the limitation of fixed-size input-output sequences in machine translation tasks, as they fail to adequately represent real-world requirements. The Unequal Size subcategory can handle different sizes of inputs and outputs. A practical application of the the Unequal Size subcategory can be observed in Machine Translation. In this scenario, the model generates a sequence of translated text outputs based on a sequence of input sentences. Unequal Size subcategory employs an encoder-decoder architecture, where the encoder adopts the Many-to-One architecture, and the decoder adopts the One-to-Many architecture. One notable contribution in this area was made by Kalchbrenner and Blunsom [19], who pioneered the approach of mapping the entire input sentence to a vector. This work is related to the study conducted by Cho et al. [20], although the latter was specifically utilized to refine hypotheses generated by a phrase-based system [21]. In this architecture, the encoder component plays a crucial role in transforming the inputs into a singular vector, commonly referred to as the context. This context vector, typically with a length of 256, 512, or 1024, encapsulates all the pertinent information detected by the encoder from the input sentence, which serves as the translation target, see Figure 4(a). Subsequently, this vector is passed on to the decoder, which generates the corresponding output sequence. It’s important to note that both the encoder and decoder components in this architecture are RNNs.
2.1.1 Attention
The evolution of attention mechanisms in neural networks represents a significant advancement in the field of deep learning, particularly in tasks related to NLP and machine translation. Initially introduced by Graves et al. [22], the concept of attention mechanisms was designed to enhance the model’s ability to focus on specific parts of the input sequence when generating an output, mimicking the human ability to concentrate on particular aspects of a task. This foundational work laid the groundwork for subsequent developments in attention mechanisms, providing a mechanism for models to dynamically assign importance to different parts of the input data.
Building on Graves’ initial concept, Bahdanau et al. [23] introduced the additive attention mechanism, which was specifically designed to improve machine translation. This approach computes the attention weights through a feed-forward neural network, allowing the model to consider the entire input sequence and determine the relevance of each part when translating a segment. This additive form of attention significantly improved the performance of sequence-to-sequence models by enabling a more nuanced understanding and alignment between the input and output sequences [21]. Following this, Luong et al. [24] proposed the multiplicative attention mechanism, also known as dot-product attention, which simplifies the computation of attention weights by calculating the dot product between the query and all keys. This method not only streamlined the attention mechanism but also offered improvements in computational efficiency and performance in various NLP tasks, marking a pivotal moment in the evolution of attention mechanisms from their inception to more sophisticated and efficient variants.
[…]
(a) without attention
[…]
(b) with attention
Figure 4: Sequence-To-Sequence RNN with and without attention mechanism.
The central idea of the Attention mechanism is to shift focus from the task of learning a single vector representation for each sentence. Instead, it adopts a strategy of selectively attending to particular input vectors in the input sequence, guided by assigned attention weights. This strategy enables the model to dynamically allocate its attention resources to the most pertinent segments of the sequence, thereby improving its capacity to process and comprehend the information more efficiently [25].
One possible explanation for the improvement is that the attention layer created memories associated with the context pattern rather than memories associated with the input itself, relieving pressure on the RNN model structure’s weights and causing the model memory to be devoted to remembering the input rather than the context pattern [26].
2.1.2 Self-Attention
To this point, attention mechanisms in sequence-transformation models have primarily relied on complex RNNs, featuring an encoder and a decoder, the most successful models in language translation yet. However, in 2017, Vaswani et al. [5] introduced a simple network architecture known as the Transformer, see Figure 5, which exclusively utilized attention mechanism, eliminating the need for RNNs. They introduced a novel attention mechanism called self-attention, which is also known as KQV-attention (Key, Query, and Value). This attention mechanism subsequently gained prominence as a central component within the Transformer architecture. The attention mechanism stands out due to its ability to provide Transformers with an extensive long-term memory. In the Transformer model, it becomes possible to focus on all previously generated tokens.
The embedding layer in a Transformer model is the initial stage where input tokens are transformed into dense vectors, capturing semantic information about each token’s meaning and context within the text. These embeddings serve as the foundation for subsequent layers to process and understand the relationships between words in the input sequence [27].
Self-attention is a mechanism that allows an input sequence to process itself in a way that each position in the sequence can attend to all positions within the same sequence. This mechanism is a cornerstone of the Transformer architecture, which has revolutionized NLP and beyond by enabling models to efficiently handle sequences of data with complex dependencies. The purpose of self-attention is to compute a representation of each element in a sequence by considering the entire sequence, thereby capturing the contextual relationships between elements regardless of their positional distance from each other. This ability to capture both local and global dependencies makes self-attention particularly powerful for tasks such as machine translation, text summarization, and sequence prediction, where understanding the context and the relationship between words or elements in a sequence is crucial [5]
The mathematical formulation of self-attention involves several key steps. First, a set of query vectors $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}^{Q}$, a set of key vectors $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}^{K}$, and a set of value vectors $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}^{V}$ are calculated through linear transformations of the input sequence, where $\boldsymbol{X}$ is the input matrix representing embeddings of tokens in a sequence, $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ are weight matrices for queries, keys, and values, respectively. The attention scores are then calculated by taking the dot product of the query vector with all key vectors, followed by scaling the result by the inverse square root of the dimension of the keys ($\sqrt{d_{k}}$) to avoid overly large values. These scores are then passed through a softmax function to obtain the attention weights, which represent the importance of each element’s contribution to the output. Finally, the output for each position is computed as a weighted sum of the value vectors,
$\boldsymbol{A(\boldsymbol{Q,K,V})}=\text{softmax}(\frac{\boldsymbol{QK^{\top}}}{\sqrt{d_{k}}})\boldsymbol{V}$ (12)
This process allows the model to dynamically focus on different parts of the input sequence, enabling the extraction of rich contextual information from the sequence.
(a) Transformer
(b) Self-Attention
Figure 5: Transformer architecture and its self-attention mechanism, adapted from [5].
2.1.3 Multi-Head-Attention
Multi-head attention is an extension of the self-attention mechanism designed to allow the model to jointly attend the information from different representation subspaces at different positions [5]. Instead of performing a single attention function, it runs the attention mechanism multiple times in parallel. The outputs of these independent attention computations are then concatenated and linearly transformed into the expected dimension. The mathematical formulation of multi-head attention can be described in four steps. First, Multi-Head Attention Computation,
$\boldsymbol{Q}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{Q},\quad\boldsymbol{K}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{K},\quad\boldsymbol{V}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{V}$ (13)
where i is the index of each self-attention head,
$\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})=\text{softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V}_{i}$ (14)
Then, concatenating outputs from all heads followed by a final linear transformation,
$\boldsymbol{A}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{Concat}(\boldsymbol{A}_{1},\boldsymbol{A}_{2},\ldots,\boldsymbol{A}_{h})\boldsymbol{W}^{O}$ (15)
where $\boldsymbol{W}^{O}$ is the weight matrix for the final linear transformation.
The multi-head attention mechanism enables the model to capture different types of information from different positions of the input sequence. By processing the sequence through multiple attention ”heads”, the model can focus on different aspects of the sequence, such as syntactic and semantic features, simultaneously. This capability enhances the model’s ability to understand and represent complex data, making multi-head attention a powerful component of Transformer-based architectures [28].
2.2 From Transformer to Vision Transformer
The journey from the inception of the Transformer model to the development of the Vision Transformer marks a pivotal advancement in deep learning, showcasing the adaptability of models initially designed for sequence data processing to the realm of image analysis. This transition underscores a significant shift in approach, from conventional image processing techniques to more sophisticated, sequence-based methodologies.
Introduced in 2017 by Vaswani et al. [5] through the seminal paper ”Attention Is All You Need,” the Transformer model revolutionized NLP by leveraging self-attention mechanisms. This innovation allowed for the processing of sequences of data without the reliance on recurrent layers, facilitating unprecedented parallelization and significantly reducing training times for large datasets. The Transformer’s success in NLP sparked curiosity about its potential applicability across different types of data, including images, setting the stage for a transformative adaptation.
The adaptation of Transformers for image data pivoted on a novel concept: treating images not as traditional 2D arrays of pixels but as sequences of smaller, discrete image patches. This approach, however, faced computational challenges due to the self-attention mechanism’s quadratic complexity with respect to input length. The breakthrough came with the introduction of the ViT by Dosovitskiy et al. [6] in 2020, which applied the Transformer architecture directly to images, see Figure 6. By dividing an image into fixed-size patches and processing these patches as if they were tokens in a text sequence, ViT was able to capture complex relationships between different parts of an image using the Transformer’s encoder.
The operational mechanics of ViT begin with the division of an input image into fixed-size patches, each of which is flattened and linearly transformed into a vector, effectively converting the 2D image into a 1D sequence of embeddings. To account for the lack of inherent positional awareness within the Transformer architecture, positional embeddings are added to these patch embeddings, ensuring the model retains spatial information. The sequence of embeddings is then processed through the Transformer encoder, which consists of layers of multi-head self-attention and feed-forward neural networks, allowing the model to dynamically weigh the importance of each patch relative to others for a given task.
For tasks like image classification, the output from the Transformer encoder is passed through a classification head, often utilizing a learnable ”class token” appended to the sequence of patch embeddings for this purpose. The model is trained on large datasets using backpropagation and, during inference, processes images through these steps to predict their classes.
The ViT not only demonstrates exceptional performance on image classification tasks, often surpassing CNNs when trained on extensive datasets, but also highlights the Transformer architecture’s capacity to capture the global context within images. Despite its advantages, ViT’s reliance on substantial computational resources for training and its need for large datasets to achieve optimal performance present challenges. Nonetheless, the development of ViT signifies a significant milestone in the application of sequence processing models to the field of computer vision, opening new avenues for research and practical applications.
Figure 6: The Vision Transformer architecture, adapted from [6].
2.2.1 Vision Transformer: From Spatial to Spatio-Temporal Tasks
The original ViT, designed for static image processing, divides images into patches and interprets these as sequences, leveraging the Transformer’s self-attention mechanism to understand complex spatial relationships. Extending this model to action recognition involves adapting it to analyze video frames sequentially to capture both spatial and temporal relationships. Several works attempted to adapt ViT in action recognition task using different methods as below.
Temporal Dimension Integration The integration of the temporal dimension is a fundamental step in adapting ViT for action recognition. Traditional ViT models process images as a series of patches, treating them essentially as sequences for the self-attention mechanism to analyze spatial relationships. By extending this concept to include the temporal dimension, the models can now treat videos as sequences of frame patches over time. This allows the models to capture the evolution of actions across frames. The work by Bertasius et al. [29] highlights the potential of incorporating temporal information into Transformers, marking a significant advancement in video analysis capabilities.
Spatio-Temporal Embeddings To effectively capture the dynamics of actions within videos, adapted ViT models generate spatio-temporal embeddings. This involves extending the traditional positional embeddings used in ViTs to also include temporal positions, thereby creating embeddings that account for both spatial and temporal information within video sequences. The discussion by Arnab et al. [30] on the creation of these spatio-temporal embeddings showcases the method’s effectiveness in enhancing the model’s understanding of action dynamics across both space and time.
Multi-Head Self-Attention Across Time The extension of self-attention mechanisms to analyze relationships between patches not just within individual frames but also across different frames is crucial for recognizing actions over time. This approach enables the model to identify relevant features and changes across the video sequences, facilitating a deeper understanding of motion and the progression of actions. The exploration by Bertasius et al. [29] of this concept demonstrates how Transformers can be effectively adapted to capture the temporal dynamics of actions, a key aspect of video analysis.
2.3 CNNs: From Spatial to Spatio-Temporal Tasks
CNNs have undergone a significant evolution since their introduction in the 1980s. The Neocognitron [31], developed by Kunihiko Fukushima, presented an early example of neural networks incorporating convolutional operations for image processing, setting the foundations for subsequent progress. Shortly after, Yann LeCun and collaborators introduced LeNet-5 [32], a key architecture designed for handwritten digit recognition, showcasing the effectiveness of convolutional layers in pattern recognition tasks.
The progress of CNNs reached a turning point in the mid-2010s with the introduction of models like AlexNet [33], showcasing their potential in image classification tasks. Alongside architectural innovations, this milestone was achieved thanks to access to large datasets, notably, ImageNet [34], and computational improvements, including the rise of Graphics Processing Units (GPUs) for parallel computing. Large-scale datasets provided the diversity and complexity necessary for training deep networks, while enhanced computational power accelerated the training of sophisticated CNN architectures.
The architectural enhancements, large datasets, and increased computational capabilities helped CNNs to be a cornerstone in deep learning methodologies, extending their applications beyond image processing to various domains. Notable architectures like VGGNet [35], distinguished by its uniform design and small convolutional filters, GoogLeNet [36], with its inception modules for capturing features at different scales efficiently, and ResNet [37], which introduced residual learning for training very deep networks, have further enriched the landscape of CNNs.
Transfer Learning. Deep Neural Networks, specifically CNNs, have been proven to work very well in the presence of large, diverse datasets. However, access to large representative datasets is a big challenge in the deep learning era. To benefit from the impressive pattern recognition capability of CNNs in the absence of diverse and large datasets, transfer learning comes into play to leverage knowledge gained from a source task to enhance the performance of a model on a related target task. Transfer learning involves pre-training a model on a large dataset through which CNNs’ early layers capture relatively general features and patterns that may be common and helpful for other related target tasks. By freezing the early layers to preserve the learned knowledge, and fine-tuning the later layers on a target task to allow the model to adapt to target task-specific features, the power of deep networks is exploited even when there is limited data. ImageNet is a widely recognised source dataset within the domain of computer vision. Several prominent architectural designs, including VGG, ResNet and GoogleNet have made their pre-trained weights on the Imagenet dataset publicly available to provide a source for subsequent researchers and practitioners. The availability of pre-trained weights accelerates the training process on diverse computer vision tasks, enabling the transfer of learned features and representations from the ImageNet dataset to target applications.
Object detection, another fundamental task in computer vision, has seen significant advancements in the last 10 years. Early methods such as Viola-Jones [38] employed handcrafted features and classifiers to detect and classify objects in a given image. The paradigm shift in the 2010s was marked by the introduction of R-CNN [39]: Region with CNN features and subsequent architectures like Fast R-CNN [39] and Faster R-CNN [40]. This pivotal development brought forth a two-stage approach, involving the proposal of regions of interest and their classification. Following that Single Shot MultiBox Detector (SSD) [41] and You Only Look Once (YOLO) [42] models that allow real-time object detection were proposed to show the efficacy of single-shot detection frameworks. One of the recent aims in the field is to achieve anchor-free methods [43, 44] to work with more flexible and efficient methods.
Segmentation. Similar to object detection, the early techniques for segmentation, a computer vision task that aims to partition an image into distinct and meaningful regions, involved handcrafted features. Early methods like thresholding and region-growing struggled with complex scenes and variations. Advancements in CNNs in the 2010s also skyrocketed the number of proposed architectures that use powerful CNNs and state-of-the-art performance was quickly surpassed by methods like Fully Convolutional Networks (FCN) [45] and U-Net [46]. FCN allowed pixel-level predictions while maintaining spatial information and U-Net helped with better feature integration with the introduction of skip connections. Skip connections, also known as shortcut connections or residual connections, were incorporated in the U-Net architecture to address the challenge of information loss during downsampling and upsampling operations. These connections enable the direct flow of information between corresponding encoder and decoder layers, facilitating the integration of low-level and high-level features. This design choice enhances the model’s ability to capture fine details and spatial information while maintaining pixel-level predictions, making it especially effective for image segmentation tasks.
Transfer learning is not limited to image classification tasks in the computer vision field. Object detection and segmentation tasks, also greatly benefit from the pre-trained weights to enhance model performance and efficiency. In object detection tasks, architectures like R-CNN [39], YOLO [42] and Mask R-CNN [47]; in segmentation tasks architectures like FCN [45] and PSPNet [48] trained on large datasets are employed as a starting point and fine-tuning happens on target tasks. In essence, transfer learning in deep learning, and particularly in CNNs, facilitates the transfer of knowledge, allowing models to achieve impressive results in diverse applications, even when faced with limited labelled data for a target task.
3 Deep Learning-Based Techniques in HAR
This section briefly recalls the most commonly used deep learning-based HAR approaches.
3.1 CNN-based approaches in HAR
This section recalls the most prominent CNN-based approaches in HAR, organized chronologically.
Deep learning was still in its early stages in 2012, and CNNs, or CNNs, had not yet gained significant popularity in the field of HAR. The focus was primarily on traditional machine learning approaches, such as Support Vector Machines (SVMs) [49], and handcrafted features, such as Histogram of Oriented Gradients (HOG) [50] and Histogram of Optical Flow (HOF) [51]. A few studies did, nevertheless, start looking into neural networks for action recognition.
In 2012, Ji et al. [52] presented an exploration into the use of 3D CNNs for recognizing human actions in surveillance videos. This model innovatively extends traditional 2D CNNs by performing 3D convolutions, thereby extracting features from both spatial and temporal dimensions to effectively capture motion information. Evaluated on the TREC Video Retrieval Evaluation (TRECVID) dataset and the KTH dataset, the 3D CNN model demonstrated significant improvements over traditional 2D CNNs and other baseline methods in action recognition tasks. The paper’s novelty lies in its unique approach to integrating spatial-temporal features for action recognition, offering enhanced performance in real-world surveillance scenarios. For papers specifically from 2012 that only used CNN architectures for HAR, there may not be a significant number of high-impact papers due to the fact that the use of CNNs in this field was still emerging and not as established as in later years. The seminal works that laid the groundwork for using CNNs in action recognition often came slightly later, as the computational power and dataset availability became more conducive to deep learning approaches.
In 2013, Valle et al. [53] address the challenge of HAR from video data, focusing specifically on the recognition of walking and running actions. They propose an innovative approach using a variant of a 2D CNN, a type of multilayer perceptron, which is distinct for its ability to learn feature extraction directly from raw, unprocessed images during the error backpropagation phase. This methodology represents a significant advancement over traditional methods that rely on pre-processed images. The authors developed a specialized dataset comprising 4000 images per activity to train and evaluate their model. The paper’s contribution lies in its novel approach to feature extraction in HAR, demonstrating the potential of CNNs in efficiently and accurately recognizing complex human actions directly from video sequences, a task of growing importance in fields such as surveillance and automated monitoring systems.
In 2014, the use of CNNs in action recognition was at a pivotal stage, marking a shift from hand-crafted feature-based methods to deep learning approaches. The key points of the use of CNNs in action recognition at that period of time is the following. Emergence of Deep Learning, deep learning, particularly CNNs, had started to dominate image classification tasks, thanks to their ability to learn feature representations directly from raw pixel data. This success in static images paved the way for applying CNNs to video data for action recognition. Challenges in Video data, unlike 2D images, videos incorporate a third dimension which represents the temporal patterns, making action recognition more complex. CNNs had to be adapted to not only recognize spatial patterns but also capture motion information over time dimension. Datasets and Benchmarks, the adoption of large-scale video datasets like UCF-101 [54] and HMDB-51 [55] became more common. These datasets provided diverse sets of actions and were large enough to train deep networks. The performance on these benchmarks has been becoming a key measure of progress for action recognition models. Transfer Learning, due to the computational expense of training CNNs from scratch and the relatively smaller size of video datasets compared to image datasets, transfer learning became a popular strategy. Networks pre-trained on large image datasets like ImageNet [34] were fine-tuned on video frames for action recognition tasks. Computational Constraints, despite the promise of CNNs, computational constraints were a significant challenge. Training deep networks required significant GPU power, and processing video data with CNNs was resource-intensive. This limited the complexity of the models that could be trained and the size of the datasets that could be used.
In 2014, Simonyan et al. [35] presented an innovative approach to recognizing actions in video sequences by using a two-stream CNN architecture. This approach divides the task into two distinct problems: recognizing spatial features from single frames and capturing temporal features across frames. The spatial stream CNN processes static visual information, while the temporal stream CNN handles motion by analysing optical flow. The model was tested on benchmark datasets like UCF-101 and HMDB-51, where it achieved state-of-the-art results, showcasing the effectiveness of this dual-stream method. The novelty of this work lies in the separation of motion and appearance features, which allows for more specialized networks that can better capture the complexities of video-based action recognition. The success of this model has made a significant impact on the field, influencing many future research directions in video understanding.
Sun et al. [56] presented a novel approach to action recognition by integrating deep learning with Slow Feature Analysis (SFA). SFA is a computational technique used to extract features from a signal that change slowly over time. It is based on the principle that while the sensory data may change rapidly, the underlying causes of the data often change at a slower rate. By focusing on these slowly varying features, SFA aims to capture the inherent properties of the observed data that are more stable and consistent over time. In the context of visual data, such as video sequences, SFA would aim to identify features that remain constant or change slowly as the video progresses. These could be the overall shape of an object, its identity, or its trajectory through space, as opposed to the rapidly changing texture or lighting conditions. The authors developed a CNN that processes video data to extract a hierarchy of complex features for action recognition tasks. The research stands out for applying a deep learning structure to SFA, which traditionally uses a single-layer approach. This deepened SFA is shown to be superior in learning more generalizable features from video sequences. The model was evaluated on Hollywood2, KTH, and UCF Sports datasets, which are standard benchmarks with considerable intra-class variation challenges. The paper reports that the deep learning-enhanced SFA model surpasses traditional methods and some state-of-the-art techniques in accuracy, without relying on additional processing tricks such as dense sampling. This suggests that the model’s performance is due to its robust feature learning capabilities.
Ijjina et al. [57] presented a novel approach to HAR by leveraging the patterns found in action bank features through a specially designed CNN. Action bank features are computed by considering a fixed set of videos known as an action bank. These videos serve as templates to measure the similarity of new videos against them. The CNN is uniquely tailored with specific masks to capture these patterns within the UCF50 dataset. The novelty lies in this approach of using action bank features for action recognition. The results show that the CNN effectively learns from the data, with a decreasing error rate over iterations, indicating good generalization. The paper suggests the potential for applying this method to other datasets and exploring different CNN models.
In 2015, the use of CNNs in action recognition tasks was at the forefront of research in computer vision, particularly due to their ability to learn hierarchical feature representations directly from raw data. The field was rapidly evolving, with new architectures and techniques being proposed to leverage the power of CNNs for this complex task.
Cheron et al. [58] presented an innovative approach to action recognition in videos. The core contribution is the development of the Pose-based CNN (P-CNN) descriptor, which uniquely integrates motion and appearance information along the trajectories of human body parts. This methodology marks a departure from traditional action recognition techniques that predominantly rely on statistical representations of local motion descriptors. By focusing on the detailed analysis of individual body parts through CNNs, the P-CNN model demonstrates a heightened sensitivity to subtle and fine-grained actions, a notable advancement in the field. The paper validates the efficacy of this approach using challenging datasets like JHMDB and MPII Cooking Activities, where it shows a marked improvement over existing methods, underscoring the potential of pose-based features in enhancing the accuracy of action recognition systems.
Geng et al. [59] addressed the limitations of traditional classifiers that rely on complex, handcrafted features, the authors propose a model that utilizes CNNs to directly process raw video inputs. A key innovation of their work is the introduction of a pre-training strategy employing a Convolutional Auto-Encoder (CAE), aimed at reducing the computational burden associated with kernel training. This method not only facilitates more efficient training but also enhances the model’s ability to generalize from raw data, circumventing the need for manually designed features. Evaluated on the KTH action database, the proposed model demonstrates superior performance in recognizing human actions compared to existing algorithms that depend on handcrafted features, showcasing its potential for real-world applications in video surveillance, virtual reality, and intelligent human-computer interactions. This approach marks a significant advancement in the field of computer vision and deep learning, offering a robust solution to the challenges of HAR.
Du et al. [60] propose a novel, CNN-based framework for action recognition, leveraging the dynamics of human skeletal movements. This research departs from traditional methods by transforming skeleton sequences into matrices, which are then converted into images, encapsulating both spatial structures and temporal dynamics. These images are fed into a specially designed CNN, where simple max-pooling plays a crucial role in spatial feature selection and temporal frequency adjustment. This approach not only preserves the intricate motion characteristics of each body part but also addresses the challenges posed by variable sequence lengths and action frequencies. The model’s efficacy is demonstrated through its superior performance on the Berkeley MHAD [61] and the ChaLearn gesture recognition [62] datasets, where it achieves state-of-the-art results, particularly excelling with a significant margin over existing methods in terms of computational efficiency and accuracy. This advancement in skeleton-based action recognition represents a significant leap forward, offering robust, efficient solutions for applications in human-computer interaction, surveillance, and beyond.
Gkioxari et al. [63] introduced a novel approach to action recognition in static images, significantly enhancing the accuracy by incorporating contextual cues from the surrounding environment. The core innovation lies in the adaptation of the Region-based Convolutional Network (RCNN) to the RCNN model, which uniquely utilizes both a primary region, encompassing the person performing the action, and a secondary region that automatically identifies relevant contextual information. This dual-region approach allows R*CNN to capture a more comprehensive understanding of the action by considering additional elements like objects, other people, and the scene itself. The model’s effectiveness is demonstrated through its remarkable performance on the PASCAL VOC Actions dataset, where it achieves a mean AP of 90.2%, significantly surpassing existing methods. Additionally, its versatility is showcased by its application to the MPII Human Pose dataset and the Berkeley Attributes of People dataset, where it achieves state-of-the-art results in attribute classification. This paper’s contribution is pivotal, as it not only advances the field of action recognition by integrating contextual understanding into the recognition process but also opens avenues for fine-grained classification tasks, underscoring the model’s adaptability and robustness.
The year 2016 witnessed a surge in the use of CNN models for action recognition tasks. Numerous models emerged, employing multiple streams and modalities. Additionally, there was a notable development of models utilizing 3D-CNNs, as well as hybrid models that combined 2D-CNNs with RNNs to analyze temporal information in videos.
Bilen et al. [64] introduced a groundbreaking approach to video analysis, particularly for action recognition, using CNNs. The central innovation is the concept of ”dynamic images,” a compact video representation that encodes the temporal evolution of video frames into a single RGB image. This is achieved through an efficient process known as rank pooling. The paper presents a novel static/dynamic neural network architecture that enables end-to-end training from videos. This architecture uniquely combines static appearance information from still frames with dynamic information from the entire video sequence. A significant aspect of this model is the introduction of an approximate rank pooling CNN layer, which greatly speeds up the process compared to traditional rank pooling. The paper claims state-of-the-art performance in action recognition on standard benchmarks (HMDB51 and UCF101). The dynamic images primarily focus on the acting objects in the video, effectively summarizing actions and motions. However, the compression of video content into a single image, while efficient, might oversimplify complex actions or subtle nuances in videos.
Zhang et al. [65] introduces a significant advancement in the field of video-based action recognition by optimizing the deep two-stream architecture for real-time processing. The core innovation lies in replacing the computationally intensive optical flow with motion vectors, which are readily available from compressed videos, thus eliminating the need for additional calculations. However, motion vectors inherently lack the fine-grained structural details and contain noisy motion patterns, which could degrade recognition performance. To address this, the authors propose a novel approach of transferring knowledge from an optical flow CNN (OF-CNN) to a motion vector CNN (MV-CNN). This transfer is achieved through three strategies: initialization transfer, supervision transfer, and a combination of both. By doing so, the enhanced MV-CNN effectively learns from the more accurate motion patterns captured by OF-CNN, leading to a significant boost in performance. The paper demonstrates that this method achieves comparable recognition accuracy to state-of-the-art techniques while processing at an impressive speed of 390.7 frames per second, which is 27 times faster than the original two-stream method. This breakthrough not only provides a feasible solution for real-time action recognition but also opens up new possibilities for efficient video analysis in various applications. The evaluation of this method on large datasets like UCF101, HMDB51, and THUMOS14 further underscores its effectiveness and practical utility in real-world scenarios.
Rahmani et al. [66] introduced a novel approach for recognizing human actions in 3D from various viewpoints, a significant challenge in video-based HAR due to factors like viewpoint variations, illumination, and visual appearance. Central to their methodology is a deep CNN that transforms human poses from different, unknown views into a view-invariant high-level space. Uniquely, the training data for this model is synthetically generated by fitting 3D human models to real motion capture data and rendering these poses from multiple viewpoints, circumventing the need for extensive and expensive multiview training data. This approach is innovative in its use of synthetic data for training while maintaining the ability to generalize to real depth images of unseen poses without re-training. The paper also introduces a group sparse Fourier Temporal Pyramid for robust spatio-temporal representation, enhancing the discriminative power of the model for action recognition. The proposed method demonstrates significant improvements over existing methods, achieving higher accuracies on both multiview and single-view benchmark datasets, thereby marking a substantial advancement in the field of 3D action recognition.
Xin et al. [67] introduces a novel approach to action recognition in video, addressing the complexities of real-world environments. The core of the paper is the development of the Adaptive Recurrent-Convolutional Hybrid (ARCH) Networks, a hybrid deep neural network (DNN) that integrates static, dynamic, and sequential mixed features for robust action recognition. This approach is designed to tackle three key challenges: spatial domain variations (such as noise and occlusion), temporal domain polytropes (differences in action velocities and durations), and intra- and inter-class diversities (similarities and differences within and between action categories). The ARCH model, evaluated on the UCF101 dataset and further validated through transfer learning on other datasets, demonstrates superior performance in recognizing human actions, outperforming existing deep-learning-based methods. This advancement is significant in its comprehensive handling of the spatial, temporal, and class diversity challenges in action recognition, offering a more effective and adaptable solution for long-term action recognition in uncontrolled and realistic environments.
By 2017, CNNs had become a key technology in the field of action recognition within computer vision. Their strength lay in their ability to autonomously extract meaningful spatial features from images or video frames. However, the complexity of action recognition required not just spatial but also temporal analysis. This led to the integration of CNNs with architectures capable of temporal processing, like RNNs or LSTMs, or the development of specialized CNN variants like 3D CNNs. Challenges such as handling diverse and complex video data, as well as the computational demands of training these models, were significant. Nonetheless, the advancements in CNN-based action recognition systems opened up numerous practical applications, ranging from surveillance to interactive gaming, and continued to be a vibrant area of research and development in the field of computer vision.
Zhao et al. [68] introduces a novel approach for action recognition in 3D videos, a field with significant applications in health monitoring, assisted living, and surveillance. The study’s core innovation is the integration of RNNs and CNNs in a unique two-stream fusion method. The RNNs, utilizing gated-recurrent-units, excel in processing long-term information from optical tracking data, while the 3D-CNNs are adept at handling detailed, recent information from video data. This combination allows for a comprehensive analysis of both spatial and temporal data aspects. The paper employs the NTU RGB+D Dataset, one of the largest datasets for 3D action recognition, containing over 56,000 action samples in 60 different classes. The novelty of the research lies in its two-stream fusion approach, which merges the outputs of RNN and CNN models, significantly enhancing action recognition performance. This approach includes two fusion structures: decision fusion and feature fusion, with the latter demonstrating superior results. The proposed method shows a remarkable improvement over state-of-the-art methods, outperforming them by more than 14% in both cross-subject and cross-view settings. Additionally, the novel RNN structure proposed in the study is optimized for faster convergence and reduced computational demand, making it a significant advancement in the field of action recognition in 3D videos.
Ullah et al. [69] presents an innovative approach to action recognition in video sequences, a challenging area in computer vision. The core of their methodology is the integration of Deep Bidirectional Long Short-Term Memory (DB-LSTM) networks with CNN features. This combination is designed to effectively capture and analyze the sequential information in video frames, addressing key challenges such as visual content similarity, changes in viewpoint, and varying illumination conditions. The model utilizes deep features extracted from every sixth frame using a pre-trained AlexNet, significantly reducing redundancy and computational complexity. The method is rigorously evaluated on three benchmark datasets: UCF-101, YouTube 11 Actions, and HMDB51, demonstrating its effectiveness in recognizing a wide range of actions. The novelty of this research lies in its efficient use of DB-LSTM for processing sequential frame changes in conjunction with CNN for deep visual data analysis, making it particularly adept at handling lengthy videos. The experimental results show that this approach significantly outperforms existing state-of-the-art methods in action recognition, highlighting its potential for applications in areas like surveillance, human-computer interaction, and sports analysis.
Li et al. [70] introduced an innovative approach for action recognition in 3D skeleton videos. The core of this research is the development of a method that transforms 3D skeleton data into color images, ensuring translation-scale invariance and independence from the dataset. This transformation is crucial as it effectively handles variations in translation and scale, common challenges in 3D skeleton data analysis. The study then employs a multi-scale deep CNN, built upon pre-trained models like ResNet, to classify these transformed images. This multi-scale CNN is specifically designed to enhance the model’s ability to adjust temporal frequencies, a key aspect when dealing with action recognition. The methodology is rigorously tested on the NTU RGB-D dataset, known for its complexity and diversity, and the results demonstrate that the proposed approach significantly outperforms existing methods. The novelty of this work lies in its unique image mapping technique and the application of a multi-scale deep CNN, making it a substantial contribution to the field of action recognition from 3D skeleton videos.
Varol et al. [71] introduced a novel approach to action recognition in videos through the use of Long-term Temporal Convolutions (LTC) within CNNs. Addressing the limitation of existing methods that typically analyze actions over short video intervals, this study proposes a model that captures the full temporal extent of human actions, which often last several seconds and encompass a wide range of frames. The LTC-CNN model comprises five space-time convolutional layers followed by three fully connected layers, designed to extend the temporal reach of the network while balancing complexity by reducing spatial resolution. This approach is evaluated on two challenging HAR benchmarks, UCF101 and HMDB51, known for their diverse and complex action categories. The paper’s innovation lies in its focus on long-term temporal patterns, a critical aspect often overlooked in action recognition. The study not only demonstrates the effectiveness of LTC in capturing extensive temporal information but also underscores the importance of high-quality optical flow estimation for accurate video representation. The results show state-of-the-art performance on both datasets, highlighting the model’s superiority in handling the temporal dynamics of human actions.
In 2018, Zhigang Tu et al. [72] introduces a novel approach to video-based HAR using a multi-stream CNN architecture. This model, termed Human-Related Multi-Stream CNN (HR-MSCNN), innovatively combines three motion-related and three appearance-related streams to analyze human actions in videos. The model employs Block-Sparse Robust Principal Component Analysis (B-RPCA) for effective human detection, addressing challenges like background motion and illumination changes, and uses a motion saliency measure to refine foreground regions and identify primary moving body parts. Evaluated on benchmark datasets such as JHMDB, HMDB51, UCF Sports, and UCF101, the HR-MSCNN demonstrates state-of-the-art performance. The paper’s novelty lies in its end-to-end CNN that considers both the detected person and discriminative motion regions, a new method for obtaining static and dynamic features, and a fusion method to efficiently combine information from six streams for action prediction. The findings highlight the model’s enhanced ability to focus on the most informative parts of the human body for accurate action recognition, marking a significant advancement in the field.
Arunnehru et al. [73] introduced an approach to improving surveillance systems through the application of 3D-CNNs integrated with 3D motion cuboids. This method is designed to enhance the accuracy of suspicious action recognition in video surveillance, a critical aspect of ensuring public safety and preventing potential crimes. By utilizing the spatial-temporal data encapsulated in 3D motion cuboids and processing this information through 3D-CNNs, the model is capable of detecting and recognizing complex human actions with higher precision than previously established methods. The research is validated through extensive testing on two benchmark datasets, KTH and Weizmann, where it demonstrated superior performance in terms of accuracy, thereby outperforming existing action recognition techniques. This advancement not only marks a significant contribution to the fields of intelligent video surveillance and computer vision but also opens new avenues for the development of more sophisticated and reliable surveillance systems capable of preemptively identifying and mitigating risks associated with suspicious activities.
Miao Ma et al. [74] introduced a novel approach for recognizing both general and fine-grained human actions in video sequences using a six-stream CNN. This method significantly focuses on fine-grained actions, which are more subtle and occur in smaller image regions compared to general actions. The core of the approach involves estimating human poses and body parts positions in video sequences, followed by cropping image patches at various scales to capture diverse appearance and motion cues. Uniquely, the model employs the outputs of the pooling layer of the CNN, which retains more spatial information than the fully connected layers typically used in such analyses. This is complemented by an effective encoding method that generates discriminative action descriptors for classification. The model’s efficacy is demonstrated on two public datasets, the Sub-JHMDB and MPII Cooking datasets, where it outperforms several state-of-the-art methods. The novelty of this work lies in its combination of coarse human pose tracking with multi-scale patch extraction and advanced pooling layer encoding, making it a significant contribution to the field of fine-grained action recognition in videos.
Zhang et al. [75] introduces a novel approach to real-time action recognition in videos. Addressing the limitations of traditional two-stream CNNs, which are hindered by the time-consuming computation of optical flows (OFs), this study proposes the use of motion vectors (MVs) as a more efficient alternative. MVs, inherent in video streams, can be extracted with minimal computational overhead, offering a pathway to real-time processing. However, MVs inherently contain noise and lack the fine details present in OFs, leading to potential accuracy degradation when used directly in CNNs. To overcome this, the paper presents four innovative training strategies that leverage the knowledge from OF-based CNNs to enhance the performance of MV-based CNNs. This knowledge transfer is rooted in the structural similarities between MV and OF, enabling the application of insights gained from OF CNNs to improve the generalization ability of MV CNNs. The core of this approach is the deeply transferred MV CNN, which employs a method called ”Deeply Connected Transfer” for multi-layer knowledge transfer from OF CNN to MV CNN. This method significantly boosts the accuracy of MV CNNs while maintaining the high-speed processing capabilities essential for real-time action recognition. The effectiveness of this approach is demonstrated through extensive experiments on various datasets, including UCF101, HMDB51, and THUMOS14, showcasing the model’s ability to achieve high accuracy and processing speed, thus marking a significant advancement in the field of real-time video-based action recognition.
Liu et al. [76] introduced a novel architecture for efficient and real-time action recognition in videos. The core of this architecture is the Temporal Convolutional 3D Network (T-C3D), which innovatively combines a residual 3D CNN with a temporal encoding method to capture both the appearance of individual frames and the motion between consecutive frames in a video. This approach addresses the computational inefficiencies of existing methods, particularly those reliant on optical flow extraction, by focusing on hierarchical multi-granularity learning. The T-C3D model is evaluated on two widely recognized action recognition datasets, UCF101 and HMDB51, where it demonstrates significant improvements in both accuracy and processing speed compared to state-of-the-art real-time methods. Notably, the T-C3D achieves over 5.4% higher accuracy and is twice as fast in terms of inference speed, reaching up to 969 frames per second. The model’s ability to process RGB frames in real-time while maintaining high accuracy and capturing both short-term and long-term temporal dynamics marks a substantial advancement in the field of video-based action recognition. The paper’s contribution lies in its novel integration of 3D-CNN with temporal encoding, enabling efficient modeling of video dynamics and facilitating real-time application potential.
In 2019, Jiang et al. [77] an innovative approach for video action recognition by efficiently integrating spatiotemporal and motion features within a unified 2D CNN framework, a significant departure from the traditional reliance on 3D convolution and optical flow pre-calculation. Central to this approach is the STM network, which comprises two novel components: the Channel-wise SpatioTemporal Module (CSTM) and the Channel-wise Motion Module (CMM). The CSTM focuses on capturing spatiotemporal features through channel-wise 1D convolution, allowing for effective temporal information fusion with minimal computational overhead. In contrast, the CMM is designed to encode motion features by learning distinct motion characteristics for each channel between adjacent time steps. This method was rigorously tested on various datasets, including Something-Something v1 & v2, Jester, Kinetics-400, UCF-101, and HMDB-51, demonstrating superior performance over existing state-of-the-art methods. The novelty of this research lies in its unique integration of spatiotemporal and motion features in a 2D framework, efficiently handling different types of action recognition scenarios and showcasing its potential in practical applications.
Crasto et al. [78] introduced a groundbreaking approach to action recognition by ingeniously integrating motion information into a standard 3D CNN that processes RGB frames, thereby circumventing the computationally intensive and latency-prone process of calculating optical flow, a common necessity in conventional two-stream action recognition frameworks. Central to this innovation are two novel learning strategies: MERS (Motion-Emulated RGB Stream) and MARS (Motion-Augmented RGB Stream). MERS trains the network to emulate features from the motion stream, significantly reducing the computational burden, while MARS further advances this concept by synergistically combining appearance and motion information into a singular, efficient stream. This approach not only achieves superior performance, outstripping individual RGB and Flow streams with 72.7% accuracy on the Kinetics dataset and 80.1% on HMDB51 (split-1), but also maintains the high standards of two-stream architectures without their inherent computational inefficiencies. The paper’s innovative use of a feature-based loss, alongside the standard cross-entropy loss, for network training marks a significant leap forward, showcasing the effective integration of temporal information into a 3D CNN and setting a new benchmark for future research in action recognition.
Hussein et al.[79] introduced a novel framework aimed at enhancing the recognition of complex human activities in videos. Recognizing the limitations of traditional spatiotemporal 3D convolutions, which struggle with the variability and length of complex actions, the authors propose Timeception convolution layers. These layers are uniquely designed to handle minute-long temporal patterns, extending the temporal modeling capacity up to 1024 timesteps, significantly surpassing existing methods. Timeception achieves this by employing multi-scale temporal convolutions and reducing the complexity inherent in 3D convolutions, making it particularly adept at capturing the nuances of complex actions that consist of multiple, variably-timed one-actions. The model’s efficacy is demonstrated through impressive accuracy on challenging datasets like Charades, Breakfast Actions, and MultiTHUMOS, marking a substantial advancement in the field of action recognition. The paper’s innovative approach in decomposing and efficiently managing the temporal aspect of videos paves the way for more nuanced and accurate video analysis, particularly in understanding and interpreting complex, multi-faceted human activities.
He et al. [80] presents a sophisticated approach to action recognition in videos through the novel StNet architecture, which adeptly captures both local and global spatial-temporal relationships. StNet ingeniously stacks successive video frames into a super-image and applies 2D convolution to efficiently encode local spatial-temporal information, a method that contrasts with traditional CNN+RNN or pure 3D convolution-based approaches. To grasp the global spatial-temporal structure, the architecture incorporates temporal convolutions and introduces the innovative Temporal Xception Block (TXB), which separately processes channel-wise and temporal-wise convolutions, optimizing the model for size and computational efficiency. Evaluated on the Kinetics dataset, StNet demonstrates superior performance over existing 2D and 3D convolution-based solutions, striking a commendable balance between recognition accuracy and model complexity. Furthermore, the model’s robustness and generalizability are underscored by its successful application to the UCF101 dataset, showcasing its potential in advancing the field of video-based action recognition.
Hao Yang et al. [81] presented a novel approach to action recognition, addressing the computational and storage challenges associated with traditional 3D CNNs. The authors introduce an innovative model that utilizes efficient asymmetric one-directional 3D convolutions to approximate the conventional 3D convolutions, significantly reducing the number of parameters and computational cost. Central to the model are MicroNets, a series of local 3D convolutional networks incorporating multi-scale 3D convolution branches, designed to enhance the feature learning capacity of asymmetric 3D convolutions without increasing computational demands. The model also features a multi-source enhanced input method, which fuses information from RGB and Flow frames at the pre-processing stage, thereby circumventing the need for training separate networks for each frame type and further reducing computational requirements. Evaluated on two challenging benchmarks, UCF-101 and HMDB-51, the asymmetric 3D-CNN model not only outperforms traditional 3D-CNN models in terms of efficiency and effectiveness but also achieves comparable results with state-of-the-art action recognition methods, marking a significant advancement in the field of computer vision.
In 2020, Sudhakaran et al. [82] introduces a novel approach to video action recognition by proposing the Gate-Shift Module (GSM), a lightweight and efficient module designed to transform standard 2D CNNs into powerful spatio-temporal feature extractors. The GSM innovatively applies spatial gating to the spatial-temporal decomposition of 3D kernels, allowing the network to adaptively route and combine features through time with minimal additional computational cost and parameter overhead. This approach addresses the challenges of the high dimensionality and computational complexity typically associated with 3D CNNs in video action recognition. The paper demonstrates the effectiveness of the GSM through extensive evaluations, where the model achieves state-of-the-art results on the Something Something-V1 and Diving48 datasets and competitive results on the EPIC-Kitchens dataset, all while maintaining a significantly lower model complexity compared to existing methods. This breakthrough presents a significant advancement in the field, offering a more efficient and scalable solution for video action recognition tasks.
Liu x et al. [83] investigated the application of the VGG-16 Convolutional Neural Network model to the recognition of human actions from video data. This research employs deep learning techniques to analyze kinematic behavior, utilizing two datasets: the KTH dataset, featuring actions like walking and boxing, and the Weizmann dataset, which includes activities such as bending and waving. The study’s novelty lies in adapting the VGG-16 model, typically used for static image recognition, to dynamically recognize human behaviors, highlighting its efficacy compared to traditional methods like the histogram of optical flow orientation. Findings indicate that the VGG-16 model significantly outperforms other methods, achieving an average accuracy of 91.93% on the KTH dataset and 95.75% on the Weizmann dataset, with particularly high accuracy in recognizing boxing and bending movements. These results underscore the potential of advanced CNN models in the real-time analysis of complex human movements, paving the way for improvements in fields such as sports analytics and health monitoring.
Xiong et al. [84] presented a novel approach to enhance Human-Robot Collaboration (HRC) in smart manufacturing by improving the accuracy and reliability of HAR. Recognizing the challenges in effectively leveraging temporal information and the scarcity of large training datasets in manufacturing environments, the paper introduces an integrated method that combines optical flow and CNN-based transfer learning. Optical flow images, encapsulating the temporal dynamics of human motion, are fed into a two-stream CNN structure, meticulously designed to parse both spatial and temporal information concurrently. To circumvent the issue of limited training data, the study explores transfer learning, enabling the adaptation of a pretrained CNN to the specific context of manufacturing scenarios. The efficacy of this innovative approach is validated through its application to engine block assembly tasks, showcasing its potential in advancing the capabilities of HRC systems and contributing significantly to the evolution of smart manufacturing practices. The paper’s novel integration of optical flow with a two-stream CNN, augmented by transfer learning, marks a substantial stride in addressing the intricate demands of HAR within the realm of industrial automation.
Sahoo et al. [85] innovates in HAR by merging depth information with sequential learning. The HAR-Depth model integrates a Deep Bidirectional Long Short Term Memory (DBiLSTM) for capturing the temporal dynamics of actions and employs depth history images (DHI) to enhance shape recognition, where depth-estimated frames are projected onto the X-Y plane. Importantly, the model harnesses CNNs for initial feature extraction from action frames, which are then processed through the DBiLSTM. Evaluated on datasets such as KTH, UCF Sports, JHMDB, UCF101, and HMDB51, the model demonstrated superior accuracy, notably achieving 97.67% on KTH and 95.00% on UCF Sports. This approach highlights the model’s capability to distinguish similar actions through an effective combination of depth-based shape information and sophisticated temporal sequence processing, positioning it as a potent solution for advanced surveillance and interactive systems.
Abdelbakya et al. [86] introduced an approach to HAR using the Principal Component Analysis Network (PCANet). This research addresses the computational inefficiencies and the extensive training data requirements commonly associated with traditional CNNs in the context of HAR. By employing PCANet, a simplified deep learning model that utilizes Principal PCA for filter learning in convolutional layers, the authors propose a method that significantly reduces computational complexity while maintaining high accuracy in action recognition tasks. The methodology involves selecting a subset of frames from each video, extracting feature vectors via PCANet, reducing dimensionality through Whitening Principal Component Analysis (WPCA), and classifying actions using Support Vector Machines (SVM). The effectiveness of this approach is demonstrated on the KTH Human Action Dataset, showcasing superior performance compared to existing methods. This paper not only presents a computationally efficient alternative for HAR but also opens new avenues for applying simplified deep learning architectures in real-time video analysis and surveillance applications.
Jaouedi et al. [87] intoduced a hybrid deep learning model for human action recognition that incorporates Gated Recurrent Neural Networks (GRNN) as a core component. While the primary focus is on the integration of GRNN, Gaussian Mixture Models (GMM), and Kalman Filters (KF) for motion detection and tracking, the model is rooted in deep learning principles similar to those underlying CNNs. The approach leverages the sequential data processing capabilities of GRNNs, which are a variant of RNNs, to capture temporal dependencies in video data for action recognition. Although the paper emphasizes GRNN for its ability to handle temporal sequences effectively, the overall architecture aligns with deep learning methodologies, including those used in CNNs, for feature extraction and classification tasks in video analysis. This innovative approach aims to overcome the inherent challenges in HAR, such as variations in viewpoint, background clutter, and environmental conditions, by leveraging the temporal processing strengths of GRNNs for sequential data, the motion detection capabilities of GMMs, and the object tracking precision of KFs. Evaluated on three challenging datasets—UCF Sports, UCF101, and KTH—the model demonstrates its effectiveness by achieving an impressive average accuracy of 96.3% on the KTH dataset, highlighting its potential applicability in a wide range of domains, from intelligent video surveillance to human-computer interaction. This research marks a significant advancement in the field of computer vision by proposing a comprehensive solution that not only enhances the accuracy of human action recognition but also opens new avenues for future exploration in deep learning and video analysis technologies.
Asghari et al. [88] introduced a groundbreaking approach to HAR by intricately capturing the dynamics of human body motion in videos, a critical aspect that has been underexplored in previous research. Central to this approach is a novel dynamic encoder model that processes pose heat-map sequences, extracted from video frames using a pose estimator, to generate a fixed-size representation called DynaMotion. This representation encapsulates the temporal evolution of body joint heat-maps, providing a rich, comprehensive video-level representation for action classification. The authors demonstrate the efficacy of this method by integrating the DynaMotion representation with appearance and motion streams in a two-stream network architecture, leading to superior performance on several benchmark datasets including HMDB, JHMDB, UCF-101, and AVA. The paper’s innovative approach of leveraging the dynamical information from each pixel in the video marks a significant advancement in the field of HAR, showcasing the potential of combining appearance and motion information with temporal dynamics for enhanced action recognition accuracy.
In 2021 Wang et al. [89] introduced a novel approach to video action recognition by embedding a multipath excitation module, named ACTION, into 2D CNNs. This innovative module addresses the computational limitations of 3D CNNs while effectively capturing spatial-temporal, channel-wise, and motion patterns crucial for action recognition. The ACTION module comprises three distinct paths: Spatio-Temporal Excitation (STE) for characterizing spatio-temporal representations using 3D convolution, Channel Excitation (CE) for adaptively recalibrating channel-wise feature responses, and Motion Excitation (ME) for highlighting motion-sensitive channels by calculating feature-level temporal differences. The integration of these paths into the ACTION-Net model allows for a comprehensive and efficient analysis of video data, leading to significant performance improvements on three benchmark datasets: Something-Something V2, Jester, and EgoGesture. The paper demonstrates that ACTION-Net, equipped with the ACTION module, consistently outperforms its 2D CNN counterparts, offering a computationally efficient yet highly effective solution for action recognition in videos.
Wu et al. [90] presented an approach to video-based action recognition by extending 3D CNNs to not just RGB data but also depth and pose data modalities. Recognizing the limitations of single-modality (RGB) in capturing the comprehensive dynamics of actions, the authors propose a novel multimodal two-stream 3D network framework. This framework innovatively integrates Depth Residual Dynamic Image Sequence (DRDIS) and Pose Estimation Map Sequence (PEMS) as two distinct yet complementary data streams. DRDIS effectively encapsulates the spatial-temporal progression of actions in depth videos, while PEMS focuses on intuitively representing the spatial configuration of body parts, free from the distractions of cluttered backgrounds. The fusion of these streams aims to harness the rich, multimodal data, thereby significantly enhancing the action recognition performance. The model’s superiority is empirically validated through extensive experiments on four challenging datasets, namely SBU Kinect Interaction, UTD-MHAD, and NTU RGB+D 60 & 120, showcasing its potential to set new benchmarks in the domain of video-based action recognition. The paper’s novelty lies in its pioneering application of 3D CNNs to multimodal action data, marking a significant stride in the field and opening new avenues for future research.
Lee et al. [91] introduced an advanced video prediction model that integrates Long-term Motion Context Memory (LMC-Memory) with a memory alignment learning process, significantly improving the capability to utilize long-term motion contexts from brief input sequences. A notable aspect of the model’s architecture is the use of 3D Convolutional Networks (C3D) to extract motion features, which are then processed through this novel memory system to align and recall dynamic contexts effectively. The model was evaluated using both synthetic (Moving-MNIST) and real-world (KTH Action and Human 3.6M) datasets, demonstrating its superior performance in predicting future frames accurately, particularly in scenarios that require understanding extended activities from limited data. This innovation addresses a critical gap in traditional video prediction techniques, which often struggle with long-term temporal dependencies, by enhancing the model’s ability to predict complex motions over time through the strategic integration of CNNs with memory-based learning.
Wang et al. [92] introduced a novel video architecture, the Temporal Difference Network (TDN), aimed at enhancing action recognition in videos by efficiently capturing multi-scale temporal information. Addressing the challenges of temporal modeling, the TDN employs an innovative Temporal Difference Module (TDM) that operates on two levels: local and global motion modeling. This approach refines motion patterns for 2D CNNs using temporal differences over consecutive frames and captures long-range structures through temporal differences across segments. The TDN framework is designed to be integrated seamlessly with existing CNNs, offering a principled and efficient solution for action recognition. The paper demonstrates the model’s superiority through extensive experiments on the Something-Something V1 & V2 and Kinetics-400 datasets, where TDN sets new benchmarks, showcasing its ability to capture nuanced motion patterns and long-range temporal structures. The introduction of TDN marks a significant advancement in the field, providing a robust alternative to 3D convolutions and pushing the boundaries of video-based action recognition.
Liu et al. [93] introduced an advanced approach for recognizing human actions, particularly focusing on abnormal behaviors in video surveillance. The core of the methodology is a two-stream CNN model that synergistically combines motion history images (MHI) and RGB frames to enhance the accuracy of action recognition. The model comprises two primary components: VMHI, which employs the VGG-16 network trained on MHIs to capture motion information, and FRGB, which utilizes the Faster R-CNN algorithm trained on RGB frames with Kalman filter-assisted data annotation to capture appearance information. The integration of these two streams allows the model to effectively recognize not only individual human actions but also complex interactions between people, significantly improving the differentiation of similar actions. The model’s robustness and superior performance are demonstrated through extensive experiments on the KTH, Weizmann, UT-interaction, and TenthLab datasets, where it outperforms existing state-of-the-art methods. This innovative approach addresses critical challenges in HAR, offering significant improvements in accuracy and applicability in real-world surveillance scenarios.
In 2022, Sanchez et al. [94] presented 3DFCNN, a novel 3D fully CNN designed for real-time HAR using raw depth image sequences. This end-to-end trainable model is optimized for high accuracy and real-time performance, distinguishing itself by its ability to automatically encode spatio-temporal patterns from depth sequences while ensuring privacy protection, as identities cannot be discerned from depth data. The 3DFCNN model, characterized by its reduced architecture and computational efficiency, outperforms non-DNN-based state-of-the-art methods and competes closely with DNN-based approaches, achieving a maximum accuracy of 83.6% with a significantly lower computational cost of 1.09 seconds. Evaluated on three widely recognized datasets—NTU RGB+D Action Recognition Dataset, NorthWest-UCLA Multiview Action 3D, and UWA3D Multiview Activity II—the model demonstrates its robustness and applicability in real-world scenarios, marking a significant advancement in the field of action recognition, especially in contexts where privacy is paramount.
Anshul Shah et al. [95] presented a novel Joint-Motion Reasoning Network (JMRN) for action recognition, focusing on the dynamics of human joints. Unlike traditional methods that rely on RGB or optical flow features, JMRN emphasizes joint-based action recognition. The model operates by first extracting motion features from each joint independently through a shared motion encoder and then performing collective reasoning. A distinctive aspect of JMRN is the joint selector module, which re-weights joint information to highlight the most discriminative joints for the task at hand. Additionally, the paper introduces a joint-contrastive loss to ensure that per-joint motion signatures are expressive of the action, and a geometry-aware data augmentation technique to enhance the model’s generalization capabilities. Evaluated on several datasets including JHMDB, HMDB, Charades, AVA, and Mimetics, JMRN demonstrates substantial improvements over state-of-the-art joint-based approaches. The model’s robustness is further validated by its performance on the Mimetics dataset, which consists of out-of-context actions. The paper’s approach, focusing on the independent and collective reasoning of joint movements, marks a significant advancement in the field of action recognition, offering a new perspective and methodology that diverges from conventional RGB and optical flow-based models.
Basha et al. [96] introduced a sophisticated approach to enhancing HAR through an innovative sampling technique and deep learning integration. Specifically, the study develops a novel Gaussian Weighting Function to aggregate video frames, significantly preserving both spatial and temporal information, which is then processed using a newly proposed 3D CNN combined with LSTM networks. This architecture is adept at extracting and classifying spatio-temporal features from reduced datasets without losing critical motion details. The efficacy of this model is validated against benchmark datasets like KTH, WEIZMANN, and CASIA-B, achieving notably high accuracy rates—95.78%, 95.27%, and 95.27% respectively. These results demonstrate a considerable improvement over existing methods, underlining the potential of the proposed model in real-world applications such as surveillance and interactive systems, where efficient and accurate action recognition is crucial. This study not only addresses key challenges in human action recognition by reducing data volume and enhancing information retention but also sets a new benchmark for future research in video-based action recognition.
Senthilkumar et al. [97] introduced a sophisticated approach to HAR, emphasizing the integration of spatial-temporal models with a bi-temporal end-to-end framework. The core of this approach is the fusion of Dilated CNNs (DCNN) and deep Bi-LSTM networks, complemented by residual blocks, to effectively detect human activities from video frames. The model ingeniously combines the features extracted by CNN with functioning features using skip connections, and employs a temporal dependent relational model to preserve spatial-temporal information during data extraction. Evaluated on standard datasets like J-HMDB, UCF11, and Sports, the model demonstrates remarkable performance, achieving up to 98% accuracy in recognizing a wide range of human actions. The paper’s novelty lies in its unique integration of LSTM and CNN functions, focusing on the temporal characteristics of skeletal joints, and its ability to organize skeleton information effectively. The findings highlight the model’s superior performance and potential in real-world applications, making significant strides in the field of HAR.
In 2023, Noor et al. [98] introduced a novel, computationally efficient 3D-CNN model tailored for real-time action recognition and fall detection. Addressing the computational constraints of existing deep learning models in practical applications, this research innovates by proposing a lightweight skeleton-based 3D-CNN framework, significantly enhancing processing speed and accuracy. The model’s robustness is further fortified by an augmented human pose dataset, meticulously curated from the AIHub senior abnormal behavior dataset, focusing on fall motion and enriched with the COCO-pose dataset to improve pose estimation accuracy. This integration not only optimizes the model for real-world multi-person environments but also sets new benchmarks in fall detection, showcasing the model’s capability to efficiently process extracted pose keypoints and maintain high accuracy in action recognition. The paper’s contribution is pivotal in advancing the practical deployment of action recognition systems, promising enhanced performance in real-time applications without compromising on computational efficiency.
Zong et al. [99] presented an innovative approach to video-based action recognition, addressing the challenges of complex video backgrounds and the integration of spatial and temporal information. The authors introduce a novel four-stream network model, incorporating spatial and temporal saliency streams alongside traditional appearance and motion streams. The model, referred to as the 3M model, is further enhanced with a multi-task learning-based LSTM, adept at capturing long-term dependencies between consecutive frames and ensuring a comprehensive understanding of video dynamics. The model’s unique structure and the integration of spatial and temporal saliency streams, coupled with multi-task learning-based LSTM, significantly enhance the model’s ability to recognize and categorize actions in videos, demonstrating superior performance on popular video action recognition datasets. If this summary does not match the content of the new document, please provide additional details or attempt to access the document again.
Table 1: CNN-based Approaches.
Paper Year Model Dataset Novelty
Ji et al. [52] 2012 3D CNN KTH and TREC Video Retrieval Evaluation (TRECVID) dataset The paper introduces a unique 3D CNN architecture and proposes regularizing the 3D CNN models with auxiliary outputs computed as high-level motion features
Valle et al. [53] 2013 CNN A custom dataset The novelty lies in the CNN’s ability to process unprocessed images directly, learning to extract features during training.
Simonyan et al. [35] 2014 Two-stream CNN architecture UCF-101 and HMDB-51 The novelty lies in its two-stream architecture, which separately but concurrently processes spatial and temporal information.
Sun et al. [56] 2014 Two-layer SFA structure with 3D convolution Hollywood2, KTH, and UCF Sports The novelty lies in the combination of SFA with deep learning to learn hierarchical ’slow’ features from video data.
Ijjina et al. [57] 2014 CNN architecture with linear masks UCF50 dataset The novelty lies in its use of action bank features in combination with a CNN classifier
Cheron et al. [58] 2015 CNN JHMDB and MPII Cooking Activities The introduction of a P-CNN descriptor that uniquely combines motion and appearance information along tracks of human body parts for enhanced HAR.
Geng et al. [59] 2015 CNN KTH The combination of CNNs with a CAE for pre-training, which addresses the computational inefficiency of kernel training in deep learning models for HAR.
Du et al. [60] 2015 CNN Berkeley MHAD and ChaLearn gesture recognition datasets The novelty lies in its unique approach of transforming skeleton sequences into image representations for efficient and accurate action recognition using a specialized CNN.
Gkioxari et al. [63] 2015 CNN PASCAL VOC Actions, MPII Human Pose and Berkeley Attributes of People Datasets The novelty lies in its unique integration of secondary contextual cues with primary action regions in images using an adapted RCNN model for enhanced action recognition.
Bilen et al. [64] 2016 CNN UCF101 and HMDB51 The novelty lies in its introduction of ’dynamic images,’ a new video representation method that encodes the temporal evolution of frames into a single image for efficient HAR using CNNs
Zhang et al. [65] 2016 CNN UCF101, HMDB51, and THUMOS14 The novelty lies in its use of motion vectors instead of optical flow in a two-stream CNN architecture, enabling real-time action recognition with comparable accuracy to state-of-the-art methods
Rahmani et al. [66] 2016 3D CNN Synthetically generated dataset The novelty lies in its use of a deep CNN trained on synthetically generated depth images for view-invariant human pose representation in 3D action recognition
Xin et al. [67] 2016 Hybrid CNN-RNN model UCF101 The novelty lies in its introduction of the ARCH Networks, a hybrid DNN that innovatively combines CNN and RNN elements for robust long-term HAR in complex, real-world environments
Zhao et al. [68] 2017 Two-stream gated-RNNs and 3D-CNNs NTU RGB+D Dataset The novelty lies in its integration of gated RNNs with 3D-CNNs in a two-stream fusion method for enhanced action recognition in 3D videos data
Ullah et al. [69] 2017 Deep Bidirectional LSTM with CNN UCF-101, YouTube 11 Actions, and HMDB51 The novelty lies in its integration of Deep Bidirectional LSTM with CNN features for efficient and effective HAR in video sequences
Li et al. [70] 2017 Multi-scale deep CNN NTU RGB-D dataset The novelty lies in its translation-scale invariant image mapping method for 3D skeleton video data and the use of a multi-scale deep CNN for HAR.
Varol et al. [71] 2017 3D CNN UCF101 and HMDB51 The novelty lies in its implementation of LTC within 3D CNNs for capturing the full temporal extent of human actions in videos, significantly enhancing HAR accuracy
Arunnehru et al. [73] 2018 3D CNN KTH and Weizmann Datasets The integration of 3D motion cuboids with 3D-CNNs for action recognition, a method not extensively explored in previous studies. This approach allows for a more detailed by considering the spatial-temporal dynamics of movements
Zhigang Tu et al. [72] 2018 Human-Related Multi-Stream CNN JHMDB, HMDB51, UCF Sports and UCF101 The novelty lies in its end-to-end multi-stream CNN that uniquely combines human detection and motion saliency for enhanced HAR.
Miao Ma et al. [74] 2018 Six-stream CNN Sub-JHMDB dataset and MPII Cooking dataset The novelty lies in its unique approach of using a six-stream CNN that leverages coarse human pose tracking and multi-scale patch extraction, combined with advanced pooling layer encoding, to enhance both general and fine-grained HAR in videos
Zhang et al. [75] 2018 Two-stream-based CNN UCF101, HMDB51 and THUMOS14 The novelty lies in its use of motion vectors instead of optical flows for real-time HAR
Liu et al. [76] 2018 A novel temporal-residual 3D CNN UCF101 and HMDB51 The novelty lies in its unique integration of a residual 3D CNN with a temporal encoding method, enabling efficient and real-time HAR by capturing both short-term and long-term temporal dynamics in videos
Jiang et al. [77] 2019 Spatio-Temporal and Motion (STM) CNN-based network Something-Something v1 & v2, Jester, Kinetics-400, UCF-101 and HMDB-51 The novelty lies in its integration of spatiotemporal and motion features within a unified CNN framework, avoiding the computational intensity of 3D convolutions and optical flow calculations
Crasto et al. [78] 2019 3D CNN Kinetics and HMDB51 The novelty lies in introducing MARS, a single-stream approach that integrates motion information into RGB frames without the need for explicit optical flow computation.
Hussein et al.[79] 2019 Integrated 2D CNN and 3D CNN architecture Charades, Breakfast Actions, and MultiTHUMOS The novelty lies in introducing Timeception CNN layers, which leverage multi-scale temporal convolutions to efficiently model minute-long temporal patterns in complex actions.
He et al. [80] 2019 Integrated 2D CNN and 3D CNN architecture Kinetics and UCF101 The novelty lies in the StNet architecture, which uniquely combines super-images with 2D convolutions and a Temporal Xception Block for efficient and effective local and global spatial-temporal modelling in HAR.
Hao Yang et al. [81] 2019 Asymmetric 3D CNN UCF-101 and HMDB-51 The paper introduces efficient asymmetric one-directional 3D convolutions and MicroNets, significantly reducing computational costs and parameters while enhancing feature learning in HAR
Sudhakaran et al. [82] 2020 2D CNN Something Something-V1, Diving48 and EPIC-Kitchens The paper introduces the GSM, a lightweight module that transforms 2D-CNNs into efficient spatio-temporal feature extractors for video HAR with minimal computational overhead
Liu x et al. [83] 2020 2D CNN KTH and Weizmann datasets The paper introduced the adaptation of the VGG-16 deep learning model to kinematic human behavior recognition from video data, demonstrating enhanced accuracy compared to traditional dynamic recognition methods.
Xiong et al. [84] 2020 Two-stream 2D CNN structure Engine block assembly The paper introduces a novel transferable two-stream CNN integrating optical flow and transfer learning for robust and accurate HAR in smart manufacturing environments
Sahoo et al. [85] 2020 2D CNN with DBiLSTM KTH, UCF Sports, JHMDB, UCF101, and HMDB51 The paper introduced the integration of DHI with a CNN and a DBiLSTM network, enhancing human action recognition by capturing both depth-informed shape details and temporal dynamics
Abdelbakya et al. [86] 2020 PCA KTH dataset The novelty lies in its application of PCANet (a form of Convolutional Neural Network) for HAR, presenting a less computationally intensive alternative to traditional CNNs.
Jaouedi et al. [87] 2020 CNN KTH the paper introduces a hybrid deep learning model for human action recognition that incorporates GRNN as a core component. While the primary focus is on the integration of GRNN, Gaussian Mixture Models (GMM), and Kalman Filters (KF) for motion detection and tracking, the model is rooted in deep learning principles similar to those underlying CNNs.
Asghari et al. [88] 2020 2D CNN HMDB, JHMDB, UCF-101, and AVA The paper introduces DynaMotion, a dynamic motion representation that captures the temporal evolution of body joint heat-maps, significantly enhancing HAR in videos
Wang et al. [89] 2021 Incorporated 2D-CNN and 3D-CNN model Something-Something V2, Jester, and EgoGesture The paper introduces ACTION-Net, a novel network embedding the ACTION module that enhances 2D CNNs with multipath excitation, capturing spatial-temporal, channel-wise, and motion information efficiently for superior HAR
Lee et al. [91] 2021 3D CNN Moving-MNIST for synthetic tests, and KTH Action and Human 3.6M datasets The paper introduced the integrating LMC-Memory with 3D CNN and a memory alignment learning technique, enabling the model to precisely recall complex long-term motion contexts from short video sequences for enhanced video prediction.
Wu et al. [90] 2021 Multimodal two-stream 3D-CNN network SBU Kinect Interaction, UTD-MHAD, NTU RGB+D 60 & 120 The novelty lies in its pioneering extension of 3D CNNs to depth and pose data modalities for video HAR, introducing a multimodal two-stream 3D
Wang et al. [92] 2021 CNN model Something-Something V1 & V2 and Kinetics-400 The novelty lies in the introduction of the TDN, which efficiently captures multi-scale temporal information for action recognition by leveraging a novel TDM at both local and global motion levels
Liu et al. [93] 2021 Two-stream 2D-CNN network KTH, Weizmann, UT-interaction and TenthLab The novelty lies in its integration of motion history images with RGB frames in a two-stream CNN, significantly enhancing the accuracy of human abnormal behavior recognition in video surveillance
Sanchez et al. [94] 2022 3D-CNN NTU RGB+D, NorthWest-UCLA and UWA3D The novelty lies in the introduction of 3DFCNN, a real-time, privacy-preserving 3D fully CNN for HAR using raw depth data
Anshul Shah et al. [95] 2022 Not mentioned JHMDB, HMDB, Charades, AVA and Mimetics datasets The novelty lies in the introduction of JMRN architecture, which uniquely combines independent per-joint motion feature extraction with collective reasoning and a joint-contrastive loss, enhancing the accuracy and robustness of joint-based HAR
Basha et al. [96] 2022 3D CNN with LSTM KTH, WEIZMANN, and CASIA-B Human Activity and Gait datasets The paper introduced the use of a Gaussian Weighting Function for video frame sampling combined with a 3D CNN-LSTM architecture, significantly enhancing the preservation and classification of spatio-temporal features in human action recognition.
Senthilkumar et al. [97] 2022 Dilated CNN with deep Bi-LSTM J-HMDB and UCF-11 The novelty lies in its integration of Dilated CNNs with deep Bi-LSTM networks and a temporal dependent relational model, offering a robust spatial-temporal framework for accurate and efficient HAR
Noor et al. [98] 2023 3D-CNN COCO-pose, AIHub senior abnormal behavior and its Augmented Human Pose dataset The novelty is the introduction of a lightweight skeleton-based 3D-CNN and an augmented human pose dataset, improving real-time HAR and fall detection accuracy while reducing computational requirements
Zong et al. [99] 2023 A combination of 2D CNNs and LSTMs architeture UCF101 Action, HMDB51 Action and Large-scale Kinetics Action Datasets The novelty lies in its introduction of a spatial and temporal saliency-based four-stream network with multi-task learning, enhancing HAR by capturing salient object and motion and modelling long-term dependencies
3.2 Transformer-based Approaches in HAR
The ViT was conceptualized and introduced in the academic domain through the paper authored by Dosovitskiy et al. [6] in 2020. The ViT marked a paradigm shift in still image recognition methodologies, applying the Transformer model, predominantly known for its success in NLP, to the realm of computer vision. The application of ViTs in action recognition, a more specific and complex task within the field of computer vision, followed the initial introduction of ViT. Specifically, in 2021 and beyond, subsequent research and publications have explored and expanded the use of ViTs for action recognition tasks, demonstrating their efficacy in capturing spatial-temporal features within video data. They employ attention mechanisms to minimize redundant information and to model interactions over long distances in both space and time [100]. The adaptation of ViT to action recognition signifies the model’s versatility and its potential for broader applications in computer vision beyond static image analysis.
In 2021, Arnab et al. [30] introduced a novel approach to video classification by leveraging the capabilities of Transformer models, traditionally successful in image classification. ViViT, a Pure-transformer based model, ingeniously processes spatio-temporal tokens extracted from videos through a series of Transformer layers, addressing the challenge of managing the inherently long sequences of tokens characteristic of video data. The model’s novelty lies in its efficient variants that factorize the spatial and temporal dimensions of the input, significantly enhancing computational efficiency and scalability. Moreover, ViViT pioneers in regularizing the model during training and ingeniously utilizes pretrained image models, enabling effective training on relatively smaller datasets—a notable advancement given the typically large datasets required by Transformer models. The paper’s comprehensive ablation studies and methodical analysis culminate in ViViT achieving state-of-the-art results on several prestigious video classification benchmarks, including Kinetics 400 and 600, Epic Kitchens, Something-Something v2, and Moments in Time, thereby setting new standards and outperforming existing methods based on deep 3D convolutional networks.
Wang et al. [101] revolutionized the field of video action recognition by transitioning from conventional unimodal methodologies to a multimodal learning framework that capitalizes on the semantic nuances of label texts. This innovative approach, embodied in ActionCLIP, redefines the task as a video-text matching challenge, thereby facilitating zero-shot action recognition capabilities without necessitating additional labeled data or parameter modifications. Central to ActionCLIP is the ”pre-train, prompt, and fine-tune” paradigm, a strategy that effectively leverages extensive web image-text or video-text data for pre-training. This paradigm ingeniously aligns the action recognition task with pre-training challenges through prompt engineering and ensures robust performance through end-to-end fine-tuning on specific datasets. ActionCLIP, with its ViT-B/16 backbone, not only demonstrates exceptional zero-shot/few-shot transfer abilities but also achieves groundbreaking performance, notably securing a top-1 accuracy of 83.8% on the Kinetics-400 dataset. This paradigm shift underscores the potential of integrating rich semantic information from label texts into video representation, marking a significant leap forward in the domain of action recognition.
Bertasius et al. [29] delves into the TimeSformer model, a novel approach to video understanding that leverages the Transformer architecture, renowned for its success in NLP, and adapts it for video data. The TimeSformer stands out by employing a divided space-time attention mechanism, allowing the model to separately and efficiently process the spatial and temporal dimensions of videos. This approach is meticulously detailed in the supplementary materials, which cover the model’s implementation, training, and inference procedures, alongside additional ablations and qualitative results. The TimeSformer is rigorously evaluated on large-scale video datasets such as Kinetics-400, Kinetics-600, Something-Something-V2, and Diving-48, demonstrating its robust performance and the model’s ability to focus on relevant parts of the video for nuanced spatiotemporal reasoning. The visualizations of space-time attention in the paper further illustrate the model’s capacity to understand and interpret complex video content, marking a significant advancement in the field of video understanding.
Sharir et al. [102] introduced STAM (Space Time Attention Model), a groundbreaking approach to action recognition in videos. STAM addresses the computational inefficiencies prevalent in traditional methods that rely on 3D convolutions by significantly reducing the number of frames required for inference without compromising accuracy. The model capitalizes on a temporal Transformer that applies global attention over video frames, effectively capturing salient information and enabling input efficiency. This novel approach allows STAM to achieve state-of-the-art results on the Kinetics dataset, notably achieving 80.5% top-1 accuracy on Kinetics-400 with approximately 30% fewer frames per video and 40% faster inference compared to existing leading methods. The paper’s innovation lies in its unique application of self-attention layers for both spatial and temporal dimensions, marking a significant advancement in video action recognition by offering a solution that is not only accurate but also computationally efficient and practical for real-world applications.
Fan et al. [103] introduced Multiscale ViTs (MViT), a novel Transformer architecture designed for video and image recognition tasks, effectively integrating the concept of multiscale feature hierarchies with Transformer models. MViT operates through several channel-resolution scale stages, progressively expanding channel capacity while reducing spatial resolution, thereby creating a multiscale pyramid of features. This design allows early layers to model simple low-level visual information at high spatial resolution, while deeper layers focus on spatially coarse but complex high-dimensional features, effectively addressing the dense nature of visual signals. The model demonstrates a significant performance gain over concurrent ViTs in video recognition tasks without relying on large-scale external pre-training, and it’s particularly efficient in computation and parameters. For instance, on the Kinetics-400 dataset, MViT achieves 80.5% top-1 accuracy with approximately 30% fewer frames per video and 40% faster inference than the current leading method. The approach’s novelty lies in its ability to leverage the Transformer architecture for dense visual signals, offering an efficient and scalable solution for video understanding tasks. The paper’s findings suggest that MViT can serve as a robust framework for various visual recognition tasks, setting new benchmarks in performance and efficiency.
In 2022, a significant shift in video recognition modeling is presented by Ze liu et al. [104] , moving from conventional CNNs to Transformer-based architectures. the research presents the Video Swin Transformer, a novel adaptation of the Swin Transformer architecture tailored for video recognition tasks, leveraging the inherent spatiotemporal locality of videos. Unlike traditional Transformer models that compute self-attention globally, the Video Swin Transformer employs local self-attention within non-overlapping windows across both spatial and temporal dimensions, significantly reducing computational overhead and model size. This approach not only capitalizes on the local nature of video data but also maintains the hierarchical structure of the original Swin Transformer, extending local attention to the spatiotemporal domain. The model demonstrates exceptional performance on various video recognition benchmarks, including Kinetics-400, Kinetics-600, and Something-Something v2, achieving state-of-the-art accuracy with a smaller model size and less pre-training data compared to its predecessors. The Video Swin Transformer’s innovative design, focusing on local spatiotemporal relationships and efficient computation, marks a substantial advancement in the field of video understanding, offering a robust and scalable solution for video-based recognition tasks.
Jiawei Chen et al. [105] introduced a pioneering Transformer-based approach, MM-ViT, for video action recognition, distinctively operating within the compressed video domain. This approach capitalizes on the rich, readily available multi-modal data in videos, including I-frames, motion vectors, residuals, and audio waveforms, diverging from traditional methods that primarily depend on decoded RGB frames. MM-ViT’s architecture is meticulously designed to manage the extensive spatiotemporal tokens from these modalities, employing scalable model variants and innovative cross-modal attention mechanisms to factorize self-attention across space, time, and modality dimensions efficiently. The model’s efficacy is rigorously validated on prominent action recognition benchmarks like UCF-101, Something-Something-v2, and Kinetics-600, where it demonstrates superior performance. Notably, MM-ViT not only surpasses existing video Transformers in efficiency and accuracy but also competes favorably with state-of-the-art CNN models that rely on computationally intensive optical flow, marking a significant advancement in the domain of video action recognition.
Hashiguchi et al. [106] introduced a novel approach to enhance action recognition in videos through the innovative Multi-head Self/Cross-Attention (MSCA) mechanism, integrated within the ViT framework. This research addresses the limitations of previous methods like Temporal Shift Module (TSM) and TokenShift by fully leveraging the multi-head self-attention structure of ViT for temporal feature interaction. MSCA innovatively allows attention heads to focus not only on the current frame but also on adjacent frames, enriching the temporal context of the action recognition process. Evaluated on the Kinetics400 dataset, the MSCA mechanism, particularly the MSCA-KV variant, demonstrates superior performance, outperforming the baseline ViT and TokenShift models. This improvement underscores the model’s ability to effectively capture the dynamic temporal relationships in videos, marking a significant advancement in the field of action recognition. The paper’s contribution lies in its novel approach to temporal interaction within the Transformer block, achieving enhanced recognition performance without increasing computational complexity or altering the underlying model architecture.
Momal Ijaz et al. [107] introduced an innovative multimodal Transformer-based network aimed at improving elderly patient care in healthcare facilities. This network adeptly fuses features from skeletal joints and acceleration data to recognize nursing activities, addressing the critical need for automated and accurate monitoring of nurse care. The model’s distinctiveness lies in its dual modality approach, combining acceleration and skeletal joints data, a method not previously explored in the Nurse Care Activity Recognition Challenge (NCRC) dataset. The proposed network not only outperforms traditional single modality networks but also sets a new benchmark by surpassing state-of-the-art methods like ST-GCNs, GRUs, and classifiers based on handcrafted features, achieving an impressive accuracy of 81.8%. The research underscores the potential of Transformer-based models in healthcare, particularly in creating a fair and efficient care environment by ensuring consistent patient care and alleviating the documentation burden on nursing staff.
In 2023, Wensel et al. [108] introduced a groundbreaking ViT-ReT framework, integrating ViT and Recurrent Transformer (ReT) to address the challenges in HAR in videos. This innovative approach aims to overcome the limitations of traditional activity recognition systems that heavily rely on CNNs for feature extraction and RNNs for analyzing time-dependent data. The ViT-ReT framework replaces CNNs with ViT for efficient and salient feature extraction from images and supplants RNNs with ReT to effectively handle sequences of data, addressing the high model complexity and lack of parallelizability inherent in traditional models. The framework is designed to be lightweight and efficient, making it particularly suitable for resource-constrained and real-time environments like IoT and edge devices. Extensive comparative experiments across four publicly available human action datasets demonstrate that ViT-ReT not only achieves a significant speedup over the baseline ResNet50-LSTM approach but also outperforms state-of-the-art HAR methods in both accuracy and runtime, marking a significant advancement in the field of computer vision and activity recognition.
Chen et al. [109] presented a novel approach to HAR from still images, a notably challenging task in computer vision. The Swin-Fusion model, at the core of this research, leverages the Swin-Transformer’s hierarchical structure and introduces a feature fusion module to effectively extract and integrate multi-scale behavioral information. This innovative model addresses the inherent limitations of CNNs, such as the lack of long-range perception and difficulty in understanding global structures, by utilizing the Transformer’s ability to focus on discriminative features and calculate relationships between distant pixels. Evaluated on five distinct still image-based HAR datasets, the Swin-Fusion model demonstrates superior performance, outperforming advanced CNN-based models. It achieves this by enhancing behavior recognition through the strategic sharing and reuse of feature maps at various stages, without significantly increasing the computational resources required for training. The research marks a significant advancement in still image-based HAR, offering a robust solution that combines the strengths of Swin-Transformer architecture and feature fusion techniques.
Xing et al. [110] introduces a novel approach to semi-supervised action recognition by leveraging Transformer models, a departure from the conventional use of CNNs in this domain. Addressing the challenge of efficiently utilizing the abundance of unlabeled video data for action recognition, the authors propose SVFormer, a model that incorporates a steady pseudo-labeling mechanism (EMA-Teacher) and introduces innovative data augmentation strategies tailored for video data: Tube TokenMix and Temporal Warping Augmentation. These augmentations are designed to capture the complex temporal dynamics inherent in video content, a critical aspect that has been challenging for existing semi-supervised learning (SSL) methods. Evaluated on three benchmark datasets—Kinetics-400, UCF-101, and HMDB-51—SVFormer demonstrates a significant improvement in action recognition performance, outperforming state-of-the-art methods by a notable margin, especially under conditions of limited labeled data. This breakthrough underscores the potential of Transformers in video understanding tasks and sets a new precedent for future research in semi-supervised learning for video action recognition, highlighting the effectiveness of SVFormer’s methodological innovations in leveraging unlabeled video data.
Mondal et al. [111] introduced an approach to actor-agnostic, multi-modal, multi-label action recognition that significantly diverges from traditional actor-specific models. By ingeniously integrating visual and textual modalities through a Transformer-based detection framework inspired by DETR, and leveraging the capabilities of a pretrained vision-language model like CLIP, MSQNet is designed to recognize a multitude of actions within a single video across different actors, including both humans and animals. This integration not only facilitates a richer representation of action classes but also obviates the necessity for actor-specific pose estimation, thereby streamlining the model’s complexity and reducing its maintenance demands. The novelty of MSQNet lies in its ability to address the concurrent occurrence of multiple actions and utilize additional information sources such as class name text, which have been largely overlooked in existing methods. Demonstrated through extensive experimentation across five publicly available benchmarks, MSQNet’s superior performance, achieving up to a 50% improvement over prior arts, underscores its potential to set a new standard in the field of action recognition by offering a versatile, efficient, and more generalized solution.
Table 2: Transformer-based Approaches.
Paper Year Model Dataset Novelty
Arnab et al. [30] 2021 Pure-Transformer architecture Kinetics 400 and 600, Epic Kitchens 100, Something-Something v2 and Moments in Time The novelty lies in its introduction of a Transformer based model for video classification, efficiently handling spatio-temporal tokens through factorized spatial and temporal dimensions, and achieving state-of-the-art results on multiple benchmarks
Wang et al. [101] 2021 A ViT-based “ActionClip” architecture Kinetics-400 The novelty lies in its multimodal learning framework that transforms video HAR into a video-text matching problem, enabling robust zero-shot recognition and introducing the ”pre-train, prompt, and fine-tune” paradigm for leveraging large-scale web data
Bertasius et al. [29] 2021 ViT-based “TimeSformer” architecture Kinetics-400, Kinetics-600, Something-Something-V2 and Diving-48 The novelty lies in its innovative application of divided space-time attention, enabling efficient and effective video understanding by separately processing spatial and temporal components within the Transformer architecture
Sharir et al. [102] 2021 Transformer architecture Kinetics-400, Kinetics-600, Something-Something-V2 and Diving-48 The novelty lies in introducing STAM, a Transformer-based model for video HAR that leverages global attention over video frames, significantly reducing the required number of frames and computational load while maintaining or surpassing state-of-the-art accuracy
Fan et al. [103] 2021 ViT-based “MViT” architecture Kinetics-400, Kinetics-600, Something-Something-V2, Charadesand AVA The novelty lies in its integration of multiscale feature hierarchies with Transformer models, creating a scalable and efficient architecture for advanced video and image recognition tasks
architecture Ze liu et al. [104] 2022 ViT-based architecture “Video Swin Transformer” Kinetics-400, Kinetics-600 and Something-Something v. The novelty lies in its hierarchical, local self-attention mechanism within non-overlapping windows across both spatial and temporal dimensions
Jiawei Chen et al. [105] 2022 Transformer-based “MM-ViT” model UCF-101, Something-Something-v2 and Kinetics-600 The novelty lies in its unique approach of operating exclusively in the compressed video domain, leveraging a pure Transformer-based model to exploit all available modalities (I-frames, motion vectors, residuals, and audio) for enhanced video action recognition
Hashiguchi et al. [106] 2022 ViT-based “MSCA” architecture Kinetics400 The novelty is a mechanism within the ViT, enabling enhanced temporal interaction across frames for HAR without increasing computational complexity
Momal Ijaz et al. [107] 2022 Pure Transformer-based architecture Nurse Care Activity Recognition Challenge (NCRC) dataset The novelty of the study lies in its development of a multimodal Transformer network that uniquely fuses skeletal joints and acceleration data for enhanced nursing activity recognition
Wensel et al. [108] 2023 Integrated ViT and Recurrent Transformer (ReT) - The novelty lies in the ViT-ReT framework, which innovatively combines ViT and ReT to enhance speed and accuracy in HAR in videos, particularly in resource-constrained and real-time environments
Chen et al. [109] 2023 ViT-based architecture Li’s action, Stanford-40, PPMI-24, AUC-V1, and AUC-V2 datasets The novelty lies in the Swin-Fusion model, which innovatively combines the Swin-Transformer with a feature fusion module for enhanced HAR from still images, surpassing the performance of traditional CNN-based models
Xing et al. [110] 2023 ViT architecture Kinetics-400, UCF-101, and HMDB-51 The novelty lies in its pioneering use of Transformers for semi-supervised video recognition, enhanced by data augmentation strategies specifically designed for video data: Tube TokenMix, Temporal Warping Augmentation
3.3 CNN-Transformer Hybrid Architectures
In 2021, Neimark et al. [112] presented a transformative approach to video recognition by introducing a Transformer-based framework that diverges from the traditional reliance on 3D ConvNets. Video Transformer Networks (VTN) capitalizes on the strengths of Transformers to process entire video sequences, employing a modular architecture that includes a 2D spatial backbone for feature extraction, a temporal attention-based encoder leveraging Longformer for integrating temporal dynamics, and a classification MLP head for action prediction. This innovative framework demonstrates remarkable efficiency, achieving training speeds 16.1 times faster and inference speeds 5.1 times faster than existing state-of-the-art methods, while requiring significantly fewer GFLOPs. Evaluated on the Kinetics-400 and Moments in Time benchmarks, VTN not only maintains competitive accuracy but also offers a scalable solution for analyzing long video sequences in an end-to-end manner. The paper’s findings suggest that VTN could serve as a new baseline for video recognition research, highlighting the potential of Transformer models in efficiently handling complex video data and paving the way for future advancements in the field.
In 2022, Liu et al. [113] presented TadTR, an innovative end-to-end framework for Temporal Action Detection (TAD) utilizing Transformer technology, addressing the complexity and rigidity of previous TAD methods that rely on multi-stage pipelines and hand-designed components. TadTR simplifies the TAD process by introducing action queries, learnable embeddings that adaptively extract relevant temporal context from videos for direct action instance prediction. This approach is enhanced with three key innovations: a temporal deformable attention module for focused context extraction, a segment refinement mechanism for accurate boundary detection, and an actionness regression head for confident instance prediction. Demonstrating its effectiveness, TadTR achieves state-of-the-art performance on benchmark datasets like THUMOS14 and HACS Segments, and competitive results on ActivityNet-1.3 when combined with an additional action classifier. This streamlined, efficient framework marks a significant advancement in video understanding, offering a flexible, powerful solution for TAD with lower computational costs and improved end-to-end learning capabilities.
In 2023, Dasom Ahn et al. [114] presented an innovative approach to HAR by integrating spatio-temporal video and skeleton data into a cohesive model, addressing the complexity and inefficiency of handling cross-modal data with separate models. This method innovatively aggregates video frames and skeleton sequences into global grid tokens and joint map tokens, respectively, which are then processed by the STAR-Transformer. The core of the STAR-Transformer includes a unique encoder-decoder structure with full spatio-temporal attention, zigzag spatio-temporal attention, and binary spatio-temporal attention modules, designed to enhance the model’s ability to capture and represent multi-feature spatio-temporal data effectively. Evaluated on benchmark datasets such as Penn-Action, NTU-RGB+D 60, and NTU-RGB+D 120, the STAR-Transformer demonstrates superior performance over existing state-of-the-art methods, showcasing its potential to significantly advance the field of action recognition. This approach not only simplifies the pipeline for processing cross-modal data but also offers a more efficient and flexible solution for action recognition tasks, highlighting the effectiveness of integrating video and skeleton features through a Transformer-based model.
Surek et al. [115] introduces an innovative approach by combining a residual network (ResNet) and a ViT architecture with a semi-supervised learning method, using DINO (self-DIstillation with NO labels). This novel methodology is tested using the Human Motion Database (HMDB51), chosen for its diverse range of human actions. The study’s results are particularly promising, with the bi-dimensional ViT combined with long short-term memory (LSTM) demonstrating high accuracy rates. These findings indicate a substantial improvement in the field of video-based HAR, showcasing the effectiveness of integrating advanced deep learning techniques. The research’s approach to HAR, utilizing a blend of supervised and semi-supervised learning models, is a noteworthy contribution to the domain.
Wang et al. [116] focused on developing an efficient model capable of operating on edge devices with limited computing power. The core innovation is the 3D-ShuffleViT network, which ingeniously integrates the self-attention mechanism with convolution, incorporating the ACISA module to enhance performance. This model was rigorously evaluated on various datasets, including EgoGesture, Jester, and UCF-101, covering a broad spectrum of action recognition scenarios. The 3D-ShuffleViT network stands out for its computational efficiency, achieving a significant reduction in computational cost (only 6% of that of SlowFast-ResNet101) while maintaining high accuracy (98% of SlowFast-ResNet101’s Top1 accuracy on the EgoGesture dataset). Additionally, it offers a remarkable speed advantage, being 2.5 times faster than SlowFast-ResNet101 on the same CPU. The 3D-ShuffleViT network’s ability to balance accuracy and computational efficiency makes it an ideal solution for deployment on edge devices. Its performance on various datasets demonstrates its versatility and effectiveness in different action recognition contexts. The model’s innovative design, combining the strengths of self-attention and convolution, represents a significant step forward in the development of lightweight yet powerful video action recognition systems.
Djenouri et al. [117] introduced the CVTN (Convolution Visual Transformer Network), a novel hybrid model that synergizes CNNs and Visual Transformers (VTs) to enhance the analysis and comprehension of human activities. The model is designed to leverage sensor data in both spatial and temporal dimensions, with the CNN component learning spatial visual features from diverse sensor data, followed by the VT segment capturing temporal insights. This integration aims to provide a more robust and comprehensive understanding of human activities. The efficacy of the CVTN model is evaluated using the Kinetics dataset, which simulates real-world HAR scenarios. The results of this evaluation are significant, with the CVTN model demonstrating a remarkable performance, achieving a recognition rate of 95%, thereby outperforming recent baseline HAR solutions. The CVTN model’s approach to combining spatial and temporal data through a hybrid of CNNs and VTs represents a significant advancement in HAR.
Ye et al. [118] introduced a transformative approach to video prediction tasks such as future frame prediction and frame interpolation using a single, unified model. This model employs a transformer-based neural process combined with implicit neural representations (INRs), enabling continuous and task-agnostic prediction across various datasets including KTH, BAIR, KITTI, and Cityscapes. Distinctively, the model operates beyond traditional discrete prediction frameworks by leveraging INRs to handle spatio-temporal coordinates, which allows it to predict video frames at any desired frame rate, a significant advancement over existing task-specific models. The findings highlight the model’s capability to generate high-quality, temporally consistent videos across different scenarios, achieving state-of-the-art performance on standard metrics and demonstrating substantial improvements in video prediction technology. This research suggests a scalable and efficient pathway for future video processing applications, where a single model can adapt to multiple video prediction tasks without the need for task-specific adjustments.
Ahmadabadi et al. [119] delves into the fusion of CNNs and ViT models to advance the field of HAR. This innovative research introduces a methodology where knowledge distillation is employed to transfer insights from a comprehensive, high-capacity CNN model (teacher) to a more compact ViT model (student), aiming to amalgamate the teacher’s adeptness at extracting local image features with the student’s capability to apprehend global features through attention mechanisms. Utilizing the Stanford 40 dataset for empirical validation, the study showcases a significant enhancement in the performance of the student models, evidenced by improved accuracy and mean Average Precision (mAP) metrics compared to conventional training approaches. The novelty lies in its strategic approach to integrating the distinct advantages of CNNs and ViTs, leveraging the depth and breadth of feature recognition capabilities across different models, thereby setting a new precedent for efficiency and effectiveness in the recognition of human actions from still images. This approach not only underscores the potential of hybrid model architectures in overcoming the limitations of individual models but also highlights the efficacy of knowledge distillation as a pivotal technique for optimizing model performance in complex vision tasks.
Table 3: Hybrid Approaches.
Paper Year Model Datase Novelty
Neimark et al. [112] 2021 “VTN” a hybrid CNN-transformer Kinetics-400 and Moments in Time The novelty of the lies in its Transformer-based framework for video recognition that processes entire video sequences efficiently, moving beyond traditional 3D ConvNets and enabling faster training and inference with competitive accuracy
Liu et al. [113] 2022 “TadTR” a hybrid CNN-transformer THUMOS14, HACS Segments and ActivityNet-1.3 TadTR introduces a novel end-to-end Transformer-based framework for Temporal Action Detection, simplifying the detection pipeline with action queries and incorporating unique enhancements like temporal deformable attention, segment refinement, and actionness regression for efficient and accurate action instance prediction
Dasom Ahn et al. [114] 2023 “STAR-Transformer” a hybrid CNN-Transformer model Penn-Action, NTU-RGB+D 60, and NTU-RGB+D 120 datasets The novelty lies in the unique integration of spatio-temporal video and skeleton data into a cohesive Transformer-based model, employing specialized attention mechanisms to efficiently process cross-modal data for HAR
Surek et al. [115] 2023 A hybrid CNN-Transformer model HMDB51 The novelty of the study lies in its integration of ResNet and ViT with DINO (self-DIstillation with NO labels) for semi-supervised learning in HAR from RGB videos
Wang et al. [116] 2023 “3D-ShuffleViT” a hybrid CNN-transformer architecture EgoGesture, Jester and UCF-101 The novelty lies in the deep integration of self-attention and convolution within a 3D framework, achieving high video HAR accuracy with significantly reduced computational costs for deployment on edge devices
Djenouri et al. [117] 2023 A hybrid CNN-transformer architecture Kinetics dataset The novelty of the study lies in the CVTN model’s unique hybrid integration of CNNs and ViTs to enhance HAR by leveraging both spatial and temporal sensor data
Ye et al. [118] 2023 A hybrid CNN-transformer architecture KTH, BAIR, KITTI and Cityscapes The novelty of the paper lies in its development of a unified, transformer-based neural process model that enables continuous conditional video prediction across multiple tasks using implicit neural representations.
Ahmadabadi et al. [119] 2023 A hybrid CNN-transformer architecture Stanford 40 dataset The novelty lies in enhancing HAR by distilling knowledge from high-capacity CNNs to compact ViTs, merging local and global image feature extraction for improved accuracy and efficiency
3.4 Discussion
In the field of action recognition, the choice of model – whether CNN-based, ViT-based, or a hybrid of CNN and ViT – significantly influences the outcome and efficiency of the task. CNN-based models are particularly adept at extracting local features due to their convolutional nature [120], making them highly effective in pattern recognition within images and videos. Their computational efficiency is a boon for real-time applications [121], and their robustness to input variations is notable [122]. However, CNNs often struggle with global contextual understanding [36] and are prone to overfitting. Moreover, their ability to model long-range temporal dependencies [123], which is crucial in action recognition, is somewhat limited.
ViT-based models, in contrast, excel in capturing global dependencies [124, 6], thanks to their self-attention mechanism. This attribute makes them particularly suited for understanding complex actions that require a broader view beyond local features. ViTs are scalable with data, benefiting significantly from larger datasets, and are flexible in processing inputs of various sizes [125]. The adaptability in processing various input sizes is a byproduct of the patch-based approach and the global receptive field of the ViTs. However, these models are computationally more intensive and require substantial training data to achieve optimal performance [11]. Unlike CNNs, ViTs are not as efficient in extracting detailed local features, which can be a critical drawback in certain action recognition scenarios.
Hybrid models that combine CNNs and ViTs aim to harness the strengths of both architectures. They offer the local feature extraction capabilities of CNNs along with the global context awareness of ViTs, potentially providing a more balanced approach to action recognition. These models can be more efficient and versatile, adapting well to a range of tasks. However, this combination brings its own challenges, including increased architectural complexity, higher resource demands, and the need for careful tuning to balance the contributions of both CNN and ViT components. The choice among these models depends on the specific requirements of the action recognition task, such as the available computational resources, the nature and size of the dataset, and the types of actions that need to be recognized. For a summary of the advantages and disadvantages of these three architectural variations, see Table 4.
Table 4: Capability Comparison of Transformer-based, CNN-based, and Hybrid Models in Action Recognition
Criteria ViT-based CNN-based Hybrid Models
Advantages
Excel at capturing global dependencies
Scalable with data
Flexible in processing various input sizes
Adept at extracting local features
Computationally efficient
Robust to input variations
Efficient and versatile
Adapts well to a range of tasks
Disadvantages
Computationally intensive
Requires substantial training data
Limited global contextual understanding
Prone to overfitting
Limited in modeling long-range dependencies
Architectural complexity
Higher resource demands
Need for careful tuning
Balancing contributions of both components can be challenging
4 Proposed CNN-ViT Hybrid Architecture
In this section, we will detail our proposed CNN-ViT architecture for human action recognition, which leverages the benefits of both approaches described above. The architecture incorporates a TimeDistributed layer with a CNN backbone, followed by a ViT model to classify actions in video sequences.
Spatial Component (TimeDistributed CNN): The CNN backbone (i.e. Mobilenet [121]) in TimeDistributed layer processes each frame $\boldsymbol{X}\in\mathbb{R}^{K\times M\times N}$, and outputs spatial features vector $\boldsymbol{v}=p_{\theta}(\boldsymbol{X})\in\mathbb{R}^{L}$ where $p_{\theta}$ is the CNN model (e.g. MobileNet or VGG16) with parameters $\theta$ wrapped by TimeDistributed layer.
Let $\mathcal{X}$ be a collection of $n$ frames $\mathcal{X}=\{\boldsymbol{X}^{i}\}_{i=1}^{n}$, which is fed into the Spatial component of our architecture in Figure 7. The trained CNN $p_{\theta}$ is then used $n$ times to process $\mathcal{X}$ frame by frame and create a set of $n$ spatial feature vectors $\mathcal{V}=\{\boldsymbol{v}^{i}\}_{i=1}^{n}$, where $\boldsymbol{v}^{i}=p_{\theta}(\boldsymbol{X}^{i})$.
Temporal Component (ViT): In this proposed hybrid CNN-ViT model, the Vit is engineered to process sequences of $N$ vectors, see Figure 7, each representing a distinct frame of the video, but in the proposed ViT-only model, each vector is representing a distinct patch, see Figure 8. These vectors are first linearly projected into a high-dimensional space, facilitating the model’s ability to learn complex patterns within the data. To ensure the model captures the sequential nature of the input, positional encodings are added to these embeddings. The core of the ViT consists of two layers, each comprising a multi-head self-attention mechanism and a feed-forward network. The self-attention mechanism allows the model to weigh the importance of different patches relative to each other, while the feed-forward network, utilizing an Exponential Linear Unit (ELU) activation function, processes each position independently to capture global context. The ViT is designed to aggregate the information from all vectors and positional encodings into a single [CLS] token, which is prepended to the input sequence. The output vector associated with this [CLS] token, after propagation through the Transformer layers, serves as a comprehensive representation of the entire input, suitable for downstream classification tasks.
The sequence of vectors $\mathcal{V}$ which is outputted by the TimeDistributed CNN is then fed into the ViT. The ViT processes this sequence and outputs a final representation $\boldsymbol{z}$ which is represented as
$\boldsymbol{z}=\text{ViT}([\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{N}]).$ (16)
The final representation $\boldsymbol{z}$ is then fed into the Softmax layer to classify the action in the video. The Transformer encoder is designed to process a sequence of vectors, each representing one frame, and aggregate information into a single vector for classification.
[…]
Figure 7: The hybrid CNN-ViT architecture.
[…]
Figure 8: The ViT-only architecture.
5 Experiments
The goal of the presented experiments is not necessarily to produce a model that outperforms the state-of-the-art models in HAR field. Rather, the aim is to conduct a comparison among the ViT-only, CNN, and hybrid models to give further insights. The KTH dataset was chosen for its balanced representation of spatial and temporal features. Six experiments were conducted, training each of the aforementioned models on three different lengths of frame sequences. Care was taken to avoid pretraining in order to ensure the neutrality of the results. The Transnet model by Alomar et al. [126], was used to represent the CNN model. The ViT model was depicted in Figure 8. For the Spatial Component of the hybrid model, we employed the spatial component of Transnet and for the Temporal Component, we employed the same ViT model that we used in ViT-only model.
Experiment Settings. We constructed our model utilizing Python 3.6, incorporating the Keras deep learning framework, OpenCV for image processing, matplotlib, and the scikit-learn library. The training and testing were performed on a computer equipped with an Intel Core i7 processor, an NVidia RTX 2070 graphics card, and 64GB of RAM.
Data. In 2004, the Royal Institute of Technology unveiled the KTH dataset, a significant and publicly accessible dataset for action recognition [127]. Renowned as a benchmark dataset, it encompasses six types of actions: walking, jogging, running, boxing, hand-waving, and hand-clapping. The dataset features performances by twenty-five different individuals, introducing a diversity in execution. Additionally, the environment for each participant’s actions was deliberately altered, including settings such as outdoors, outdoors with scale changes, outdoors with clothing variations, and indoors. The KTH dataset comprises 2,391 video sequences, all recorded at 25 frames per second using a stationary camera against uniform backgrounds.
5.1 Results and Discussion
In the exploration of HAR within video sequences, three distinct models were evaluated i.e., CNN, ViT-only, and a Hybrid model. These models were trained on the KTH dataset, focusing on three different context lengths i.e., short (12 frames), medium (18 frames) and long (24 frames). The results from these experiments provide insightful revelations into the efficacy of each model under different temporal contexts, see Table 5.
Table 5: Experimental Results of three models on the KTH Dataset using three different context lengths. Hybrid-N refers to the model trained without pre-training whereas Hybrid-I is for the model pretrained on Imagenet[34].
Context length CNN-based ViT-only Hybrid-N Hybrid-I
12 frames 94.35 92.44 94.12 96.34
18 frames 93.91 92.82 94.56 97.13
24 frames 93.49 93.69 95.78 97.89
The CNN model exhibited a decrease in accuracy as the frame length increased, recording 94.35% for 12 frames, 93.91% for 18 frames, and 93.49% for 24 frames. This descending trend suggests that CNN may struggle with processing longer sequences where temporal dynamics become more complex, potentially leading to challenges such as overfitting or difficulties in temporal feature retention over extended durations.
In contrast, the ViT model demonstrated an improvement in performance with longer sequences, achieving accuracies of 92.44% for 12 frames, 92.82% for 18 frames, and 93.69% for 24 frames. This ascending pattern supports the notion that ViT architectures, with their inherent self-attention mechanisms, are well-suited to managing longer sequences. The ability of ViTs to assign varying degrees of importance to different parts of the sequence likely contributes to their enhanced performance on longer input frames.
The Hybrid CNN-ViT model showcased the highest and continuously improving accuracy rates across all frame lengths: 94.12% for 12 frames, 94.56% for 18 frames, and an impressive 95.78% for 24 frames. This model synergistically combines CNN’s robust spatial feature extraction capabilities with ViT’s efficient handling of temporal relationships via self-attention. The results from this model indicate that such a hybrid approach is particularly effective in capturing the complexities of action recognition tasks in video sequences, especially as the sequence length increases.
These findings underscore the potential advantages of hybrid neural network architectures in video-based action recognition tasks, particularly for handling longer sequences with complex interactions. The superior performance of the Hybrid CNN-ViT model suggests that integrating the spatial acuity of CNNs with the temporal finesse of ViTs can lead to more accurate and reliable recognition systems. Future work could explore the scalability of these models to other datasets, their computational efficiency, and their robustness against variations in video quality and scene dynamics. Additionally, further research might investigate the optimal balance of CNN and ViT components within hybrid models to maximize both performance and efficiency.
Table 6: Comparison of proposed hybrid model with the state-of-the-art models on KTH dataset.
Methods Venue Accuracy
Geng et al. [59] ICCSAE ’15 92.49
Arunnehru et al. [73] RoSMa ’18 94.90
Abdelbakya et al. [86] ITCE ’20 87.52
Jaouedi et al. [87] KSUCI journal ’20 96.30
Liu X et al [83] JAIHC ’20 91.93
Sahoo et al. [85] TETCI ’20 97.67
Lee et al. [91] CVF ’21 89.40
Basha et al. [96] MTA journal ’22 96.53
Ye et al. [118] CVF ’23 90.90
Ours - 97.89
Table 6 Shows that the presented CNN-ViT hybrid model achieved an impressive 97.89% accuracy on the KTH dataset, placing it prominently among state-of-the-art models for HAR. This performance is notably superior when compared to earlier benchmarks reported in the literature such as Geng et al. [59], who achieved 92.49%, and Arunnehru et al.[73] with 94.90%. Our model utilizes an ImageNet-pre-trained MobileNet [121] as the CNN backbone in the spatial component, which enhances its robust feature extraction capabilities. Combined with the dynamic attention mechanisms of ViT, it thereby enhances both the spatial and temporal processing of video sequences. Furthermore, the model not only surpasses other contemporary approaches like Liu X et al [83] (91.93%) and Lee et al. [91] (89.40%) but also shows competitive or superior performance against some of the highest accuracies in the field, such as Jaouedi et al. [87] (96.30%) and Basha et al. [96] (96.53%). Even in comparison to the high benchmark set by Sahoo et al. [85] (97.67%), our model demonstrates a marginal but significant improvement, underscoring the efficacy of integrating CNN with ViT. This integration not only facilitates more nuanced feature extraction across both spatial and sequential dimensions but also adapts more dynamically to the varied contexts inherent in video data, making it a potent solution for realistic action recognition scenarios.
The integration of CNN with ViT is particularly advantageous for enhancing feature extraction capabilities and focusing on relevant segments dynamically through the attention mechanisms of ViTs. This not only helps in improving accuracy but also in making the model more adaptable to varied video contexts, a key requirement for action recognition in realistic scenarios. This comparative advantage suggests that hybrid models are paving the way for future explorations in human action recognition, combining the best of convolutional and ViT-based architectures for improved performance and efficiency.
6 Challenges and Future Directions
Challenges The field of HAR faces several formidable challenges that stem from the inherent complexity of interpreting human movements within diverse and dynamic environments. One of the primary obstacles is the variability in human actions themselves, which can differ significantly in speed, scale, and execution from one individual to another [2]. This variability necessitates the development of sophisticated models capable of generalizing across a wide range of actions without sacrificing accuracy [128]. Additionally, the presence of complex backgrounds and environments further complicates the task of HAR. Systems must be adept at isolating and recognizing human actions against a backdrop of potentially distracting or obstructive elements, which can vary from the bustling activity of a city street to the unpredictable conditions of outdoor settings [129, 37].
Moreover, HAR systems must navigate the fine line between inter-class similarity and intra-class variability, where actions that are similar to each other (such as running versus jogging) require nuanced differentiation, while the same action can appear markedly different when performed by different individuals or under varying circumstances [130, 131]. The challenge of temporal segmentation adds another layer of complexity, as accurately determining the start and end of an action within a continuous video stream is crucial for effective recognition [132]. Coupled with the need for computational efficiency to process video data in real-time and the difficulties associated with obtaining large, accurately annotated datasets, these challenges underscore the multifaceted nature of HAR [133]. Addressing these issues is critical for advancing the field and enhancing the practical applicability of HAR systems in real-world applications, from surveillance and security to healthcare and entertainment.
Future Directions The motivation behind this work has been driven by the compelling need to bridge the existing gap between the spatial feature extraction capabilities inherent in CNNs and the dynamic temporal processing strengths found in ViTs [30]. Through the introduction of a novel hybrid model, an attempt has been made to leverage the synergistic potential of these technologies, thereby enhancing the accuracy and efficiency of HAR systems in capturing the complex spatial-temporal dynamics of human actions.
Looking forward, a promising future for HAR is envisioned, particularly through the development of hybrid and integrated models. It is believed that the potential of these models extends beyond immediate performance improvements, inspiring new directions for research within the field. It is anticipated that future studies will focus on optimizing these hybrid architectures, aiming to make them more scalable and adaptable to real-world applications across various domains such as surveillance, healthcare, and interactive media. Furthermore, the exploration of self-attention mechanisms and the adaptation of large-scale pre-training strategies from ViTs are seen as exciting prospects for HAR. These approaches are expected to lead to the development of more sophisticated models capable of understanding and interpreting human actions with unprecedented accuracy and nuance.
The integration of CNNs and ViTs into Hybrid CNN-ViT models presents a promising avenue for overcoming the challenges faced by HAR systems. These hybrid models capitalize on the strengths of both architectures: the local feature extraction capabilities of CNNs and the global context understanding of ViTs. Future developments could focus on enhancing model adaptability to generalize across diverse actions, improving the isolation of human actions from complex backgrounds through advanced attention mechanisms, and developing nuanced differentiation techniques for closely related actions [124]. Innovations in model architecture, alongside the application of transfer learning and few-shot learning techniques, could significantly reduce the variability challenge in human actions.
Moreover, addressing the temporal segmentation challenge requires the integration of specialized temporal modules and sequence-to-sequence models to accurately determine the start and end of an action within continuous video streams. Computational efficiency remains paramount for real-time processing, necessitating ongoing efforts in model optimization and the exploration of synthetic data generation to mitigate the difficulties associated with obtaining large, accurately annotated datasets. Customizable Hybrid CNN-ViT models that can be tailored for specific applications, from surveillance to healthcare, will ensure that these advancements not only push the boundaries of academic research but also enhance practical applicability in real-world scenarios. Through these concerted efforts, Hybrid CNN-ViT models are poised to make significant contributions to the field of HAR, offering innovative solutions to its multifaceted challenges.
This work has highlighted the importance of continued innovation and cross-disciplinary collaboration in the advancement of HAR technologies. By integrating insights from computer vision, machine learning, and domain-specific knowledge, it is hoped that HAR systems will not only become more efficient and accurate but also more responsive to the complexities and variances of human behavior in natural environments. As the field moves forward, the focus is set on pushing the boundaries of what is possible in HAR, with the aim of creating systems that enhance human-computer interaction and contribute positively to society through various applications.
7 Conclusions
This survey provides a comprehensive overview of the current state of HAR by examining the roles and advancements of RNNs, CNNs, and ViTs. It delves into the evolution of these architectures, emphasizing their individual contributions to the field. The introduction of a hybrid model that combines the spatial processing capabilities of CNNs with the temporal understanding of ViTs represents a methodological advancement in HAR. This model aims to address the limitations of each architecture when used in isolation, proposing a unified approach that potentially enhances the accuracy and efficiency of action recognition tasks. The paper identifies key challenges and opportunities within HAR, such as the need for models that can effectively integrate spatial and temporal information from video data. The exploration of hybrid models, as suggested, offers a pathway for future research, particularly in improving model performance on complex video datasets. The discussion encourages further investigation into optimizing these hybrid architectures and exploring their applicability across various domains. This work sets a foundation for future studies to build upon, aiming to push the boundaries of what is currently achievable in HAR and to explore new applications of these technologies in real-world scenarios.

### v2

@@
-RNNs, CNNs and Transformers in Human Action Recognition: A Survey and A Hybrid Model
+CNNs, RNNs and Transformers in Human Action Recognition: A Survey and a Hybrid Model
 Abstract
 Human action recognition (HAR) encompasses the task of monitoring human activities across various domains, including but not limited to medical, educational, entertainment, visual surveillance, video retrieval, and the identification of anomalous activities.
-Over the past decade, the field of HAR has witnessed substantial progress by leveraging Convolutional Neural Networks (CNNs) to effectively extract and comprehend intricate information, thereby enhancing the overall performance of HAR systems.
+Over the past decade, the field of HAR has witnessed substantial progress by leveraging convolutional neural networks (CNNs) and recurrent neural networks (RNNs) to effectively extract and comprehend intricate information, thereby enhancing the overall performance of HAR systems.
 Recently, the domain of computer vision has witnessed the emergence of Vision Transformers (ViTs) as a potent solution.
@@
 Notably, within this landscape, the research community has shown keen interest in HAR, acknowledging its manifold utility and widespread adoption across various domains.
-This article aims to present an encompassing survey that focuses on CNNs and the evolution of Recurrent Neural Networks (RNNs) to ViTs given their importance in the domain of HAR.
+This article aims to present an encompassing survey that focuses on CNNs and the evolution of RNNs to ViTs given their importance in the domain of HAR.
 By conducting a thorough examination of existing literature and exploring emerging trends, this study undertakes a critical analysis and synthesis of the accumulated knowledge in this field.
@@
 keywords:
-Human action recognition; convolutional neural networks; recurrent neural networks; vision transformers; deep learning; video classification
+Human action recognition · Convolutional neural networks · Recurrent neural networks · Vision Transformers · Deep learning · Video classification
 1 Introduction
@@
 However, it is worth noting that a predominant focus of these scholarly works has been placed on the examination and evaluation of convolutional neural networks (CNNs) and traditional machine learning models within the realm of action recognition.
-The advent of Transformer[5] architecture has sparked a paradigm shift in deep learning, leading researchers to swiftly adopt and adapt these architectures to enhance the accuracy and efficiency of various computer vision tasks including action recognition.
-Consequently, there is a compelling need for a recent review that comprehensively examines the state-of-the-art research including Vision Transformers (ViTs) [6], a special Transformer for computer vision tasks, and Hybrid models in addition to CNNs.
-Such a review would serve as a crucial guiding resource to shape the future research directions with Transformer and CNN-Transformer hybrid architectures beside CNNs which previously were seen as unique and influential models for action recognition.
+The advent of Transformer architecture [96] has sparked a paradigm shift in deep learning.
 By employing a multi-head self-attention layer, the Transformer model computes sequence representations by effectively aligning words within the sequence with other words in the same sequence [93].
@@
 The incorporation of multi-head attention allows the Transformer model to collectively learn a range of representations from diverse perspectives through the collaboration of multiple attention layers.
-Inspired by Transformers, several Natural Language Processing (NLP) tasks have achieved remarkable performance, reaching human-level capabilities, as exemplified by models such as GPT [7] and BERT [8].
-The remarkable achievements of Transformers in handling sequential data, particularly in the domain of NLP, have prompted the exploration and advancement of ViTs [6].
-These novel models have demonstrated comparable or even superior performance compared to CNNs in the context of image recognition tasks, especially when operating on vast datasets such as Imagenet [9, 10, 11].
+Inspired by Transformers, many natural language processing (NLP) tasks have achieved remarkable performance, reaching human-level capabilities, as exemplified by models such as GPT [12] and BERT [24].
+The remarkable achievements of Transformers in handling sequential data, particularly in the domain of NLP, have prompted the exploration and advancement of Vision Transformer (ViT) [29] (a special Transformer for computer vision tasks).
+ViTs have demonstrated comparable or even superior performance compared to CNNs in the context of image recognition tasks, especially when operating on vast datasets such as ImageNet [41, 66, 54].
 This observation signifies a noteworthy shift in the field, wherein ViTs possess the potential to supplant the established dominance of CNNs in computer vision, mirroring the displacement witnessed in the case of recurrent neural networks (RNNs) [93].
@@
 Unlike image recognition, video recognition focuses on the complex challenge of identifying and understanding events within video sequences, including the recognition of human actions.
+Consequently, there is a compelling need for a recent review that comprehensively examines the state-of-the-art research including ViTs and hybrid models in addition to CNNs and RNNs for HAR.
+Such a review would serve as a crucial guiding resource to shape the future research directions with Transformer and CNN-Transformer hybrid architectures beside CNNs which previously were seen as unique and influential models for HAR.
 The main contributions of this paper is as follows.
-  • We present a thorough review of the CNNs, RNNs, and ViTs. This review examines the evolution from traditional methods to the latest advancements in neural network architectures.
-  • We present an extensive examination of existing literature related to action recognition.
+  • We present a thorough review of the CNNs, RNNs and ViTs. This review examines the evolution from traditional methods to the latest advancements in neural network architectures.
+  • We present an extensive examination of existing literature related to HAR.
   • We propose a novel hybrid model integrating the strengths of CNNs and ViTs.
-In addition, we provide a detailed performance comparison the proposed hybrid model against existing models.
+In addition, we provide a detailed performance comparison of the proposed hybrid model against existing models.
 The analysis highlights the model’s efficacy in handling complex HAR tasks with improved accuracy and efficiency.
-  • We also discuss emerging trends and the future direction of HAR technologies, emphasizing the importance of hybrid models in enhancing the interpretability and robustness of action recognition systems.
+  • We also discuss emerging trends and the future direction of HAR technologies, emphasizing the importance of hybrid models in enhancing the interpretability and robustness of HAR systems.
 These contributions enrich the understanding of the current state and future prospects of HAR, proposing innovative approaches and highlighting the importance of integrating different neural network architectures to advance the field.
-The paper is structured as follows: Section 2 delves into the background, covering foundational concepts and technologies crucial to HAR, including CNNs, RNNs, and ViTs.
-Section 3 reviews related work, highlighting the chronological evolution of HAR deep learning technologies.
-The novel hybrid model combining CNNs and ViTs is introduced in Section 4, followed by Section 5, which details the experimental setup and evaluation of the model.
-Section 6 discusses the challenges and their implications for future directions.
+The paper is structured as follows.
+Section 2 delves into the background, covering foundational concepts and technologies crucial to HAR, including CNNs, RNNs and ViTs, highlighting the chronological evolution of HAR deep learning technologies.
+Section 3 thoroughly reviews related HAR works with a brief discussion.
+A novel hybrid model combining CNNs and ViTs is proposed in Section 4, including the details of the experimental setup and the results.
+Section 5 discusses the challenges and their implications for future directions in HAR.
 Finally, Section 6 concludes the paper.
 2 Background
+This section provides a chronological and technical overview of three fundamental types of neural networks: CNNs, RNNs, and Transformers.
+CNNs, introduced in the late 1980s, revolutionized image processing by leveraging local connectivity and shared weights to efficiently detect spatial hierarchies in data.
+As the field progressed, RNNs emerged in the 1990s, addressing the need for modeling sequential data through their ability to maintain temporal dependencies across sequences.
+The advent of Transformers in 2017 marked a paradigm shift by utilizing self-attention mechanisms to capture global relationships in data more effectively, thereby enhancing performance in a wide array of tasks beyond sequential data.
+This background section will delve into the technical intricacies and evolutionary trajectory of these architectures, highlighting their contributions and transitions in the realm of deep learning.
+2.1 CNNs
+The evolution of CNNs has been remarkable since their introduction in the 1980s.
+Originally, CNNs were designed to process static images, primarily focusing on spatial recognition tasks such as object and pattern recognition.
+The initial idea was to build layers of convolutional filters that would apply various operations to the image to extract features like edges, textures, and shapes.
+This structure proved highly effective for tasks like image classification, object detection, image segmentation and more in computer vision.
+The Neocognitron [34], developed by Kunihiko Fukushima, presented an early example of neural networks incorporating convolutional operations for image processing, setting the foundations for subsequent progress.
+Shortly after, Yann LeCun and collaborators introduced LeNet-5 [61], a key architecture designed for handwritten digit recognition, showcasing the effectiveness of convolutional layers in pattern recognition tasks.
+The progress of CNNs reached a turning point in the mid-2010s with the introduction of models like AlexNet [58], showcasing their potential in image classification tasks.
+Alongside architectural innovations, this milestone was achieved thanks to access to large datasets, notably, ImageNet [23], and computational improvements, including the rise of graphics processing units (GPUs) for parallel computing.
+Large-scale datasets provided the diversity and complexity necessary for training deep networks, while enhanced computational power accelerated the training of sophisticated CNN architectures.
+The architectural enhancements, large datasets, and increased computational capabilities helped CNNs to be a cornerstone in deep learning methodologies, extending their applications beyond image processing to various domains.
+Notable architectures like VGGNet [81], distinguished by its uniform design and small convolutional filters, GoogLeNet [88], with its inception modules for capturing features at different scales efficiently, and ResNet [43], which introduced residual learning for training very deep networks, have further enriched the landscape of CNNs.
+2.1.1 Spatio-Temporal CNNs
+As CNNs excelled in spatial tasks, researchers began exploring their potential in handling temporal data, such as video and time-series analysis.
+The challenge was to incorporate the dimension of time into the inherently spatial architecture of CNNs.
+To address this task, spatio-temporal CNNs were developed.
+These networks extend traditional CNN architectures by adding a temporal component to analyze dynamic behaviors across time frames.
+Several approaches have been utilized and main types are as follows.
+3D convolution involves extending the 2D kernels to 3D, allowing the network to perform convolution across both spatial and temporal dimensions.
+This approach is directly applied to video data where the third dimension represents time [42, 90].
+The two-stream CNNs involve running two parallel CNN streams: one for spatial processing of individual frames and another for temporal processing, usually of optical flow, which captures motion between frames [81, 33].
+RNNs with CNNs aim to combine CNNs for spatial processing with RNNs like long short-term memory (LSTM) or gated recurrent unit (GRU) to handle temporal dependencies.
+This hybrid model leverages CNNs’ ability to extract spatial features and RNNs’ capacity to manage temporal sequences effectively [111, 28].
 2.2 From Vanilla RNN to Attention-Based Transformers
-The fundamental aim of the Transformers lies in effectively addressing time series data tasks, placing a distinct emphasis on NLP as one of its primary objectives.
-To address these tasks, RNNs have emerged as a prominent deep learning technique.
-However, the development of the Transformers did not occur in isolation; rather, it was the result of a series of iterative optimisations that were progressively integrated into the existing RNN structures.
-This section aims to elucidate the underlying RNN techniques and the subsequent advancements that were incorporated into them, ultimately leading to the final formulation of the Transformer architecture.
-Notation.
-Here, we establish common notations for RNN architectures including Vanilla RNNs, Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) to streamline discussions in subsequent sections.
+This section explores the evolution from RNNs to the Transformers, highlighting the progression in handling time series and sequence data.
+Initially, RNNs were the go-to deep learning technique for managing temporal tasks, effectively capturing sequential dependencies.
+However, the development of Transformers marked a significant leap forward, driven by a series of iterative improvements and optimizations that built upon the limitations of RNNs.
+Transformers, with their focus on NLP, introduced a novel attention mechanism that allows for more efficient and scalable processing of sequential data.
+By examining the foundational RNN techniques and the subsequent enhancements leading to the Transformer architecture, this section elucidates the transformative journey from traditional RNN models to the sophisticated attention-based frameworks that now dominate the field.
+We firstly establish common notations for RNN architectures including vanilla RNNs, LSTM and GRU to streamline discussions in subsequent sections.
 In these architectures, each iteration involves a cell that sequentially processes an input embedding $\boldsymbol{x}_{t}\in\mathbb{R}^{n_{x}}$ and retains information from the previous sequence through the hidden state $\boldsymbol{h}_{t-1}\in\mathbb{R}^{n_{h}}$ using weight matrices $\boldsymbol{W}\in\mathbb{R}^{n_{h}\times n_{h}}$ and $\boldsymbol{U}\in\mathbb{R}^{n_{h}\times n_{x}}$.
-The $\boldsymbol{W}$-like matrices encompasses all weights related to hidden-to-hidden connections, while $\boldsymbol{U}$-like matrices encompasses all weight matrices related to input-to-hidden connections.
+The $\boldsymbol{W}$-like matrices encompass all weights related to hidden-to-hidden connections, while $\boldsymbol{U}$-like matrices encompass all weight matrices related to input-to-hidden connections.
 Additionally, bias terms are represented by $\boldsymbol{b}$-like vectors.
 Each cell produces a new hidden state $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ as its output.
+2.2.1 Vanilla RNNs
 Vanilla RNNs [77, 50] lack the presence of a cell state, relying solely on the hidden states as the primary means of memory retention within the RNN framework.
@@
 Vanilla RNNs effectively incorporate short-term dependencies of temporal order and past inputs in a meaningful manner.
-However, they are characterised by certain limitations.
-Firstly, due to their intrinsic sequential nature, RNNs pose challenges in parallelizing computations [14].
+However, they are characterized by certain limitations.
+Firstly, due to their intrinsic sequential nature, RNNs pose challenges in parallelized computations [40].
 Consequently, this limitation can impose restrictions on the overall speed and scalability of the network.
@@
 […]
-(a) Vanilla
+(a) Vanilla RNN
 […]
@@
 Figure 2: Various types of RNN cells.
-LSTM Hochreiter and Schmidhuber (1997) introduced the Long Short-Term Memory (LSTM) cell as a solution to address the issue of long-term dependencies and to mitigate the challenge of interdependencies among successive steps [16].
+2.2.2 LSTM
+Hochreiter and Schmidhuber (1997) introduced the LSTM cell as a solution to address the issue of long-term dependencies and to mitigate the challenge of interdependencies among successive steps [44].
 LSTM architecture incorporates a distinct component known as the cell state $\boldsymbol{c}_{t}\in\mathbb{R}^{n_{h}}$, illustrated in Figure 1(b).
 Analogous to a freeway, this cell state facilitates the smooth flow of information, ensuring that it can readily traverse without undergoing significant alterations.
-In 2000, Gers, Schmidhuber, and Cummins [17] made modifications to the initial LSTM architecture by incorporating a forget gate within the cell structure.
+[36] made modifications to the initial LSTM architecture by incorporating a forget gate within the cell structure.
 The mathematical expressions describing this modified LSTM cell are derived from its inner connections.
-Hence, the LSTM cell can be formally represented based on the depicted interconnections as follows:
+Hence, the LSTM cell can be formally represented based on the depicted interconnections as follows.
   • Forget gate decides what information should be thrown away or kept from the cell state with the equation
     $\boldsymbol{f}_{t}=\sigma(\boldsymbol{W}_{f}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{f}\boldsymbol{x}_{t}+\boldsymbol{b}_{f}),$ (2)
-    where $\boldsymbol{f}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the forget gate and $\sigma$ is sigmoid activation function.
+    where $\boldsymbol{f}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the forget gate and $\sigma$ is the sigmoid activation function.
   • Input gate determines which new information is added to the cell state with two activation functions defined as
     $\boldsymbol{i}_{t}=\sigma(\boldsymbol{W}_{i}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{i}\boldsymbol{x}_{t}+\boldsymbol{b}_{i}),$ (3)
-    where $\boldsymbol{i}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the sigmoid activation function and
-    $\boldsymbol{\tilde{c}}_{t}=\tanh(\boldsymbol{W}_{\tilde{c}}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{\tilde{c}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{c}})$ (4)
+    where $\boldsymbol{i}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the sigmoid activation function; and
+    $\boldsymbol{\tilde{c}}_{t}=\tanh(\boldsymbol{W}_{\tilde{c}}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{\tilde{c}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{c}}),$ (4)
     where $\boldsymbol{\tilde{c}}_{t}\in\mathbb{R}^{n_{h}}$ is known as candidate value. After obtaining $\boldsymbol{i}_{t}$ and $\boldsymbol{\tilde{c}}_{t}$, we can update the cell state with
-    $\boldsymbol{c}_{t}=\boldsymbol{f}_{t}\odot\boldsymbol{c}_{t-1}+\boldsymbol{i}_{t}\odot\boldsymbol{\tilde{c}}_{t}$ (5)
-    where $\boldsymbol{c}_{t-1}\in\mathbb{R}^{n_{h}}$ is the previous cell state.
+    $\boldsymbol{c}_{t}=\boldsymbol{f}_{t}\odot\boldsymbol{c}_{t-1}+\boldsymbol{i}_{t}\odot\boldsymbol{\tilde{c}}_{t},$ (5)
+    where $\boldsymbol{c}_{t-1}\in\mathbb{R}^{n_{h}}$ is the previous cell state and $\odot$ is the Hadamard operator.
   • Output gate determines the next hidden state based on the cell state and output gate’s activity
-    $\boldsymbol{o_{t}}=\sigma(\boldsymbol{W}_{o}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{o}\boldsymbol{x}_{t}+\boldsymbol{b}_{o})$ (6)
-    where $\boldsymbol{o}_{t}\in\mathbb{R}^{n_{h}}$ is the output of output gate. Finally the updated hidden state,
+    $\boldsymbol{o_{t}}=\sigma(\boldsymbol{W}_{o}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{o}\boldsymbol{x}_{t}+\boldsymbol{b}_{o}),$ (6)
+    where $\boldsymbol{o}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the output gate. Finally the updated hidden state,
     $\boldsymbol{h}_{t}=\tanh(\boldsymbol{c}_{t})\odot\boldsymbol{o}_{t}$ (7)
@@
 However, this advantage comes at the expense of increased complexity.
-GRU The LSTM cell surpasses the learning capability of the conventional recurrent cell, yet the additional number of parameters escalates the computational load.
-Consequently, to address this concern, Chung et al.
-[18] introduced the gated recurrent unit (GRU), see Figure 2.
+2.2.3 GRU
+The LSTM cell surpasses the learning capability of the conventional recurrent cell, yet the additional number of parameters escalates the computational load.
+Consequently, to address this concern, [18] introduced the GRU, see Figure 2.
 GRU demonstrates comparable performance to LSTM while offering a more computationally efficient design with fewer weights.
@@
   • Update gate decides how much of the past information needs to be passed along with
-    $\boldsymbol{z}_{t}=\sigma(\boldsymbol{W}_{z}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{z}\boldsymbol{x}_{t}+\boldsymbol{b}_{z})$ (8)
-    where $\boldsymbol{z}_{t}\in\mathbb{R}^{n_{h}}$ is the output of update gate. A candidate activation for the subsequent step
+    $\boldsymbol{z}_{t}=\sigma(\boldsymbol{W}_{z}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{z}\boldsymbol{x}_{t}+\boldsymbol{b}_{z}),$ (8)
+    where $\boldsymbol{z}_{t}\in\mathbb{R}^{n_{h}}$ is the output of the update gate. The output of the reset gate $\boldsymbol{r}_{t}\in\mathbb{R}^{n_{h}}$ is obtained by
+    $\boldsymbol{r}_{t}=\sigma(\boldsymbol{W}_{r}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{r}\boldsymbol{x}_{t}+\boldsymbol{b}_{r}).$ (9)
+    A candidate activation for the subsequent step is
     $\boldsymbol{\tilde{h}}_{t}=\tanh(\boldsymbol{W}_{\tilde{h}}(\boldsymbol{r}_{t}\odot\boldsymbol{h}_{t-1})+\boldsymbol{U}_{\tilde{h}}\boldsymbol{x}_{t}+\boldsymbol{b}_{\tilde{h}})$ (10)
-    where $\boldsymbol{\tilde{h}}_{t}\in\mathbb{R}^{n_{h}}$ and
-    $\boldsymbol{r}_{t}=\sigma(\boldsymbol{W}_{r}\boldsymbol{h}_{t-1}+\boldsymbol{U}_{r}\boldsymbol{x}_{t}+\boldsymbol{b}_{r})$ (10)
-    where $\boldsymbol{r}_{t}\in\mathbb{R}^{n_{h}}$ is the output of reset gate.
-  • Final Activation is a blend of the previous hidden state and the candidate activation, weighted by the update gate.
-This mechanism allows the GRU to effectively retain or replace old information with new information.
+    where $\boldsymbol{\tilde{h}}_{t}\in\mathbb{R}^{n_{h}}$.
+  • The final activation is a blend of the previous hidden state and the candidate activation, weighted by the update gate, i.e.,
     $\boldsymbol{h}_{t}=\boldsymbol{z}_{t}\odot\boldsymbol{\tilde{h}}_{t}+(1-\boldsymbol{z}_{t})\odot\boldsymbol{h}_{t-1}$ (11)
-    where $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ is the updated hidden state.
-Types of RNNs RNNs were created with an internal memory mechanism that allows them to store and use information from previous outputs.
+    where $\boldsymbol{h}_{t}\in\mathbb{R}^{n_{h}}$ is the updated hidden state. This mechanism allows the GRU to effectively retain or replace old information with new information.
+[…]
+Figure 3: Types of RNN structures based on input-output pairs.
+2.2.4 Types of RNNs
+RNNs were created with an internal memory mechanism that allows them to store and use information from previous outputs.
 This unique trait enables RNNs to retain important contextual information over time, enabling reasoned decision-making based on past results.
@@
 It operates with fixed input and output sizes, functioning similarly to a standard neural network.
-[…]
-Figure 3: Types of RNN structures based on input-output pairs.
 One-to-many represents a specific category of RNNs that is characterized by its ability to produce multiple outputs based on a single input provided to the model.
 This type of RNN is particularly useful in applications like image captioning, where a fixed input size results in a series of data outputs.
-The Many-to-one RNNs merge a sequence of inputs into a single output through a series of hidden layers that learn relevant features.
+Many-to-one RNNs merge a sequence of inputs into a single output through a series of hidden layers that learn relevant features.
 An illustrative instance of this RNN type is sentiment analysis, where the model analyzes a sequence of text inputs and produces a single output indicating the sentiment expressed in the text.
-The Many-to-many RNNs are employed to generate a sequence of output data from a sequence of input units.
+Many-to-many RNNs are employed to generate a sequence of output data from a sequence of input units.
 It can be categorized into two subcategories: equal size and unequal size.
-In the Equal Size subcategory, the input and output layers have the same size, see Many-to-many architecture in Figure 3.
-Several research efforts have emerged to tackle the limitation of fixed-size input-output sequences in machine translation tasks, as they fail to adequately represent real-world requirements.
+In the equal size subcategory, the input and output layers have the same size, see many-to-many architecture in Figure 3c.
+Several research efforts have emerged to tackle the limitation of the fixed-size input-output sequences in machine translation tasks, as they fail to adequately represent real-world requirements.
 The unequal size subcategory can handle different sizes of inputs and outputs.
-A practical application of the the Unequal Size subcategory can be observed in Machine Translation.
+A practical application of the unequal size subcategory can be observed in machine translation.
 In this scenario, the model generates a sequence of translated text outputs based on a sequence of input sentences.
 Unequal size subcategory employs an encoder-decoder architecture, where the encoder adopts the many-to-one architecture, and the decoder adopts the one-to-many architecture.
-One notable contribution in this area was made by Kalchbrenner and Blunsom [19], who pioneered the approach of mapping the entire input sentence to a vector.
-This work is related to the study conducted by Cho et al.
-[20], although the latter was specifically utilized to refine hypotheses generated by a phrase-based system [21].
+One notable contribution in this area was made by [51], who pioneered the approach of mapping the entire input sentence to a vector.
+This work is related to the study conducted by [17], although the latter was specifically utilized to refine hypotheses generated by a phrase-based system [87].
 In this architecture, the encoder component plays a crucial role in transforming the inputs into a singular vector, commonly referred to as the context.
-This context vector, typically with a length of 256, 512, or 1024, encapsulates all the pertinent information detected by the encoder from the input sentence, which serves as the translation target, see Figure 4(a).
+This context vector, typically with a length of 256, 512 or 1024, encapsulates all the pertinent information detected by the encoder from the input sentence, which serves as the translation target, see Figure 4(a).
 Subsequently, this vector is passed on to the decoder, which generates the corresponding output sequence.
-It’s important to note that both the encoder and decoder components in this architecture are RNNs.
-2.1.1 Attention
-The evolution of attention mechanisms in neural networks represents a significant advancement in the field of deep learning, particularly in tasks related to NLP and machine translation.
-Initially introduced by Graves et al.
-[22], the concept of attention mechanisms was designed to enhance the model’s ability to focus on specific parts of the input sequence when generating an output, mimicking the human ability to concentrate on particular aspects of a task.
-This foundational work laid the groundwork for subsequent developments in attention mechanisms, providing a mechanism for models to dynamically assign importance to different parts of the input data.
-Building on Graves’ initial concept, Bahdanau et al.
-[23] introduced the additive attention mechanism, which was specifically designed to improve machine translation.
-This approach computes the attention weights through a feed-forward neural network, allowing the model to consider the entire input sequence and determine the relevance of each part when translating a segment.
-This additive form of attention significantly improved the performance of sequence-to-sequence models by enabling a more nuanced understanding and alignment between the input and output sequences [21].
-Following this, Luong et al.
-[24] proposed the multiplicative attention mechanism, also known as dot-product attention, which simplifies the computation of attention weights by calculating the dot product between the query and all keys.
-This method not only streamlined the attention mechanism but also offered improvements in computational efficiency and performance in various NLP tasks, marking a pivotal moment in the evolution of attention mechanisms from their inception to more sophisticated and efficient variants.
+It is important to note that both the encoder and decoder components in this architecture are RNNs.
+Different from Figure 4(a), Figure 4(b) gives the encoder-decoder architecture with attention which will be introduced in the next section.
 […]
@@
 (b) With attention
-Figure 4: Sequence-To-Sequence RNN with and without attention mechanism.
+Figure 4: Sequence-to-sequence RNN with and without the attention mechanism.
+2.2.5 Attention
+The evolution of attention mechanisms in neural networks represents a significant advancement in the field of deep learning, particularly in tasks related to NLP and machine translation.
+Initially introduced by [39], the concept of attention mechanisms was designed to enhance the model’s ability to focus on specific parts of the input sequence when generating an output, mimicking the human ability to concentrate on particular aspects of a task.
+This foundational work laid the groundwork for subsequent developments in attention mechanisms, providing a mechanism for models to dynamically assign importance to different parts of the input data.
+Building on Graves’ initial concept, [6] introduced the additive attention mechanism, which was specifically designed to improve machine translation.
+This approach computes the attention weights through a feed-forward neural network, allowing the model to consider the entire input sequence and determine the relevance of each part when translating a segment.
+This additive form of attention significantly improved the performance of sequence-to-sequence models by enabling a more nuanced understanding and alignment between the input and output sequences [87].
+Following this, [69] proposed the multiplicative attention mechanism, also known as dot-product attention, which simplifies the computation of attention weights by calculating the dot product between the query and all keys.
+This method not only streamlined the attention mechanism but also offered improvements in computational efficiency and performance in various NLP tasks, marking a pivotal moment in the evolution of attention mechanisms from their inception to more sophisticated and efficient variants.
 The central idea of the attention mechanism is to shift focus from the task of learning a single vector representation for each sentence.
@@
 One possible explanation for the improvement is that the attention layer created memories associated with the context pattern rather than memories associated with the input itself, relieving pressure on the RNN model structure’s weights and causing the model memory to be devoted to remembering the input rather than the context pattern [46].
+(a) Transformer
+(b) Self-Attention
+Figure 5: Transformer architecture and its self-attention mechanism (adapted from [96]).
 2.2.6 Self-Attention
 To this point, attention mechanisms in sequence-transformation models have primarily relied on complex RNNs, featuring an encoder and a decoder, the most successful models in language translation yet.
-However, in 2017, Vaswani et al.
-[5] introduced a simple network architecture known as the Transformer, see Figure 5, which exclusively utilized attention mechanism, eliminating the need for RNNs.
+However, [96] introduced a simple network architecture known as the Transformer, see Figure 5, which exclusively utilized attention mechanism, eliminating the need for RNNs.
 They introduced a novel attention mechanism called self-attention, which is also known as KQV-attention (Key, Query, and Value).
@@
 The purpose of self-attention is to compute a representation of each element in a sequence by considering the entire sequence, thereby capturing the contextual relationships between elements regardless of their positional distance from each other.
-This ability to capture both local and global dependencies makes self-attention particularly powerful for tasks such as machine translation, text summarization, and sequence prediction, where understanding the context and the relationship between words or elements in a sequence is crucial [5]
+This ability to capture both local and global dependencies makes self-attention particularly powerful for tasks such as machine translation, text summarization, and sequence prediction, where understanding the context and the relationship between words or elements in a sequence is crucial [96].
 The mathematical formulation of self-attention involves several key steps.
-First, a set of query vectors $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}^{Q}$, a set of key vectors $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}^{K}$, and a set of value vectors $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}^{V}$ are calculated through linear transformations of the input sequence, where $\boldsymbol{X}$ is the input matrix representing embeddings of tokens in a sequence, $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ are weight matrices for queries, keys, and values, respectively.
-The attention scores are then calculated by taking the dot product of the query vector with all key vectors, followed by scaling the result by the inverse square root of the dimension of the keys ($\sqrt{d_{k}}$) to avoid overly large values.
+First, a set of query vectors $\boldsymbol{Q}=\boldsymbol{X}\boldsymbol{W}^{Q}$, a set of key vectors $\boldsymbol{K}=\boldsymbol{X}\boldsymbol{W}^{K}$, and a set of value vectors $\boldsymbol{V}=\boldsymbol{X}\boldsymbol{W}^{V}$ are calculated through linear transformations of the input sequence, where $\boldsymbol{X}$ is the input matrix representing embeddings of tokens in a sequence, and $\boldsymbol{W}^{Q}$,$\boldsymbol{W}^{K}$, and $\boldsymbol{W}^{V}$ are weight matrices for queries, keys, and values, respectively.
+The attention scores are then calculated by taking the dot product of the query vector with all key vectors, followed by scaling the result by the inverse square root of the dimension of the keys (say $\sqrt{d_{k}}$) to avoid overly large values.
 These scores are then passed through a softmax function to obtain the attention weights, which represent the importance of each element’s contribution to the output.
-Finally, the output for each position is computed as a weighted sum of the value vectors,
-$\boldsymbol{A(\boldsymbol{Q,K,V})}=\text{softmax}(\frac{\boldsymbol{QK^{\top}}}{\sqrt{d_{k}}})\boldsymbol{V}$ (12)
+Finally, the output say $\boldsymbol{A}$ is computed as a weighted sum of the value vectors, i.e.,
+$\boldsymbol{A(\boldsymbol{Q,K,V})}=\text{softmax}(\frac{\boldsymbol{QK^{\top}}}{\sqrt{d_{k}}})\boldsymbol{V}.$ (12)
 This process allows the model to dynamically focus on different parts of the input sequence, enabling the extraction of rich contextual information from the sequence.
-(a) Transformer
-(b) Self-Attention
-Figure 5: Transformer architecture and its self-attention mechanism, adapted from [5].
 2.2.7 Multi-Head-Attention
@@
 The outputs of these independent attention computations are then concatenated and linearly transformed into the expected dimension.
-The mathematical formulation of multi-head attention can be described in four steps.
-First, Multi-Head Attention Computation,
-$\boldsymbol{Q}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{Q},\quad\boldsymbol{K}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{K},\quad\boldsymbol{V}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{V}$ (13)
-where i is the index of each self-attention head,
-$\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})=\text{softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V}_{i}$ (14)
-Then, concatenating outputs from all heads followed by a final linear transformation,
-$\boldsymbol{A}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{Concat}(\boldsymbol{A}_{1},\boldsymbol{A}_{2},\ldots,\boldsymbol{A}_{h})\boldsymbol{W}^{O}$ (15)
-where $\boldsymbol{W}^{O}$ is the weight matrix for the final linear transformation.
+The mathematical formulation of the multi-head attention can be described in the following steps.
+First, for the $i$-th self-attention head, find
+$\boldsymbol{Q}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{Q},\quad\boldsymbol{K}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{K},\quad\boldsymbol{V}_{i}=\boldsymbol{X}\boldsymbol{W}_{i}^{V},$ (13)
+and then compute
+$\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})=\text{softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V}_{i}.$ (14)
+The multi-head attention is obtained by concatenating all $\boldsymbol{A}_{i}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{i})$.
 The multi-head attention mechanism enables the model to capture different types of information from different positions of the input sequence.
-By processing the sequence through multiple attention ”heads”, the model can focus on different aspects of the sequence, such as syntactic and semantic features, simultaneously.
+By processing the sequence through multiple attention “heads”, the model can focus on different aspects of the sequence, such as syntactic and semantic features, simultaneously.
 This capability enhances the model’s ability to understand and represent complex data, making multi-head attention a powerful component of Transformer-based architectures [25].
 2.3 From Transformer to Vision Transformer
-The journey from the inception of the Transformer model to the development of the Vision Transformer marks a pivotal advancement in deep learning, showcasing the adaptability of models initially designed for sequence data processing to the realm of image analysis.
-This transition underscores a significant shift in approach, from conventional image processing techniques to more sophisticated, sequence-based methodologies.
-Introduced in 2017 by Vaswani et al.
-[5] through the seminal paper ”Attention Is All You Need,” the Transformer model revolutionized NLP by leveraging self-attention mechanisms.
+The journey from the inception of the Transformer model to the development of the ViT marks a pivotal advancement in deep learning, showcasing the adaptability of models initially designed for sequence data processing to the realm of image analysis.
+This transition underscores a significant shift in approach, from conventional image processing techniques to more sophisticated sequence-based methodologies.
+Introduced by [96] through the seminal paper “Attention Is All You Need”, the Transformer model revolutionized NLP by leveraging self-attention mechanisms.
 This innovation allowed for the processing of sequences of data without the reliance on recurrent layers, facilitating unprecedented parallelization and significantly reducing training times for large datasets.
 The Transformer’s success in NLP sparked curiosity about its potential applicability across different types of data, including images, setting the stage for a transformative adaptation.
-The adaptation of Transformers for image data pivoted on a novel concept: treating images not as traditional 2D arrays of pixels but as sequences of smaller, discrete image patches.
+The adaptation of Transformers for image data pivoted on a novel concept: treating images not as traditional 2D arrays of pixels but as sequences of smaller and discrete image patches.
 This approach, however, faced computational challenges due to the self-attention mechanism’s quadratic complexity with respect to input length.
-The breakthrough came with the introduction of the ViT by Dosovitskiy et al.
-[6] in 2020, which applied the Transformer architecture directly to images, see Figure 6.
+The breakthrough came with the introduction of the ViT by [29], which applied the Transformer architecture directly to images, see Figure 6.
 By dividing an image into fixed-size patches and processing these patches as if they were tokens in a text sequence, ViT was able to capture complex relationships between different parts of an image using the Transformer’s encoder.
@@
 The sequence of embeddings is then processed through the Transformer encoder, which consists of layers of multi-head self-attention and feed-forward neural networks, allowing the model to dynamically weigh the importance of each patch relative to others for a given task.
-For tasks like image classification, the output from the Transformer encoder is passed through a classification head, often utilizing a learnable ”class token” appended to the sequence of patch embeddings for this purpose.
+For tasks like image classification, the output from the Transformer encoder is passed through a classification head, often utilizing a learnable “class token” appended to the sequence of patch embeddings for this purpose.
 The model is trained on large datasets using backpropagation and, during inference, processes images through these steps to predict their classes.
@@
 Nonetheless, the development of ViT signifies a significant milestone in the application of sequence processing models to the field of computer vision, opening new avenues for research and practical applications.
-Figure 6: The Vision Transformer architecture, adapted from [6].
-2.2.1 Vision Transformer: From Spatial to Spatio-Temporal Tasks
+Figure 6: The ViT architecture (adapted from [29]).
 The original ViT, designed for static image processing, divides images into patches and interprets these as sequences, leveraging the Transformer’s self-attention mechanism to understand complex spatial relationships.
@@
 Several works attempted to adapt ViT in action recognition task using different methods as below.
-Temporal Dimension Integration The integration of the temporal dimension is a fundamental step in adapting ViT for action recognition.
+Temporal dimension integration.
+The integration of the temporal dimension is a fundamental step in adapting ViT for action recognition.
 Traditional ViT models process images as a series of patches, treating them essentially as sequences for the self-attention mechanism to analyze spatial relationships.
@@
 This allows the models to capture the evolution of actions across frames.
-The work by Bertasius et al.
-[29] highlights the potential of incorporating temporal information into Transformers, marking a significant advancement in video analysis capabilities.
-Spatio-Temporal Embeddings To effectively capture the dynamics of actions within videos, adapted ViT models generate spatio-temporal embeddings.
+The work by [10] highlights the potential of incorporating temporal information into Transformers, marking a significant advancement in video analysis capabilities.
+Spatio-temporal embeddings.
+To effectively capture the dynamics of actions within videos, adapted ViT models generate spatio-temporal embeddings.
 This involves extending the traditional positional embeddings used in ViTs to also include temporal positions, thereby creating embeddings that account for both spatial and temporal information within video sequences.
-The discussion by Arnab et al.
-[30] on the creation of these spatio-temporal embeddings showcases the method’s effectiveness in enhancing the model’s understanding of action dynamics across both space and time.
-Multi-Head Self-Attention Across Time The extension of self-attention mechanisms to analyze relationships between patches not just within individual frames but also across different frames is crucial for recognizing actions over time.
+The discussion by [4] on the creation of these spatio-temporal embeddings showcases the method’s effectiveness in enhancing the model’s understanding of action dynamics across both space and time.
+Multi-head self-attention across time.
+The extension of self-attention mechanisms to analyze relationships between patches not just within individual frames but also across different frames is crucial for recognizing actions over time.
 This approach enables the model to identify relevant features and changes across the video sequences, facilitating a deeper understanding of motion and the progression of actions.
-The exploration by Bertasius et al.
-[29] of this concept demonstrates how Transformers can be effectively adapted to capture the temporal dynamics of actions, a key aspect of video analysis.
-2.3 CNNs: From Spatial to Spatio-Temporal Tasks
-CNNs have undergone a significant evolution since their introduction in the 1980s.
-The Neocognitron [31], developed by Kunihiko Fukushima, presented an early example of neural networks incorporating convolutional operations for image processing, setting the foundations for subsequent progress.
-Shortly after, Yann LeCun and collaborators introduced LeNet-5 [32], a key architecture designed for handwritten digit recognition, showcasing the effectiveness of convolutional layers in pattern recognition tasks.
-The progress of CNNs reached a turning point in the mid-2010s with the introduction of models like AlexNet [33], showcasing their potential in image classification tasks.
-Alongside architectural innovations, this milestone was achieved thanks to access to large datasets, notably, ImageNet [34], and computational improvements, including the rise of Graphics Processing Units (GPUs) for parallel computing.
-Large-scale datasets provided the diversity and complexity necessary for training deep networks, while enhanced computational power accelerated the training of sophisticated CNN architectures.
-The architectural enhancements, large datasets, and increased computational capabilities helped CNNs to be a cornerstone in deep learning methodologies, extending their applications beyond image processing to various domains.
-Notable architectures like VGGNet [35], distinguished by its uniform design and small convolutional filters, GoogLeNet [36], with its inception modules for capturing features at different scales efficiently, and ResNet [37], which introduced residual learning for training very deep networks, have further enriched the landscape of CNNs.
-Transfer Learning.
-Deep Neural Networks, specifically CNNs, have been proven to work very well in the presence of large, diverse datasets.
-However, access to large representative datasets is a big challenge in the deep learning era.
-To benefit from the impressive pattern recognition capability of CNNs in the absence of diverse and large datasets, transfer learning comes into play to leverage knowledge gained from a source task to enhance the performance of a model on a related target task.
-Transfer learning involves pre-training a model on a large dataset through which CNNs’ early layers capture relatively general features and patterns that may be common and helpful for other related target tasks.
-By freezing the early layers to preserve the learned knowledge, and fine-tuning the later layers on a target task to allow the model to adapt to target task-specific features, the power of deep networks is exploited even when there is limited data.
-ImageNet is a widely recognised source dataset within the domain of computer vision.
-Several prominent architectural designs, including VGG, ResNet and GoogleNet have made their pre-trained weights on the Imagenet dataset publicly available to provide a source for subsequent researchers and practitioners.
-The availability of pre-trained weights accelerates the training process on diverse computer vision tasks, enabling the transfer of learned features and representations from the ImageNet dataset to target applications.
-Object detection, another fundamental task in computer vision, has seen significant advancements in the last 10 years.
-Early methods such as Viola-Jones [38] employed handcrafted features and classifiers to detect and classify objects in a given image.
-The paradigm shift in the 2010s was marked by the introduction of R-CNN [39]: Region with CNN features and subsequent architectures like Fast R-CNN [39] and Faster R-CNN [40].
-This pivotal development brought forth a two-stage approach, involving the proposal of regions of interest and their classification.
-Following that Single Shot MultiBox Detector (SSD) [41] and You Only Look Once (YOLO) [42] models that allow real-time object detection were proposed to show the efficacy of single-shot detection frameworks.
-One of the recent aims in the field is to achieve anchor-free methods [43, 44] to work with more flexible and efficient methods.
-Segmentation.
-Similar to object detection, the early techniques for segmentation, a computer vision task that aims to partition an image into distinct and meaningful regions, involved handcrafted features.
-Early methods like thresholding and region-growing struggled with complex scenes and variations.
-Advancements in CNNs in the 2010s also skyrocketed the number of proposed architectures that use powerful CNNs and state-of-the-art performance was quickly surpassed by methods like Fully Convolutional Networks (FCN) [45] and U-Net [46].
-FCN allowed pixel-level predictions while maintaining spatial information and U-Net helped with better feature integration with the introduction of skip connections.
-Skip connections, also known as shortcut connections or residual connections, were incorporated in the U-Net architecture to address the challenge of information loss during downsampling and upsampling operations.
-These connections enable the direct flow of information between corresponding encoder and decoder layers, facilitating the integration of low-level and high-level features.
-This design choice enhances the model’s ability to capture fine details and spatial information while maintaining pixel-level predictions, making it especially effective for image segmentation tasks.
-Transfer learning is not limited to image classification tasks in the computer vision field.
-Object detection and segmentation tasks, also greatly benefit from the pre-trained weights to enhance model performance and efficiency.
-In object detection tasks, architectures like R-CNN [39], YOLO [42] and Mask R-CNN [47]; in segmentation tasks architectures like FCN [45] and PSPNet [48] trained on large datasets are employed as a starting point and fine-tuning happens on target tasks.
-In essence, transfer learning in deep learning, and particularly in CNNs, facilitates the transfer of knowledge, allowing models to achieve impressive results in diverse applications, even when faced with limited labelled data for a target task.
-3 Deep Learning-Based Techniques in HAR
+The exploration by [10] of this concept demonstrates how Transformers can be effectively adapted to capture the temporal dynamics of actions, a key aspect of video analysis.
+3 Literature Review
 This section briefly recalls the most commonly used deep learning-based HAR approaches.
 3.1 CNN-Based Approaches in HAR
-This section recalls the most prominent CNN-based approaches in HAR, organized chronologically.
-Deep learning was still in its early stages in 2012, and CNNs, or CNNs, had not yet gained significant popularity in the field of HAR.
-The focus was primarily on traditional machine learning approaches, such as Support Vector Machines (SVMs) [49], and handcrafted features, such as Histogram of Oriented Gradients (HOG) [50] and Histogram of Optical Flow (HOF) [51].
+This section recalls the most prominent CNN-based approaches in HAR based on the model type (i.e., the two-stream CNN, 3D CNN, and RNNs with CNNs), organized chronologically.
+Deep learning was still in its early stages in 2012, and CNNs or RNNs had not yet gained significant popularity in the field of HAR.
+The focus was primarily on traditional machine learning approaches, such as support vector machines [19], and handcrafted features, such as histogram of oriented gradients [21] and histogram of optical flow [7].
 A few studies did, nevertheless, start looking into neural networks for action recognition.
-In 2012, Ji et al.
-[52] presented an exploration into the use of 3D CNNs for recognizing human actions in surveillance videos.
-This model innovatively extends traditional 2D CNNs by performing 3D convolutions, thereby extracting features from both spatial and temporal dimensions to effectively capture motion information.
-Evaluated on the TREC Video Retrieval Evaluation (TRECVID) dataset and the KTH dataset, the 3D CNN model demonstrated significant improvements over traditional 2D CNNs and other baseline methods in action recognition tasks.
-The paper’s novelty lies in its unique approach to integrating spatial-temporal features for action recognition, offering enhanced performance in real-world surveillance scenarios.
-For papers specifically from 2012 that only used CNN architectures for HAR, there may not be a significant number of high-impact papers due to the fact that the use of CNNs in this field was still emerging and not as established as in later years.
-The seminal works that laid the groundwork for using CNNs in action recognition often came slightly later, as the computational power and dataset availability became more conducive to deep learning approaches.
-In 2013, Valle et al.
-[53] address the challenge of HAR from video data, focusing specifically on the recognition of walking and running actions.
-They propose an innovative approach using a variant of a 2D CNN, a type of multilayer perceptron, which is distinct for its ability to learn feature extraction directly from raw, unprocessed images during the error backpropagation phase.
-This methodology represents a significant advancement over traditional methods that rely on pre-processed images.
-The authors developed a specialized dataset comprising 4000 images per activity to train and evaluate their model.
-The paper’s contribution lies in its novel approach to feature extraction in HAR, demonstrating the potential of CNNs in efficiently and accurately recognizing complex human actions directly from video sequences, a task of growing importance in fields such as surveillance and automated monitoring systems.
 In 2014, the use of CNNs in action recognition was at a pivotal stage, marking a shift from hand-crafted feature-based methods to deep learning approaches.
-The key points of the use of CNNs in action recognition at that period of time is the following.
-Emergence of Deep Learning, deep learning, particularly CNNs, had started to dominate image classification tasks, thanks to their ability to learn feature representations directly from raw pixel data.
+The key points of the use of CNNs in action recognition at that period of time are the following.
+(I) Emergence of deep learning: deep learning, particularly CNNs, had started to dominate image classification tasks, thanks to their ability to learn feature representations directly from raw pixel data.
 This success in static images paved the way for applying CNNs to video data for action recognition.
-Challenges in Video data, unlike 2D images, videos incorporate a third dimension which represents the temporal patterns, making action recognition more complex.
+(II) Challenges in video data: unlike 2D images, videos incorporate a third dimension which represents the temporal patterns, making action recognition more complex.
 CNNs had to be adapted to not only recognize spatial patterns but also capture motion information over time dimension.
-Datasets and Benchmarks, the adoption of large-scale video datasets like UCF-101 [54] and HMDB-51 [55] became more common.
+(III) Datasets and benchmarks: the adoption of large-scale video datasets like UCF-101 [83] and HMDB-51 [59] became more common.
 These datasets provided diverse sets of actions and were large enough to train deep networks.
 The performance on these benchmarks has been becoming a key measure of progress for action recognition models.
-Transfer Learning, due to the computational expense of training CNNs from scratch and the relatively smaller size of video datasets compared to image datasets, transfer learning became a popular strategy.
+(IV) Transfer learning: due to the computational expense of training CNNs from scratch and the relatively smaller size of video datasets compared to image datasets, transfer learning became a popular strategy.
 Networks pre-trained on large image datasets like ImageNet [23] were fine-tuned on video frames for action recognition tasks.
-Computational Constraints, despite the promise of CNNs, computational constraints were a significant challenge.
+(V) Computational constraints: despite the promise of CNNs, computational constraints were a significant challenge.
 Training deep networks required significant GPU power, and processing video data with CNNs was resource-intensive.
 This limited the complexity of the models that could be trained and the size of the datasets that could be used.
-In 2014, Simonyan et al.
-[35] presented an innovative approach to recognizing actions in video sequences by using a two-stream CNN architecture.
+3.1.1 Two-Stream CNNs
+[81] presented an innovative approach to recognize actions in video sequences by using a two-stream CNN architecture.
 This approach divides the task into two distinct problems: recognizing spatial features from single frames and capturing temporal features across frames.
-The spatial stream CNN processes static visual information, while the temporal stream CNN handles motion by analysing optical flow.
-The model was tested on benchmark datasets like UCF-101 and HMDB-51, where it achieved state-of-the-art results, showcasing the effectiveness of this dual-stream method.
+The spatial stream CNN processes static visual information, while the temporal stream CNN handles motion by analyzing optical flow.
+The model was tested on benchmark datasets like UCF-101 and HMDB-51, where it achieved state-of-the-art results, showcasing the effectiveness of this two-stream method.
 The novelty of this work lies in the separation of motion and appearance features, which allows for more specialized networks that can better capture the complexities of video-based action recognition.
 The success of this model has made a significant impact on the field, influencing many future research directions in video understanding.
-Sun et al.
-[56] presented a novel approach to action recognition by integrating deep learning with Slow Feature Analysis (SFA).
-SFA is a computational technique used to extract features from a signal that change slowly over time.
-It is based on the principle that while the sensory data may change rapidly, the underlying causes of the data often change at a slower rate.
-By focusing on these slowly varying features, SFA aims to capture the inherent properties of the observed data that are more stable and consistent over time.
-In the context of visual data, such as video sequences, SFA would aim to identify features that remain constant or change slowly as the video progresses.
-These could be the overall shape of an object, its identity, or its trajectory through space, as opposed to the rapidly changing texture or lighting conditions.
-The authors developed a CNN that processes video data to extract a hierarchy of complex features for action recognition tasks.
-The research stands out for applying a deep learning structure to SFA, which traditionally uses a single-layer approach.
-This deepened SFA is shown to be superior in learning more generalizable features from video sequences.
-The model was evaluated on Hollywood2, KTH, and UCF Sports datasets, which are standard benchmarks with considerable intra-class variation challenges.
-The paper reports that the deep learning-enhanced SFA model surpasses traditional methods and some state-of-the-art techniques in accuracy, without relying on additional processing tricks such as dense sampling.
-This suggests that the model’s performance is due to its robust feature learning capabilities.
-Ijjina et al.
-[57] presented a novel approach to HAR by leveraging the patterns found in action bank features through a specially designed CNN.
-Action bank features are computed by considering a fixed set of videos known as an action bank.
-These videos serve as templates to measure the similarity of new videos against them.
-The CNN is uniquely tailored with specific masks to capture these patterns within the UCF50 dataset.
-The novelty lies in this approach of using action bank features for action recognition.
-The results show that the CNN effectively learns from the data, with a decreasing error rate over iterations, indicating good generalization.
-The paper suggests the potential for applying this method to other datasets and exploring different CNN models.
-In 2015, the use of CNNs in action recognition tasks was at the forefront of research in computer vision, particularly due to their ability to learn hierarchical feature representations directly from raw data.
-The field was rapidly evolving, with new architectures and techniques being proposed to leverage the power of CNNs for this complex task.
-Cheron et al.
-[58] presented an innovative approach to action recognition in videos.
-The core contribution is the development of the Pose-based CNN (P-CNN) descriptor, which uniquely integrates motion and appearance information along the trajectories of human body parts.
-This methodology marks a departure from traditional action recognition techniques that predominantly rely on statistical representations of local motion descriptors.
-By focusing on the detailed analysis of individual body parts through CNNs, the P-CNN model demonstrates a heightened sensitivity to subtle and fine-grained actions, a notable advancement in the field.
-The paper validates the efficacy of this approach using challenging datasets like JHMDB and MPII Cooking Activities, where it shows a marked improvement over existing methods, underscoring the potential of pose-based features in enhancing the accuracy of action recognition systems.
-Geng et al.
-[59] addressed the limitations of traditional classifiers that rely on complex, handcrafted features, the authors propose a model that utilizes CNNs to directly process raw video inputs.
-A key innovation of their work is the introduction of a pre-training strategy employing a Convolutional Auto-Encoder (CAE), aimed at reducing the computational burden associated with kernel training.
-This method not only facilitates more efficient training but also enhances the model’s ability to generalize from raw data, circumventing the need for manually designed features.
-Evaluated on the KTH action database, the proposed model demonstrates superior performance in recognizing human actions compared to existing algorithms that depend on handcrafted features, showcasing its potential for real-world applications in video surveillance, virtual reality, and intelligent human-computer interactions.
-This approach marks a significant advancement in the field of computer vision and deep learning, offering a robust solution to the challenges of HAR.
-Du et al.
-[60] propose a novel, CNN-based framework for action recognition, leveraging the dynamics of human skeletal movements.
-This research departs from traditional methods by transforming skeleton sequences into matrices, which are then converted into images, encapsulating both spatial structures and temporal dynamics.
-These images are fed into a specially designed CNN, where simple max-pooling plays a crucial role in spatial feature selection and temporal frequency adjustment.
-This approach not only preserves the intricate motion characteristics of each body part but also addresses the challenges posed by variable sequence lengths and action frequencies.
-The model’s efficacy is demonstrated through its superior performance on the Berkeley MHAD [61] and the ChaLearn gesture recognition [62] datasets, where it achieves state-of-the-art results, particularly excelling with a significant margin over existing methods in terms of computational efficiency and accuracy.
-This advancement in skeleton-based action recognition represents a significant leap forward, offering robust, efficient solutions for applications in human-computer interaction, surveillance, and beyond.
-Gkioxari et al.
-[63] introduced a novel approach to action recognition in static images, significantly enhancing the accuracy by incorporating contextual cues from the surrounding environment.
-The core innovation lies in the adaptation of the Region-based Convolutional Network (RCNN) to the RCNN model, which uniquely utilizes both a primary region, encompassing the person performing the action, and a secondary region that automatically identifies relevant contextual information.
-This dual-region approach allows R*CNN to capture a more comprehensive understanding of the action by considering additional elements like objects, other people, and the scene itself.
-The model’s effectiveness is demonstrated through its remarkable performance on the PASCAL VOC Actions dataset, where it achieves a mean AP of 90.2%, significantly surpassing existing methods.
-Additionally, its versatility is showcased by its application to the MPII Human Pose dataset and the Berkeley Attributes of People dataset, where it achieves state-of-the-art results in attribute classification.
-This paper’s contribution is pivotal, as it not only advances the field of action recognition by integrating contextual understanding into the recognition process but also opens avenues for fine-grained classification tasks, underscoring the model’s adaptability and robustness.
-The year 2016 witnessed a surge in the use of CNN models for action recognition tasks.
-Numerous models emerged, employing multiple streams and modalities.
-Additionally, there was a notable development of models utilizing 3D-CNNs, as well as hybrid models that combined 2D-CNNs with RNNs to analyze temporal information in videos.
-Bilen et al.
-[64] introduced a groundbreaking approach to video analysis, particularly for action recognition, using CNNs.
-The central innovation is the concept of ”dynamic images,” a compact video representation that encodes the temporal evolution of video frames into a single RGB image.
-This is achieved through an efficient process known as rank pooling.
-The paper presents a novel static/dynamic neural network architecture that enables end-to-end training from videos.
-This architecture uniquely combines static appearance information from still frames with dynamic information from the entire video sequence.
-A significant aspect of this model is the introduction of an approximate rank pooling CNN layer, which greatly speeds up the process compared to traditional rank pooling.
-The paper claims state-of-the-art performance in action recognition on standard benchmarks (HMDB51 and UCF101).
-The dynamic images primarily focus on the acting objects in the video, effectively summarizing actions and motions.
-However, the compression of video content into a single image, while efficient, might oversimplify complex actions or subtle nuances in videos.
-Zhang et al.
-[65] introduces a significant advancement in the field of video-based action recognition by optimizing the deep two-stream architecture for real-time processing.
-The core innovation lies in replacing the computationally intensive optical flow with motion vectors, which are readily available from compressed videos, thus eliminating the need for additional calculations.
-However, motion vectors inherently lack the fine-grained structural details and contain noisy motion patterns, which could degrade recognition performance.
-To address this, the authors propose a novel approach of transferring knowledge from an optical flow CNN (OF-CNN) to a motion vector CNN (MV-CNN).
-This transfer is achieved through three strategies: initialization transfer, supervision transfer, and a combination of both.
-By doing so, the enhanced MV-CNN effectively learns from the more accurate motion patterns captured by OF-CNN, leading to a significant boost in performance.
-The paper demonstrates that this method achieves comparable recognition accuracy to state-of-the-art techniques while processing at an impressive speed of 390.7 frames per second, which is 27 times faster than the original two-stream method.
-This breakthrough not only provides a feasible solution for real-time action recognition but also opens up new possibilities for efficient video analysis in various applications.
-The evaluation of this method on large datasets like UCF101, HMDB51, and THUMOS14 further underscores its effectiveness and practical utility in real-world scenarios.
-Rahmani et al.
-[66] introduced a novel approach for recognizing human actions in 3D from various viewpoints, a significant challenge in video-based HAR due to factors like viewpoint variations, illumination, and visual appearance.
-Central to their methodology is a deep CNN that transforms human poses from different, unknown views into a view-invariant high-level space.
-Uniquely, the training data for this model is synthetically generated by fitting 3D human models to real motion capture data and rendering these poses from multiple viewpoints, circumventing the need for extensive and expensive multiview training data.
-This approach is innovative in its use of synthetic data for training while maintaining the ability to generalize to real depth images of unseen poses without re-training.
-The paper also introduces a group sparse Fourier Temporal Pyramid for robust spatio-temporal representation, enhancing the discriminative power of the model for action recognition.
-The proposed method demonstrates significant improvements over existing methods, achieving higher accuracies on both multiview and single-view benchmark datasets, thereby marking a substantial advancement in the field of 3D action recognition.
-Xin et al.
-[67] introduces a novel approach to action recognition in video, addressing the complexities of real-world environments.
-The core of the paper is the development of the Adaptive Recurrent-Convolutional Hybrid (ARCH) Networks, a hybrid deep neural network (DNN) that integrates static, dynamic, and sequential mixed features for robust action recognition.
-This approach is designed to tackle three key challenges: spatial domain variations (such as noise and occlusion), temporal domain polytropes (differences in action velocities and durations), and intra- and inter-class diversities (similarities and differences within and between action categories).
-The ARCH model, evaluated on the UCF101 dataset and further validated through transfer learning on other datasets, demonstrates superior performance in recognizing human actions, outperforming existing deep-learning-based methods.
-This advancement is significant in its comprehensive handling of the spatial, temporal, and class diversity challenges in action recognition, offering a more effective and adaptable solution for long-term action recognition in uncontrolled and realistic environments.
-By 2017, CNNs had become a key technology in the field of action recognition within computer vision.
-Their strength lay in their ability to autonomously extract meaningful spatial features from images or video frames.
-However, the complexity of action recognition required not just spatial but also temporal analysis.
-This led to the integration of CNNs with architectures capable of temporal processing, like RNNs or LSTMs, or the development of specialized CNN variants like 3D CNNs.
-Challenges such as handling diverse and complex video data, as well as the computational demands of training these models, were significant.
-Nonetheless, the advancements in CNN-based action recognition systems opened up numerous practical applications, ranging from surveillance to interactive gaming, and continued to be a vibrant area of research and development in the field of computer vision.
-Zhao et al.
-[68] introduces a novel approach for action recognition in 3D videos, a field with significant applications in health monitoring, assisted living, and surveillance.
-The study’s core innovation is the integration of RNNs and CNNs in a unique two-stream fusion method.
-The RNNs, utilizing gated-recurrent-units, excel in processing long-term information from optical tracking data, while the 3D-CNNs are adept at handling detailed, recent information from video data.
-This combination allows for a comprehensive analysis of both spatial and temporal data aspects.
-The paper employs the NTU RGB+D Dataset, one of the largest datasets for 3D action recognition, containing over 56,000 action samples in 60 different classes.
-The novelty of the research lies in its two-stream fusion approach, which merges the outputs of RNN and CNN models, significantly enhancing action recognition performance.
-This approach includes two fusion structures: decision fusion and feature fusion, with the latter demonstrating superior results.
-The proposed method shows a remarkable improvement over state-of-the-art methods, outperforming them by more than 14% in both cross-subject and cross-view settings.
-Additionally, the novel RNN structure proposed in the study is optimized for faster convergence and reduced computational demand, making it a significant advancement in the field of action recognition in 3D videos.
-Ullah et al.
-[69] presents an innovative approach to action recognition in video sequences, a challenging area in computer vision.
-The core of their methodology is the integration of Deep Bidirectional Long Short-Term Memory (DB-LSTM) networks with CNN features.
-This combination is designed to effectively capture and analyze the sequential information in video frames, addressing key challenges such as visual content similarity, changes in viewpoint, and varying illumination conditions.
-The model utilizes deep features extracted from every sixth frame using a pre-trained AlexNet, significantly reducing redundancy and computational complexity.
-The method is rigorously evaluated on three benchmark datasets: UCF-101, YouTube 11 Actions, and HMDB51, demonstrating its effectiveness in recognizing a wide range of actions.
-The novelty of this research lies in its efficient use of DB-LSTM for processing sequential frame changes in conjunction with CNN for deep visual data analysis, making it particularly adept at handling lengthy videos.
-The experimental results show that this approach significantly outperforms existing state-of-the-art methods in action recognition, highlighting its potential for applications in areas like surveillance, human-computer interaction, and sports analysis.
-Li et al.
-[70] introduced an innovative approach for action recognition in 3D skeleton videos.
-The core of this research is the development of a method that transforms 3D skeleton data into color images, ensuring translation-scale invariance and independence from the dataset.
-This transformation is crucial as it effectively handles variations in translation and scale, common challenges in 3D skeleton data analysis.
-The study then employs a multi-scale deep CNN, built upon pre-trained models like ResNet, to classify these transformed images.
-This multi-scale CNN is specifically designed to enhance the model’s ability to adjust temporal frequencies, a key aspect when dealing with action recognition.
-The methodology is rigorously tested on the NTU RGB-D dataset, known for its complexity and diversity, and the results demonstrate that the proposed approach significantly outperforms existing methods.
-The novelty of this work lies in its unique image mapping technique and the application of a multi-scale deep CNN, making it a substantial contribution to the field of action recognition from 3D skeleton videos.
-Varol et al.
-[71] introduced a novel approach to action recognition in videos through the use of Long-term Temporal Convolutions (LTC) within CNNs.
-Addressing the limitation of existing methods that typically analyze actions over short video intervals, this study proposes a model that captures the full temporal extent of human actions, which often last several seconds and encompass a wide range of frames.
-The LTC-CNN model comprises five space-time convolutional layers followed by three fully connected layers, designed to extend the temporal reach of the network while balancing complexity by reducing spatial resolution.
-This approach is evaluated on two challenging HAR benchmarks, UCF101 and HMDB51, known for their diverse and complex action categories.
-The paper’s innovation lies in its focus on long-term temporal patterns, a critical aspect often overlooked in action recognition.
-The study not only demonstrates the effectiveness of LTC in capturing extensive temporal information but also underscores the importance of high-quality optical flow estimation for accurate video representation.
-The results show state-of-the-art performance on both datasets, highlighting the model’s superiority in handling the temporal dynamics of human actions.
-In 2018, Zhigang Tu et al.
-[72] introduces a novel approach to video-based HAR using a multi-stream CNN architecture.
-This model, termed Human-Related Multi-Stream CNN (HR-MSCNN), innovatively combines three motion-related and three appearance-related streams to analyze human actions in videos.
-The model employs Block-Sparse Robust Principal Component Analysis (B-RPCA) for effective human detection, addressing challenges like background motion and illumination changes, and uses a motion saliency measure to refine foreground regions and identify primary moving body parts.
-Evaluated on benchmark datasets such as JHMDB, HMDB51, UCF Sports, and UCF101, the HR-MSCNN demonstrates state-of-the-art performance.
-The paper’s novelty lies in its end-to-end CNN that considers both the detected person and discriminative motion regions, a new method for obtaining static and dynamic features, and a fusion method to efficiently combine information from six streams for action prediction.
-The findings highlight the model’s enhanced ability to focus on the most informative parts of the human body for accurate action recognition, marking a significant advancement in the field.
-Arunnehru et al.
-[73] introduced an approach to improving surveillance systems through the application of 3D-CNNs integrated with 3D motion cuboids.
-This method is designed to enhance the accuracy of suspicious action recognition in video surveillance, a critical aspect of ensuring public safety and preventing potential crimes.
-By utilizing the spatial-temporal data encapsulated in 3D motion cuboids and processing this information through 3D-CNNs, the model is capable of detecting and recognizing complex human actions with higher precision than previously established methods.
-The research is validated through extensive testing on two benchmark datasets, KTH and Weizmann, where it demonstrated superior performance in terms of accuracy, thereby outperforming existing action recognition techniques.
-This advancement not only marks a significant contribution to the fields of intelligent video surveillance and computer vision but also opens new avenues for the development of more sophisticated and reliable surveillance systems capable of preemptively identifying and mitigating risks associated with suspicious activities.
-Miao Ma et al.
-[74] introduced a novel approach for recognizing both general and fine-grained human actions in video sequences using a six-stream CNN.
-This method significantly focuses on fine-grained actions, which are more subtle and occur in smaller image regions compared to general actions.
-The core of the approach involves estimating human poses and body parts positions in video sequences, followed by cropping image patches at various scales to capture diverse appearance and motion cues.
-Uniquely, the model employs the outputs of the pooling layer of the CNN, which retains more spatial information than the fully connected layers typically used in such analyses.
-This is complemented by an effective encoding method that generates discriminative action descriptors for classification.
-The model’s efficacy is demonstrated on two public datasets, the Sub-JHMDB and MPII Cooking datasets, where it outperforms several state-of-the-art methods.
-The novelty of this work lies in its combination of coarse human pose tracking with multi-scale patch extraction and advanced pooling layer encoding, making it a significant contribution to the field of fine-grained action recognition in videos.
-Zhang et al.
-[75] introduces a novel approach to real-time action recognition in videos.
-Addressing the limitations of traditional two-stream CNNs, which are hindered by the time-consuming computation of optical flows (OFs), this study proposes the use of motion vectors (MVs) as a more efficient alternative.
-MVs, inherent in video streams, can be extracted with minimal computational overhead, offering a pathway to real-time processing.
-However, MVs inherently contain noise and lack the fine details present in OFs, leading to potential accuracy degradation when used directly in CNNs.
-To overcome this, the paper presents four innovative training strategies that leverage the knowledge from OF-based CNNs to enhance the performance of MV-based CNNs.
-This knowledge transfer is rooted in the structural similarities between MV and OF, enabling the application of insights gained from OF CNNs to improve the generalization ability of MV CNNs.
-The core of this approach is the deeply transferred MV CNN, which employs a method called ”Deeply Connected Transfer” for multi-layer knowledge transfer from OF CNN to MV CNN.
-This method significantly boosts the accuracy of MV CNNs while maintaining the high-speed processing capabilities essential for real-time action recognition.
-The effectiveness of this approach is demonstrated through extensive experiments on various datasets, including UCF101, HMDB51, and THUMOS14, showcasing the model’s ability to achieve high accuracy and processing speed, thus marking a significant advancement in the field of real-time video-based action recognition.
-Liu et al.
-[76] introduced a novel architecture for efficient and real-time action recognition in videos.
-The core of this architecture is the Temporal Convolutional 3D Network (T-C3D), which innovatively combines a residual 3D CNN with a temporal encoding method to capture both the appearance of individual frames and the motion between consecutive frames in a video.
-This approach addresses the computational inefficiencies of existing methods, particularly those reliant on optical flow extraction, by focusing on hierarchical multi-granularity learning.
-The T-C3D model is evaluated on two widely recognized action recognition datasets, UCF101 and HMDB51, where it demonstrates significant improvements in both accuracy and processing speed compared to state-of-the-art real-time methods.
-Notably, the T-C3D achieves over 5.4% higher accuracy and is twice as fast in terms of inference speed, reaching up to 969 frames per second.
-The model’s ability to process RGB frames in real-time while maintaining high accuracy and capturing both short-term and long-term temporal dynamics marks a substantial advancement in the field of video-based action recognition.
-The paper’s contribution lies in its novel integration of 3D-CNN with temporal encoding, enabling efficient modeling of video dynamics and facilitating real-time application potential.
-In 2019, Jiang et al.
-[77] an innovative approach for video action recognition by efficiently integrating spatiotemporal and motion features within a unified 2D CNN framework, a significant departure from the traditional reliance on 3D convolution and optical flow pre-calculation.
-Central to this approach is the STM network, which comprises two novel components: the Channel-wise SpatioTemporal Module (CSTM) and the Channel-wise Motion Module (CMM).
-The CSTM focuses on capturing spatiotemporal features through channel-wise 1D convolution, allowing for effective temporal information fusion with minimal computational overhead.
-In contrast, the CMM is designed to encode motion features by learning distinct motion characteristics for each channel between adjacent time steps.
-This method was rigorously tested on various datasets, including Something-Something v1 & v2, Jester, Kinetics-400, UCF-101, and HMDB-51, demonstrating superior performance over existing state-of-the-art methods.
-The novelty of this research lies in its unique integration of spatiotemporal and motion features in a 2D framework, efficiently handling different types of action recognition scenarios and showcasing its potential in practical applications.
-Crasto et al.
-[78] introduced a groundbreaking approach to action recognition by ingeniously integrating motion information into a standard 3D CNN that processes RGB frames, thereby circumventing the computationally intensive and latency-prone process of calculating optical flow, a common necessity in conventional two-stream action recognition frameworks.
-Central to this innovation are two novel learning strategies: MERS (Motion-Emulated RGB Stream) and MARS (Motion-Augmented RGB Stream).
-MERS trains the network to emulate features from the motion stream, significantly reducing the computational burden, while MARS further advances this concept by synergistically combining appearance and motion information into a singular, efficient stream.
-This approach not only achieves superior performance, outstripping individual RGB and Flow streams with 72.7% accuracy on the Kinetics dataset and 80.1% on HMDB51 (split-1), but also maintains the high standards of two-stream architectures without their inherent computational inefficiencies.
-The paper’s innovative use of a feature-based loss, alongside the standard cross-entropy loss, for network training marks a significant leap forward, showcasing the effective integration of temporal information into a 3D CNN and setting a new benchmark for future research in action recognition.
-Hussein et al.[79] introduced a novel framework aimed at enhancing the recognition of complex human activities in videos.
-Recognizing the limitations of traditional spatiotemporal 3D convolutions, which struggle with the variability and length of complex actions, the authors propose Timeception convolution layers.
-These layers are uniquely designed to handle minute-long temporal patterns, extending the temporal modeling capacity up to 1024 timesteps, significantly surpassing existing methods.
-Timeception achieves this by employing multi-scale temporal convolutions and reducing the complexity inherent in 3D convolutions, making it particularly adept at capturing the nuances of complex actions that consist of multiple, variably-timed one-actions.
-The model’s efficacy is demonstrated through impressive accuracy on challenging datasets like Charades, Breakfast Actions, and MultiTHUMOS, marking a substantial advancement in the field of action recognition.
-The paper’s innovative approach in decomposing and efficiently managing the temporal aspect of videos paves the way for more nuanced and accurate video analysis, particularly in understanding and interpreting complex, multi-faceted human activities.
-He et al.
-[80] presents a sophisticated approach to action recognition in videos through the novel StNet architecture, which adeptly captures both local and global spatial-temporal relationships.
-StNet ingeniously stacks successive video frames into a super-image and applies 2D convolution to efficiently encode local spatial-temporal information, a method that contrasts with traditional CNN+RNN or pure 3D convolution-based approaches.
-To grasp the global spatial-temporal structure, the architecture incorporates temporal convolutions and introduces the innovative Temporal Xception Block (TXB), which separately processes channel-wise and temporal-wise convolutions, optimizing the model for size and computational efficiency.
-Evaluated on the Kinetics dataset, StNet demonstrates superior performance over existing 2D and 3D convolution-based solutions, striking a commendable balance between recognition accuracy and model complexity.
-Furthermore, the model’s robustness and generalizability are underscored by its successful application to the UCF101 dataset, showcasing its potential in advancing the field of video-based action recognition.
-Hao Yang et al.
-[81] presented a novel approach to action recognition, addressing the computational and storage challenges associated with traditional 3D CNNs.
-The authors introduce an innovative model that utilizes efficient asymmetric one-directional 3D convolutions to approximate the conventional 3D convolutions, significantly reducing the number of parameters and computational cost.
-Central to the model are MicroNets, a series of local 3D convolutional networks incorporating multi-scale 3D convolution branches, designed to enhance the feature learning capacity of asymmetric 3D convolutions without increasing computational demands.
-The model also features a multi-source enhanced input method, which fuses information from RGB and Flow frames at the pre-processing stage, thereby circumventing the need for training separate networks for each frame type and further reducing computational requirements.
-Evaluated on two challenging benchmarks, UCF-101 and HMDB-51, the asymmetric 3D-CNN model not only outperforms traditional 3D-CNN models in terms of efficiency and effectiveness but also achieves comparable results with state-of-the-art action recognition methods, marking a significant advancement in the field of computer vision.
-In 2020, Sudhakaran et al.
-[82] introduces a novel approach to video action recognition by proposing the Gate-Shift Module (GSM), a lightweight and efficient module designed to transform standard 2D CNNs into powerful spatio-temporal feature extractors.
-The GSM innovatively applies spatial gating to the spatial-temporal decomposition of 3D kernels, allowing the network to adaptively route and combine features through time with minimal additional computational cost and parameter overhead.
-This approach addresses the challenges of the high dimensionality and computational complexity typically associated with 3D CNNs in video action recognition.
-The paper demonstrates the effectiveness of the GSM through extensive evaluations, where the model achieves state-of-the-art results on the Something Something-V1 and Diving48 datasets and competitive results on the EPIC-Kitchens dataset, all while maintaining a significantly lower model complexity compared to existing methods.
-This breakthrough presents a significant advancement in the field, offering a more efficient and scalable solution for video action recognition tasks.
-Liu x et al.
-[83] investigated the application of the VGG-16 Convolutional Neural Network model to the recognition of human actions from video data.
-This research employs deep learning techniques to analyze kinematic behavior, utilizing two datasets: the KTH dataset, featuring actions like walking and boxing, and the Weizmann dataset, which includes activities such as bending and waving.
-The study’s novelty lies in adapting the VGG-16 model, typically used for static image recognition, to dynamically recognize human behaviors, highlighting its efficacy compared to traditional methods like the histogram of optical flow orientation.
-Findings indicate that the VGG-16 model significantly outperforms other methods, achieving an average accuracy of 91.93% on the KTH dataset and 95.75% on the Weizmann dataset, with particularly high accuracy in recognizing boxing and bending movements.
-These results underscore the potential of advanced CNN models in the real-time analysis of complex human movements, paving the way for improvements in fields such as sports analytics and health monitoring.
-Xiong et al.
-[84] presented a novel approach to enhance Human-Robot Collaboration (HRC) in smart manufacturing by improving the accuracy and reliability of HAR.
-Recognizing the challenges in effectively leveraging temporal information and the scarcity of large training datasets in manufacturing environments, the paper introduces an integrated method that combines optical flow and CNN-based transfer learning.
-Optical flow images, encapsulating the temporal dynamics of human motion, are fed into a two-stream CNN structure, meticulously designed to parse both spatial and temporal information concurrently.
-To circumvent the issue of limited training data, the study explores transfer learning, enabling the adaptation of a pretrained CNN to the specific context of manufacturing scenarios.
-The efficacy of this innovative approach is validated through its application to engine block assembly tasks, showcasing its potential in advancing the capabilities of HRC systems and contributing significantly to the evolution of smart manufacturing practices.
-The paper’s novel integration of optical flow with a two-stream CNN, augmented by transfer learning, marks a substantial stride in addressing the intricate demands of HAR within the realm of industrial automation.
-Sahoo et al.
-[85] innovates in HAR by merging depth information with sequential learning.
-The HAR-Depth model integrates a Deep Bidirectional Long Short Term Memory (DBiLSTM) for capturing the temporal dynamics of actions and employs depth history images (DHI) to enhance shape recognition, where depth-estimated frames are projected onto the X-Y plane.
-Importantly, the model harnesses CNNs for initial feature extraction from action frames, which are then processed through the DBiLSTM.
-Evaluated on datasets such as KTH, UCF Sports, JHMDB, UCF101, and HMDB51, the model demonstrated superior accuracy, notably achieving 97.67% on KTH and 95.00% on UCF Sports.
-This approach highlights the model’s capability to distinguish similar actions through an effective combination of depth-based shape information and sophisticated temporal sequence processing, positioning it as a potent solution for advanced surveillance and interactive systems.
-Abdelbakya et al.
-[86] introduced an approach to HAR using the Principal Component Analysis Network (PCANet).
-This research addresses the computational inefficiencies and the extensive training data requirements commonly associated with traditional CNNs in the context of HAR.
-By employing PCANet, a simplified deep learning model that utilizes Principal PCA for filter learning in convolutional layers, the authors propose a method that significantly reduces computational complexity while maintaining high accuracy in action recognition tasks.
-The methodology involves selecting a subset of frames from each video, extracting feature vectors via PCANet, reducing dimensionality through Whitening Principal Component Analysis (WPCA), and classifying actions using Support Vector Machines (SVM).
-The effectiveness of this approach is demonstrated on the KTH Human Action Dataset, showcasing superior performance compared to existing methods.
-This paper not only presents a computationally efficient alternative for HAR but also opens new avenues for applying simplified deep learning architectures in real-time video analysis and surveillance applications.
-Jaouedi et al.
-[87] intoduced a hybrid deep learning model for human action recognition that incorporates Gated Recurrent Neural Networks (GRNN) as a core component.
-While the primary focus is on the integration of GRNN, Gaussian Mixture Models (GMM), and Kalman Filters (KF) for motion detection and tracking, the model is rooted in deep learning principles similar to those underlying CNNs.
-The approach leverages the sequential data processing capabilities of GRNNs, which are a variant of RNNs, to capture temporal dependencies in video data for action recognition.
-Although the paper emphasizes GRNN for its ability to handle temporal sequences effectively, the overall architecture aligns with deep learning methodologies, including those used in CNNs, for feature extraction and classification tasks in video analysis.
-This innovative approach aims to overcome the inherent challenges in HAR, such as variations in viewpoint, background clutter, and environmental conditions, by leveraging the temporal processing strengths of GRNNs for sequential data, the motion detection capabilities of GMMs, and the object tracking precision of KFs.
-Evaluated on three challenging datasets—UCF Sports, UCF101, and KTH—the model demonstrates its effectiveness by achieving an impressive average accuracy of 96.3% on the KTH dataset, highlighting its potential applicability in a wide range of domains, from intelligent video surveillance to human-computer interaction.
-This research marks a significant advancement in the field of computer vision by proposing a comprehensive solution that not only enhances the accuracy of human action recognition but also opens new avenues for future exploration in deep learning and video analysis technologies.
-Asghari et al.
-[88] introduced a groundbreaking approach to HAR by intricately capturing the dynamics of human body motion in videos, a critical aspect that has been underexplored in previous research.
-Central to this approach is a novel dynamic encoder model that processes pose heat-map sequences, extracted from video frames using a pose estimator, to generate a fixed-size representation called DynaMotion.
-This representation encapsulates the temporal evolution of body joint heat-maps, providing a rich, comprehensive video-level representation for action classification.
-The authors demonstrate the efficacy of this method by integrating the DynaMotion representation with appearance and motion streams in a two-stream network architecture, leading to superior performance on several benchmark datasets including HMDB, JHMDB, UCF-101, and AVA.
-The paper’s innovative approach of leveraging the dynamical information from each pixel in the video marks a significant advancement in the field of HAR, showcasing the potential of combining appearance and motion information with temporal dynamics for enhanced action recognition accuracy.
-In 2021 Wang et al.
-[89] introduced a novel approach to video action recognition by embedding a multipath excitation module, named ACTION, into 2D CNNs.
-This innovative module addresses the computational limitations of 3D CNNs while effectively capturing spatial-temporal, channel-wise, and motion patterns crucial for action recognition.
-The ACTION module comprises three distinct paths: Spatio-Temporal Excitation (STE) for characterizing spatio-temporal representations using 3D convolution, Channel Excitation (CE) for adaptively recalibrating channel-wise feature responses, and Motion Excitation (ME) for highlighting motion-sensitive channels by calculating feature-level temporal differences.
-The integration of these paths into the ACTION-Net model allows for a comprehensive and efficient analysis of video data, leading to significant performance improvements on three benchmark datasets: Something-Something V2, Jester, and EgoGesture.
-The paper demonstrates that ACTION-Net, equipped with the ACTION module, consistently outperforms its 2D CNN counterparts, offering a computationally efficient yet highly effective solution for action recognition in videos.
-Wu et al.
-[90] presented an approach to video-based action recognition by extending 3D CNNs to not just RGB data but also depth and pose data modalities.
-Recognizing the limitations of single-modality (RGB) in capturing the comprehensive dynamics of actions, the authors propose a novel multimodal two-stream 3D network framework.
-This framework innovatively integrates Depth Residual Dynamic Image Sequence (DRDIS) and Pose Estimation Map Sequence (PEMS) as two distinct yet complementary data streams.
-DRDIS effectively encapsulates the spatial-temporal progression of actions in depth videos, while PEMS focuses on intuitively representing the spatial configuration of body parts, free from the distractions of cluttered backgrounds.
-The fusion of these streams aims to harness the rich, multimodal data, thereby significantly enhancing the action recognition performance.
-The model’s superiority is empirically validated through extensive experiments on four challenging datasets, namely SBU Kinect Interaction, UTD-MHAD, and NTU RGB+D 60 & 120, showcasing its potential to set new benchmarks in the domain of video-based action recognition.
-The paper’s novelty lies in its pioneering application of 3D CNNs to multimodal action data, marking a significant stride in the field and opening new avenues for future research.
-Lee et al.
-[91] introduced an advanced video prediction model that integrates Long-term Motion Context Memory (LMC-Memory) with a memory alignment learning process, significantly improving the capability to utilize long-term motion contexts from brief input sequences.
-A notable aspect of the model’s architecture is the use of 3D Convolutional Networks (C3D) to extract motion features, which are then processed through this novel memory system to align and recall dynamic contexts effectively.
-The model was evaluated using both synthetic (Moving-MNIST) and real-world (KTH Action and Human 3.6M) datasets, demonstrating its superior performance in predicting future frames accurately, particularly in scenarios that require understanding extended activities from limited data.
-This innovation addresses a critical gap in traditional video prediction techniques, which often struggle with long-term temporal dependencies, by enhancing the model’s ability to predict complex motions over time through the strategic integration of CNNs with memory-based learning.
-Wang et al.
-[92] introduced a novel video architecture, the Temporal Difference Network (TDN), aimed at enhancing action recognition in videos by efficiently capturing multi-scale temporal information.
-Addressing the challenges of temporal modeling, the TDN employs an innovative Temporal Difference Module (TDM) that operates on two levels: local and global motion modeling.
-This approach refines motion patterns for 2D CNNs using temporal differences over consecutive frames and captures long-range structures through temporal differences across segments.
-The TDN framework is designed to be integrated seamlessly with existing CNNs, offering a principled and efficient solution for action recognition.
-The paper demonstrates the model’s superiority through extensive experiments on the Something-Something V1 & V2 and Kinetics-400 datasets, where TDN sets new benchmarks, showcasing its ability to capture nuanced motion patterns and long-range temporal structures.
-The introduction of TDN marks a significant advancement in the field, providing a robust alternative to 3D convolutions and pushing the boundaries of video-based action recognition.
-Liu et al.
-[93] introduced an advanced approach for recognizing human actions, particularly focusing on abnormal behaviors in video surveillance.
-The core of the methodology is a two-stream CNN model that synergistically combines motion history images (MHI) and RGB frames to enhance the accuracy of action recognition.
-The model comprises two primary components: VMHI, which employs the VGG-16 network trained on MHIs to capture motion information, and FRGB, which utilizes the Faster R-CNN algorithm trained on RGB frames with Kalman filter-assisted data annotation to capture appearance information.
-The integration of these two streams allows the model to effectively recognize not only individual human actions but also complex interactions between people, significantly improving the differentiation of similar actions.
-The model’s robustness and superior performance are demonstrated through extensive experiments on the KTH, Weizmann, UT-interaction, and TenthLab datasets, where it outperforms existing state-of-the-art methods.
-This innovative approach addresses critical challenges in HAR, offering significant improvements in accuracy and applicability in real-world surveillance scenarios.
-In 2022, Sanchez et al.
-[94] presented 3DFCNN, a novel 3D fully CNN designed for real-time HAR using raw depth image sequences.
-This end-to-end trainable model is optimized for high accuracy and real-time performance, distinguishing itself by its ability to automatically encode spatio-temporal patterns from depth sequences while ensuring privacy protection, as identities cannot be discerned from depth data.
-The 3DFCNN model, characterized by its reduced architecture and computational efficiency, outperforms non-DNN-based state-of-the-art methods and competes closely with DNN-based approaches, achieving a maximum accuracy of 83.6% with a significantly lower computational cost of 1.09 seconds.
-Evaluated on three widely recognized datasets—NTU RGB+D Action Recognition Dataset, NorthWest-UCLA Multiview Action 3D, and UWA3D Multiview Activity II—the model demonstrates its robustness and applicability in real-world scenarios, marking a significant advancement in the field of action recognition, especially in contexts where privacy is paramount.
-Anshul Shah et al.
-[95] presented a novel Joint-Motion Reasoning Network (JMRN) for action recognition, focusing on the dynamics of human joints.
-Unlike traditional methods that rely on RGB or optical flow features, JMRN emphasizes joint-based action recognition.
-The model operates by first extracting motion features from each joint independently through a shared motion encoder and then performing collective reasoning.
-A distinctive aspect of JMRN is the joint selector module, which re-weights joint information to highlight the most discriminative joints for the task at hand.
-Additionally, the paper introduces a joint-contrastive loss to ensure that per-joint motion signatures are expressive of the action, and a geometry-aware data augmentation technique to enhance the model’s generalization capabilities.
-Evaluated on several datasets including JHMDB, HMDB, Charades, AVA, and Mimetics, JMRN demonstrates substantial improvements over state-of-the-art joint-based approaches.
-The model’s robustness is further validated by its performance on the Mimetics dataset, which consists of out-of-context actions.
-The paper’s approach, focusing on the independent and collective reasoning of joint movements, marks a significant advancement in the field of action recognition, offering a new perspective and methodology that diverges from conventional RGB and optical flow-based models.
-Basha et al.
-[96] introduced a sophisticated approach to enhancing HAR through an innovative sampling technique and deep learning integration.
-Specifically, the study develops a novel Gaussian Weighting Function to aggregate video frames, significantly preserving both spatial and temporal information, which is then processed using a newly proposed 3D CNN combined with LSTM networks.
-This architecture is adept at extracting and classifying spatio-temporal features from reduced datasets without losing critical motion details.
-The efficacy of this model is validated against benchmark datasets like KTH, WEIZMANN, and CASIA-B, achieving notably high accuracy rates—95.78%, 95.27%, and 95.27% respectively.
-These results demonstrate a considerable improvement over existing methods, underlining the potential of the proposed model in real-world applications such as surveillance and interactive systems, where efficient and accurate action recognition is crucial.
-This study not only addresses key challenges in human action recognition by reducing data volume and enhancing information retention but also sets a new benchmark for future research in video-based action recognition.
-Senthilkumar et al.
-[97] introduced a sophisticated approach to HAR, emphasizing the integration of spatial-temporal models with a bi-temporal end-to-end framework.
-The core of this approach is the fusion of Dilated CNNs (DCNN) and deep Bi-LSTM networks, complemented by residual blocks, to effectively detect human activities from video frames.
-The model ingeniously combines the features extracted by CNN with functioning features using skip connections, and employs a temporal dependent relational model to preserve spatial-temporal information during data extraction.
-Evaluated on standard datasets like J-HMDB, UCF11, and Sports, the model demonstrates remarkable performance, achieving up to 98% accuracy in recognizing a wide range of human actions.
-The paper’s novelty lies in its unique integration of LSTM and CNN functions, focusing on the temporal characteristics of skeletal joints, and its ability to organize skeleton information effectively.
-The findings highlight the model’s superior performance and potential in real-world applications, making significant strides in the field of HAR.
-In 2023, Noor et al.
-[98] introduced a novel, computationally efficient 3D-CNN model tailored for real-time action recognition and fall detection.
-Addressing the computational constraints of existing deep learning models in practical applications, this research innovates by proposing a lightweight skeleton-based 3D-CNN framework, significantly enhancing processing speed and accuracy.
-The model’s robustness is further fortified by an augmented human pose dataset, meticulously curated from the AIHub senior abnormal behavior dataset, focusing on fall motion and enriched with the COCO-pose dataset to improve pose estimation accuracy.
-This integration not only optimizes the model for real-world multi-person environments but also sets new benchmarks in fall detection, showcasing the model’s capability to efficiently process extracted pose keypoints and maintain high accuracy in action recognition.
-The paper’s contribution is pivotal in advancing the practical deployment of action recognition systems, promising enhanced performance in real-time applications without compromising on computational efficiency.
-Zong et al.
-[99] presented an innovative approach to video-based action recognition, addressing the challenges of complex video backgrounds and the integration of spatial and temporal information.
-The authors introduce a novel four-stream network model, incorporating spatial and temporal saliency streams alongside traditional appearance and motion streams.
-The model, referred to as the 3M model, is further enhanced with a multi-task learning-based LSTM, adept at capturing long-term dependencies between consecutive frames and ensuring a comprehensive understanding of video dynamics.
-The model’s unique structure and the integration of spatial and temporal saliency streams, coupled with multi-task learning-based LSTM, significantly enhance the model’s ability to recognize and categorize actions in videos, demonstrating superior performance on popular video action recognition datasets.
-If this summary does not match the content of the new document, please provide additional details or attempt to access the document again.
-Table 1: CNN-based Approaches.
-Paper Year Model Dataset Novelty
-Ji et al.
-[52] 2012 3D CNN KTH and TREC Video Retrieval Evaluation (TRECVID) dataset The paper introduces a unique 3D CNN architecture and proposes regularizing the 3D CNN models with auxiliary outputs computed as high-level motion features
-Valle et al. [53] 2013 CNN A custom dataset The novelty lies in the CNN’s ability to process unprocessed images directly, learning to extract features during training.
-Simonyan et al. [35] 2014 Two-stream CNN architecture UCF-101 and HMDB-51 The novelty lies in its two-stream architecture, which separately but concurrently processes spatial and temporal information.
-Sun et al.
-[56] 2014 Two-layer SFA structure with 3D convolution Hollywood2, KTH, and UCF Sports The novelty lies in the combination of SFA with deep learning to learn hierarchical ’slow’ features from video data.
-Ijjina et al. [57] 2014 CNN architecture with linear masks UCF50 dataset The novelty lies in its use of action bank features in combination with a CNN classifier
-Cheron et al.
-[58] 2015 CNN JHMDB and MPII Cooking Activities The introduction of a P-CNN descriptor that uniquely combines motion and appearance information along tracks of human body parts for enhanced HAR.
-Geng et al. [59] 2015 CNN KTH The combination of CNNs with a CAE for pre-training, which addresses the computational inefficiency of kernel training in deep learning models for HAR.
-Du et al.
-[60] 2015 CNN Berkeley MHAD and ChaLearn gesture recognition datasets The novelty lies in its unique approach of transforming skeleton sequences into image representations for efficient and accurate action recognition using a specialized CNN.
-Gkioxari et al.
-[63] 2015 CNN PASCAL VOC Actions, MPII Human Pose and Berkeley Attributes of People Datasets The novelty lies in its unique integration of secondary contextual cues with primary action regions in images using an adapted RCNN model for enhanced action recognition.
-Bilen et al.
-[64] 2016 CNN UCF101 and HMDB51 The novelty lies in its introduction of ’dynamic images,’ a new video representation method that encodes the temporal evolution of frames into a single image for efficient HAR using CNNs
-Zhang et al.
-[65] 2016 CNN UCF101, HMDB51, and THUMOS14 The novelty lies in its use of motion vectors instead of optical flow in a two-stream CNN architecture, enabling real-time action recognition with comparable accuracy to state-of-the-art methods
-Rahmani et al.
-[66] 2016 3D CNN Synthetically generated dataset The novelty lies in its use of a deep CNN trained on synthetically generated depth images for view-invariant human pose representation in 3D action recognition
-Xin et al.
-[67] 2016 Hybrid CNN-RNN model UCF101 The novelty lies in its introduction of the ARCH Networks, a hybrid DNN that innovatively combines CNN and RNN elements for robust long-term HAR in complex, real-world environments
-Zhao et al.
-[68] 2017 Two-stream gated-RNNs and 3D-CNNs NTU RGB+D Dataset The novelty lies in its integration of gated RNNs with 3D-CNNs in a two-stream fusion method for enhanced action recognition in 3D videos data
-Ullah et al.
-[69] 2017 Deep Bidirectional LSTM with CNN UCF-101, YouTube 11 Actions, and HMDB51 The novelty lies in its integration of Deep Bidirectional LSTM with CNN features for efficient and effective HAR in video sequences
-Li et al.
-[70] 2017 Multi-scale deep CNN NTU RGB-D dataset The novelty lies in its translation-scale invariant image mapping method for 3D skeleton video data and the use of a multi-scale deep CNN for HAR.
-Varol et al.
-[71] 2017 3D CNN UCF101 and HMDB51 The novelty lies in its implementation of LTC within 3D CNNs for capturing the full temporal extent of human actions in videos, significantly enhancing HAR accuracy
-Arunnehru et al.
-[73] 2018 3D CNN KTH and Weizmann Datasets The integration of 3D motion cuboids with 3D-CNNs for action recognition, a method not extensively explored in previous studies.
-This approach allows for a more detailed by considering the spatial-temporal dynamics of movements
-Zhigang Tu et al.
-[72] 2018 Human-Related Multi-Stream CNN JHMDB, HMDB51, UCF Sports and UCF101 The novelty lies in its end-to-end multi-stream CNN that uniquely combines human detection and motion saliency for enhanced HAR.
-Miao Ma et al.
-[74] 2018 Six-stream CNN Sub-JHMDB dataset and MPII Cooking dataset The novelty lies in its unique approach of using a six-stream CNN that leverages coarse human pose tracking and multi-scale patch extraction, combined with advanced pooling layer encoding, to enhance both general and fine-grained HAR in videos
-Zhang et al. [75] 2018 Two-stream-based CNN UCF101, HMDB51 and THUMOS14 The novelty lies in its use of motion vectors instead of optical flows for real-time HAR
-Liu et al.
-[76] 2018 A novel temporal-residual 3D CNN UCF101 and HMDB51 The novelty lies in its unique integration of a residual 3D CNN with a temporal encoding method, enabling efficient and real-time HAR by capturing both short-term and long-term temporal dynamics in videos
-Jiang et al.
-[77] 2019 Spatio-Temporal and Motion (STM) CNN-based network Something-Something v1 & v2, Jester, Kinetics-400, UCF-101 and HMDB-51 The novelty lies in its integration of spatiotemporal and motion features within a unified CNN framework, avoiding the computational intensity of 3D convolutions and optical flow calculations
-Crasto et al.
-[78] 2019 3D CNN Kinetics and HMDB51 The novelty lies in introducing MARS, a single-stream approach that integrates motion information into RGB frames without the need for explicit optical flow computation.
-Hussein et al.[79] 2019 Integrated 2D CNN and 3D CNN architecture Charades, Breakfast Actions, and MultiTHUMOS The novelty lies in introducing Timeception CNN layers, which leverage multi-scale temporal convolutions to efficiently model minute-long temporal patterns in complex actions.
-He et al.
-[80] 2019 Integrated 2D CNN and 3D CNN architecture Kinetics and UCF101 The novelty lies in the StNet architecture, which uniquely combines super-images with 2D convolutions and a Temporal Xception Block for efficient and effective local and global spatial-temporal modelling in HAR.
-Hao Yang et al.
-[81] 2019 Asymmetric 3D CNN UCF-101 and HMDB-51 The paper introduces efficient asymmetric one-directional 3D convolutions and MicroNets, significantly reducing computational costs and parameters while enhancing feature learning in HAR
-Sudhakaran et al.
-[82] 2020 2D CNN Something Something-V1, Diving48 and EPIC-Kitchens The paper introduces the GSM, a lightweight module that transforms 2D-CNNs into efficient spatio-temporal feature extractors for video HAR with minimal computational overhead
-Liu x et al.
-[83] 2020 2D CNN KTH and Weizmann datasets The paper introduced the adaptation of the VGG-16 deep learning model to kinematic human behavior recognition from video data, demonstrating enhanced accuracy compared to traditional dynamic recognition methods.
-Xiong et al.
-[84] 2020 Two-stream 2D CNN structure Engine block assembly The paper introduces a novel transferable two-stream CNN integrating optical flow and transfer learning for robust and accurate HAR in smart manufacturing environments
-Sahoo et al.
-[85] 2020 2D CNN with DBiLSTM KTH, UCF Sports, JHMDB, UCF101, and HMDB51 The paper introduced the integration of DHI with a CNN and a DBiLSTM network, enhancing human action recognition by capturing both depth-informed shape details and temporal dynamics
-Abdelbakya et al.
-[86] 2020 PCA KTH dataset The novelty lies in its application of PCANet (a form of Convolutional Neural Network) for HAR, presenting a less computationally intensive alternative to traditional CNNs.
-Jaouedi et al.
-[87] 2020 CNN KTH the paper introduces a hybrid deep learning model for human action recognition that incorporates GRNN as a core component.
-While the primary focus is on the integration of GRNN, Gaussian Mixture Models (GMM), and Kalman Filters (KF) for motion detection and tracking, the model is rooted in deep learning principles similar to those underlying CNNs.
-Asghari et al.
-[88] 2020 2D CNN HMDB, JHMDB, UCF-101, and AVA The paper introduces DynaMotion, a dynamic motion representation that captures the temporal evolution of body joint heat-maps, significantly enhancing HAR in videos
-Wang et al.
-[89] 2021 Incorporated 2D-CNN and 3D-CNN model Something-Something V2, Jester, and EgoGesture The paper introduces ACTION-Net, a novel network embedding the ACTION module that enhances 2D CNNs with multipath excitation, capturing spatial-temporal, channel-wise, and motion information efficiently for superior HAR
-Lee et al.
-[91] 2021 3D CNN Moving-MNIST for synthetic tests, and KTH Action and Human 3.6M datasets The paper introduced the integrating LMC-Memory with 3D CNN and a memory alignment learning technique, enabling the model to precisely recall complex long-term motion contexts from short video sequences for enhanced video prediction.
-Wu et al.
-[90] 2021 Multimodal two-stream 3D-CNN network SBU Kinect Interaction, UTD-MHAD, NTU RGB+D 60 & 120 The novelty lies in its pioneering extension of 3D CNNs to depth and pose data modalities for video HAR, introducing a multimodal two-stream 3D
-Wang et al.
-[92] 2021 CNN model Something-Something V1 & V2 and Kinetics-400 The novelty lies in the introduction of the TDN, which efficiently captures multi-scale temporal information for action recognition by leveraging a novel TDM at both local and global motion levels
-Liu et al.
-[93] 2021 Two-stream 2D-CNN network KTH, Weizmann, UT-interaction and TenthLab The novelty lies in its integration of motion history images with RGB frames in a two-stream CNN, significantly enhancing the accuracy of human abnormal behavior recognition in video surveillance
-Sanchez et al. [94] 2022 3D-CNN NTU RGB+D, NorthWest-UCLA and UWA3D The novelty lies in the introduction of 3DFCNN, a real-time, privacy-preserving 3D fully CNN for HAR using raw depth data
-Anshul Shah et al.
-[95] 2022 Not mentioned JHMDB, HMDB, Charades, AVA and Mimetics datasets The novelty lies in the introduction of JMRN architecture, which uniquely combines independent per-joint motion feature extraction with collective reasoning and a joint-contrastive loss, enhancing the accuracy and robustness of joint-based HAR
-Basha et al.
-[96] 2022 3D CNN with LSTM KTH, WEIZMANN, and CASIA-B Human Activity and Gait datasets The paper introduced the use of a Gaussian Weighting Function for video frame sampling combined with a 3D CNN-LSTM architecture, significantly enhancing the preservation and classification of spatio-temporal features in human action recognition.
-Senthilkumar et al.
-[97] 2022 Dilated CNN with deep Bi-LSTM J-HMDB and UCF-11 The novelty lies in its integration of Dilated CNNs with deep Bi-LSTM networks and a temporal dependent relational model, offering a robust spatial-temporal framework for accurate and efficient HAR
-Noor et al.
-[98] 2023 3D-CNN COCO-pose, AIHub senior abnormal behavior and its Augmented Human Pose dataset The novelty is the introduction of a lightweight skeleton-based 3D-CNN and an augmented human pose dataset, improving real-time HAR and fall detection accuracy while reducing computational requirements
-Zong et al.
-[99] 2023 A combination of 2D CNNs and LSTMs architeture UCF101 Action, HMDB51 Action and Large-scale Kinetics Action Datasets The novelty lies in its introduction of a spatial and temporal saliency-based four-stream network with multi-task learning, enhancing HAR by capturing salient object and motion and modelling long-term dependencies
-3.2 Transformer-based Approaches in HAR
-The ViT was conceptualized and introduced in the academic domain through the paper authored by Dosovitskiy et al.
-[6] in 2020.
+Consequently, numerous methods have been proposed to enhance the the two-stream model [99, 33, 100, 75, 104].
+In 2016, building on the the two-stream CNN, [33] focused on improving the two-stream CNN by exploring various fusion strategies for combining spatial and temporal streams, resulting in better performance on the UCF-101 and HMDB-51 datasets.
+By enhancing fusion techniques, this work addressed the limitations of the initial two-stream model, leading to more effective integration of spatial and temporal information.
+[100] introduced temporal segment networks (TSN).
+This work aimed to capture long-range temporal structures for action recognition, achieving significant improvements on the UCF-101 and HMDB-51 datasets by dividing videos into segments for comprehensive analysis.
+The introduction of TSN extended the temporal analysis capabilities of the two-stream CNN, enabling the capture of long-range dependencies.
+In 2017, derived from the two-stream CNN, [20] proposed a three-stream method by using spatio-temporal vectors, with locally max-pooled features to enhance performance.
+Tested on the UCF-101 and HMDB-51 datasets, the approach demonstrated improved recognition accuracy by efficiently capturing spatio-temporal dynamics.
+In 2018, the efficient convolutional network for online video understanding (ECO) was introduced by [114], combining the two-stream CNN approach with lightweight 3D CNNs, and focusing on efficiency and real-time processing, with high efficiency and competitive accuracy demonstrated on the Kinetics and UCF-101 datasets.
+[32] introduced the SlowFast network which processes video data at varying frame rates to capture both spatial semantics and motion dynamics, achieving state-of-the-art results on the Kinetics-400 and Charades datasets.
+By introducing different temporal resolutions, this work innovated on the two-stream concept, capturing fine and coarse temporal details.
+[101] expanded on their previous work with TSN, developing a multi-stream approach that incorporated RGB, optical flow, and warped optical flow streams to model long-range temporal structures more effectively.
+This approach achieved state-of-the-art results by capturing both spatial and temporal information across various time scales.
+In 2021, temporal difference networks (TDN) were introduced by [98], leveraging the multi-stream CNN with a focus on capturing motion dynamics efficiently.
+Using the UCF-101 and HMDB-51 datasets, TDN achieved notable improvements by effectively modeling temporal differences.
+By emphasizing temporal differences, this work advanced the ability of the two-stream CNN to capture motion dynamics more effectively.
+Table LABEL:table:TwoStreamNetworks presents the works discussed in this section that utilized two or more stream CNNs approaches.
+Table 1: Two-stream CNN-based approaches in HAR.
+Paper Model Dataset Novelty
+[81] Two-stream CNN UCF-101, HMDB-51 Introduced the two-stream architecture separating spatial and temporal streams for effective action recognition.
+[33] Two-stream CNN UCF-101, HMDB-51 Explored various fusion strategies to combine spatial and temporal streams, and improved performance.
+[100] Two-stream CNN + TSN UCF-101, HMDB-51 Introduced TSN to capture long-range temporal structures by dividing videos into segments.
+[20] Three-Stream CNN UCF-101, HMDB-51 Proposed a three-stream method using spatio-temporal vectors with locally max-pooled features for enhanced performance.
+[114] Two-stream CNN + 3D CNN Kinetics, UCF-101 Combined the two-stream CNN with lightweight 3D CNNs for efficient real-time processing.
+[32] Two-stream CNN + SlowFast Kinetics-400, Charades Introduced SlowFast networks processing video data at varying frame rates to capture both spatial and motion dynamics.
+[101] CNN-RNN, (Multi-stream TSN) UCF101, HMDB51 Expanded on TSN by developing a multi-stream approach that incorporated RGB, optical flow, and warped optical flow streams to model long-range temporal structures more effectively.
+[98] Multi-stream CNN + TDN Something-Something V1 and V2 Introduced TDN focusing on capturing motion dynamics efficiently.
+3.1.2 3D CNN-Based Approaches
+The foundational work conducted by [49] introduced 3D CNNs for HAR, demonstrating their effectiveness in capturing spatio-temporal features on the KTH and UCF-101 datasets and outperforming traditional 2D CNNs.
+The work paved the way for further research on enhancing 3D convolutional models.
+[90] introduced C3D, a generic 3D CNN for spatio-temporal feature learning, achieving state-of-the-art performance on the Sports-1M and UCF-101 datasets and highlighting the scalability and effectiveness of 3D convolutions.
+Building on the work by [49], C3D demonstrated the potential of 3D CNNs across diverse datasets, influencing subsequent research in 3D CNNs.
+[95] introduced long-term temporal convolutions to capture extended motion patterns.
+This work improved the accuracy on the UCF-101 and HMDB-51 datasets and emphasized the importance of long-term motion information.
+Moreover, this study extended the temporal scope of 3D CNNs, highlighting the need for capturing long-term motion for accurate action recognition.
+In the same year, [76] proposed pseudo-3D residual networks (P3D), which combined 2D and 3D convolutions to balance the accuracy and computational complexity.
+This work achieved competitive performance on the Kinetics and UCF-101 datasets.
+Moreover, P3D networks offered a more efficient approach by blending 2D and 3D convolutions, further refining the capabilities of 3D CNNs.
+Additionally, [15] introduced I3D by inflating 2D convolutions to 3D, achieving significant improvements on the Kinetics dataset by leveraging ImageNet pre-training, thereby setting new performance benchmarks.
+I3D bridged the gap between 2D and 3D CNNs, demonstrating the benefits of transfer learning in 3D convolutional models.
+[42] evaluated the scalability of 3D CNNs with increased data and model sizes, demonstrating that deeper 3D CNNs can achieve better performance on the Kinetics and UCF-101 datasets, paralleling the success of 2D CNNs on ImageNet.
+This study emphasized the need for larger datasets and deeper models in 3D convolutional research, highlighting the potential of 3D CNNs to retrace the historical success of 2D CNNs.
+Building on these insights, [26] introduced a new temporal 3D ConvNet architecture with enhanced transfer learning capabilities, demonstrating superior performance on the UCF-101 and HMDB-51 datasets through architectural innovations and effective transfer learning.
+This work underscored the importance of architectural innovation and transfer learning, pushing the boundaries of 3D CNN performance and further advancing the field of action recognition.
+[92] further contributed by conducting a comprehensive analysis of spatio-temporal convolutions, highlighting the benefits of factorizing 3D convolutions into separate spatial and temporal components, achieving state-of-the-art results on the Kinetics and UCF-101 datasets.
+This dissection provided insights that informed subsequent model designs and optimizations.
+In the same year, [107] explored the trade-offs between speed and accuracy in spatio-temporal feature learning, proposing efficient 3D CNN variants that balance computational cost and recognition performance on the Kinetics and UCF-101 datasets.
+Their work highlighted the practical considerations of deploying 3D CNNs, emphasizing the need to balance speed and accuracy, thereby refining the approach to spatio-temporal feature learning.
+Additionally, [103] introduced non-local neural networks to capture long-range dependencies, demonstrating that non-local operations significantly improve the modeling of complex temporal relationships and enhance action recognition performance on the Kinetics and Something-Something datasets.
+By integrating non-local operations, this study advanced the ability of 3D CNNs to capture complex temporal patterns, further pushing the boundaries of spatio-temporal modeling.
+[32] introduced SlowFast Networks, a novel approach that processes video at different frame rates to capture both slow and fast motion dynamics, and achieved state-of-the-art results on the Kinetics-400 and Charades datasets.
+This innovation highlighted the importance of capturing varied motion dynamics for improved video recognition.
+In the same year, [91] presented channel-separated convolutional networks (CSN), which reduced computational complexity by separating convolutions by channel, demonstrating efficiency without sacrificing accuracy on the Kinetics and Sports-1M datasets.
+This approach contributed to the development of more computationally feasible models.
+Concurrently, [37] leveraged large-scale weakly-supervised pre-training on video data, significantly boosting performance on the IG-65M and Kinetics datasets and underscoring the potential of massive datasets in enhancing 3D CNN capabilities.
+Additionally, [57] proposed resource-efficient 3D CNNs using depthwise separable convolutions and achieved competitive accuracy with significantly reduced computational requirements on the Kinetics-400 and UCF-101 datasets.
+This work emphasized the importance of optimizing 3D CNNs for computational efficiency, further advancing the field of action recognition.
+[31] proposed X3D, a family of efficient video models by expanding architectures along multiple axes.
+It achieved state-of-the-art performance with reduced model complexity on the Kinetics-400 and Charades datasets.
+X3D highlighted the significance of model efficiency in balancing performance and computational demands.
+In the same year, [63] introduced an efficient 3D CNN with a temporal attention mechanism and achieved high accuracy with efficient computation by focusing on salient temporal features on the Kinetics-400 and UCF-101 datasets.
+This work demonstrated the potential of selectively focusing on important temporal features to enhance the efficiency and accuracy of 3D CNNs, further advancing the field of action recognition.
+Table LABEL:table:3DCNNBased presents the works discussed in this section that utilized 3D CNN approaches.
+Table 2: 3D CNN-based approaches in HAR.
+Paper Model Dataset Novelty
+[49] 3D CNN UCF-101, HMDB-51 Introduced 3D CNNs for HAR, effectively capturing spatio-temporal features and outperforming 2D CNNs.
+[90] 3D CNN Sports-1M, UCF-101 Introduced C3D, a generic 3D CNN for spatio-temporal feature learning, and achieved state-of-the-art performance.
+[95] 3D CNN UCF-101, HMDB-51 Introduced long-term temporal convolutions to capture extended motion patterns, and improved accuracy.
+[76] 3D CNN Kinetics, UCF-101 Proposed P3D networks combining 2D and 3D convolutions, balancing accuracy and computational complexity.
+[15] 3D CNN Kinetics Introduced I3D by inflating 2D convolutions to 3D, leveraging ImageNet pre-training for significant improvements.
+[42] 3D CNN Kinetics, UCF-101 Evaluated the scalability of 3D CNNs with increased data and model sizes, and showed parallels to 2D CNN success.
+[26] 3D CNN UCF-101, HMDB-51 Introduced a new temporal 3D ConvNet architecture with enhanced transfer learning capabilities.
+[92] 3D CNN Kinetics, UCF-101 Conducted a comprehensive analysis of spatio-temporal convolutions, and highlighted the benefits of factorizing 3D convolutions.
+[107] 3D CNN Kinetics, UCF-101 Explored speed-accuracy trade-offs in spatio-temporal feature learning, and proposed efficient 3D CNN variants.
+[103] 3D CNN Kinetics, Something-Something Introduced non-local operations to capture long-range dependencies, and improved modeling of complex temporal relationships.
+[32] 3D CNN Kinetics-400, Charades Proposed SlowFast networks to process video at different frame rates, capturing both slow and fast motion dynamics.
+[91] 3D CNN Kinetics, Sports-1M Introduced CSN to reduce computational complexity without sacrificing accuracy.
+[37] 3D CNN IG-65M, Kinetics Leveraged large-scale weakly-supervised pre-training on video data, and significantly boosted performance.
+[57] 3D CNN Kinetics-400, UCF-101 Proposed resource-efficient 3D CNNs using depthwise separable convolutions, and achieved competitive accuracy with reduced computational requirements.
+[31] 3D CNN Kinetics-400, Charades Proposed X3D, a family of efficient video models by expanding architectures along multiple axes.
+[63] 3D CNN Kinetics-400, UCF-101 Introduced a temporal attention mechanism to enhance efficiency and accuracy in 3D CNNs.
+3.1.3 CNN-RNN-Based Approaches
+The integration of CNNs and RNNs for HAR was significantly advanced by the work of [28], who introduced long-term recurrent convolutional networks (LRCN).
+This approach effectively combined the spatial feature extraction capabilities of CNNs with the temporal dynamics modeling of LSTMs, demonstrating substantial improvements in action recognition tasks on datasets like UCF-101 and HMDB-51.
+Building on this foundation, [111] extended the application of deep networks to video classification by integrating deep CNNs with LSTMs to handle longer video sequences.
+Their method, tested on the Sports-1M and UCF-101 datasets, highlighted the importance of capturing extended temporal dependencies for improved performance in complex video classification tasks.
+Further pushing the boundaries, [84] explored unsupervised learning of video representations using LSTMs.
+By leveraging LSTMs to learn spatio-temporal features without labeled data, their approach demonstrated effective video representation learning on the UCF-101 dataset, showcasing the versatility and potential of CNN-RNN architectures in both supervised and unsupervised learning scenarios for HAR.
+The development of CNN-RNN architectures for HAR saw significant advancements in 2016.
+[105] proposed a hybrid deep learning framework that modeled spatial-temporal clues by combining CNNs for spatial feature extraction with RNNs for temporal sequence modeling.
+Their approach, tested on the UCF-101 and HMDB-51 datasets, demonstrated substantial improvements in video classification accuracy.
+Additionally, [64] expanded the application of CNN-RNN architectures to real-time scenarios with their approach for online human action detection using joint classification-regression RNNs.
+Combining CNNs for spatial features and RNNs for temporal dynamics, their method, tested on the J-HMDB and UCF-101 datasets, achieved notable improvements in accuracy and efficiency, showcasing the practicality of CNN-RNN models in real-time action detection.
+Building on these advancements, 2017 and 2018 witnessed further refinements and innovations in CNN-RNN architectures for HAR.
+[65] introduced VideoLSTM, integrating convolutions, attention mechanisms and optical flow within a recurrent framework, and demonstrating improved performance on the UCF101 and HMDB51 datasets.
+[15] made a significant contribution with the two-stream Inflated 3D ConvNet (I3D), which inflated 2D CNN architectures into 3D and combined them with RNNs for temporal modeling.
+The model was evaluated on the Kinetics dataset, as well as UCF101 and HMDB51.
+[94] proposed a novel architecture combining CNNs with bi-directional LSTMs, effectively utilizing both spatial and temporal information from video sequences and showing superior performance on the UCF-101 and HMDB-51 datasets.
+In 2020, in the realm of human activity recognition using sensor data, [106] proposed an LSTM-CNN architecture that effectively captured both temporal dependencies and local feature patterns, showing improved accuracy on the WISDM, UCI HAR, and OPPORTUNITY datasets.
+Similarly, [72] developed a CNN-LSTM approach for smartphone sensor-based activity recognition, demonstrating high accuracy on the UCI HAR dataset and further validating the effectiveness of combining CNNs and RNNs for processing time-series data in activity recognition tasks.
+Recent advancements in HAR have leveraged sophisticated CNN-RNN architectures to enhance performance and reduce computational complexity.
+[71] introduced an attention-based LSTM network combined with dilated CNN features, and significantly improved the recognition accuracy on the UCF-101 and HMDB-51 datasets by capturing essential spatial features through dilated convolutions and temporal patterns with attention mechanisms.
+Building on this, [70] focused on multiview HAR; utilizing a CNN-LSTM architecture to cascade pose features, they achieved high accuracy (94.4% on the MCAD dataset and 91.67% on the IXMAS dataset) while reducing the computational load by targeting pose data rather than entire images.
+Table LABEL:table:CNNRNNbasedApproaches presents the works discussed in this section that utilized CNN-RNN approaches.
+Table 3: CNN-RNN-based approaches in HAR.
+Paper Model Dataset Novelty
+[28] CNN-RNN, (LRCN) UCF-101, HMDB-51 Combined CNNs for spatial feature extraction with LSTMs for temporal dynamics.
+[111] CNN-RNN Sports-1M, UCF-101 Integrated deep CNNs with LSTMs to handle longer video sequences, capturing extended temporal dependencies.
+[84] CNN-RNN, (Unsupervised LSTM) UCF-101 Explored unsupervised learning of video representations using LSTMs, leveraging spatiotemporal features.
+[105] CNN-RNN UCF-101, HMDB-51 Modeled spatial-temporal clues by combining CNNs for spatial features with RNNs for temporal sequence modeling.
+[64] CNN-RNN J-HMDB, UCF-101 Applied CNN-RNN architectures to real-time scenarios for online human action detection.
+[65] CNN-RNN (VideoLSTM) UCF-101, HMDB-51 Integrated convolutions, attention mechanisms, and optical flow within a recurrent framework.
+[15] 3D CNN-RNN Kinetics, UCF101, HMDB51 Inflated 2D CNN architectures into 3D, and combined them with RNNs for temporal modeling.
+[94] CNN-RNN, (CNN-BiLSTM) UCF101, HMDB51 Combined CNNs with bi-directional LSTMs to utilize both spatial and temporal information.
+[106] CNN-RNN WISDM, UCI, OPPORTUNITY Captured both temporal dependencies and local feature patterns for human activity recognition using sensor data.
+[72] CNN-RNN UCI Developed a CNN-LSTM approach for smartphone sensor-based activity recognition, and demonstrated high accuracy.
+[71] CNN-RNN, (CNN-Attention-LSTM) UCF-101, HMDB-51 Improved recognition accuracy with attention-based LSTM network combined with dilated CNN features.
+[70] CNN-RNN MCAD, IXMAS Achieved high accuracy in multiview HAR by cascading pose features using a CNN-LSTM architecture.
+3.2 ViT-Based Approaches in HAR
+In 2020, the ViT was conceptualized and introduced in the academic domain through the paper authored by [29].
 The ViT marked a paradigm shift in still image recognition methodologies, applying the Transformer model, predominantly known for its success in NLP, to the realm of computer vision.
@@
 The adaptation of ViT to action recognition signifies the model’s versatility and its potential for broader applications in computer vision beyond static image analysis.
-In 2021, Arnab et al.
-[30] introduced a novel approach to video classification by leveraging the capabilities of Transformer models, traditionally successful in image classification.
-ViViT, a Pure-transformer based model, ingeniously processes spatio-temporal tokens extracted from videos through a series of Transformer layers, addressing the challenge of managing the inherently long sequences of tokens characteristic of video data.
-The model’s novelty lies in its efficient variants that factorize the spatial and temporal dimensions of the input, significantly enhancing computational efficiency and scalability.
-Moreover, ViViT pioneers in regularizing the model during training and ingeniously utilizes pretrained image models, enabling effective training on relatively smaller datasets—a notable advancement given the typically large datasets required by Transformer models.
-The paper’s comprehensive ablation studies and methodical analysis culminate in ViViT achieving state-of-the-art results on several prestigious video classification benchmarks, including Kinetics 400 and 600, Epic Kitchens, Something-Something v2, and Moments in Time, thereby setting new standards and outperforming existing methods based on deep 3D convolutional networks.
-Wang et al.
-[101] revolutionized the field of video action recognition by transitioning from conventional unimodal methodologies to a multimodal learning framework that capitalizes on the semantic nuances of label texts.
-This innovative approach, embodied in ActionCLIP, redefines the task as a video-text matching challenge, thereby facilitating zero-shot action recognition capabilities without necessitating additional labeled data or parameter modifications.
-Central to ActionCLIP is the ”pre-train, prompt, and fine-tune” paradigm, a strategy that effectively leverages extensive web image-text or video-text data for pre-training.
-This paradigm ingeniously aligns the action recognition task with pre-training challenges through prompt engineering and ensures robust performance through end-to-end fine-tuning on specific datasets.
-ActionCLIP, with its ViT-B/16 backbone, not only demonstrates exceptional zero-shot/few-shot transfer abilities but also achieves groundbreaking performance, notably securing a top-1 accuracy of 83.8% on the Kinetics-400 dataset.
-This paradigm shift underscores the potential of integrating rich semantic information from label texts into video representation, marking a significant leap forward in the domain of action recognition.
-Bertasius et al.
-[29] delves into the TimeSformer model, a novel approach to video understanding that leverages the Transformer architecture, renowned for its success in NLP, and adapts it for video data.
-The TimeSformer stands out by employing a divided space-time attention mechanism, allowing the model to separately and efficiently process the spatial and temporal dimensions of videos.
-This approach is meticulously detailed in the supplementary materials, which cover the model’s implementation, training, and inference procedures, alongside additional ablations and qualitative results.
-The TimeSformer is rigorously evaluated on large-scale video datasets such as Kinetics-400, Kinetics-600, Something-Something-V2, and Diving-48, demonstrating its robust performance and the model’s ability to focus on relevant parts of the video for nuanced spatiotemporal reasoning.
-The visualizations of space-time attention in the paper further illustrate the model’s capacity to understand and interpret complex video content, marking a significant advancement in the field of video understanding.
-Sharir et al.
-[102] introduced STAM (Space Time Attention Model), a groundbreaking approach to action recognition in videos.
-STAM addresses the computational inefficiencies prevalent in traditional methods that rely on 3D convolutions by significantly reducing the number of frames required for inference without compromising accuracy.
-The model capitalizes on a temporal Transformer that applies global attention over video frames, effectively capturing salient information and enabling input efficiency.
-This novel approach allows STAM to achieve state-of-the-art results on the Kinetics dataset, notably achieving 80.5% top-1 accuracy on Kinetics-400 with approximately 30% fewer frames per video and 40% faster inference compared to existing leading methods.
-The paper’s innovation lies in its unique application of self-attention layers for both spatial and temporal dimensions, marking a significant advancement in video action recognition by offering a solution that is not only accurate but also computationally efficient and practical for real-world applications.
-Fan et al.
-[103] introduced Multiscale ViTs (MViT), a novel Transformer architecture designed for video and image recognition tasks, effectively integrating the concept of multiscale feature hierarchies with Transformer models.
-MViT operates through several channel-resolution scale stages, progressively expanding channel capacity while reducing spatial resolution, thereby creating a multiscale pyramid of features.
-This design allows early layers to model simple low-level visual information at high spatial resolution, while deeper layers focus on spatially coarse but complex high-dimensional features, effectively addressing the dense nature of visual signals.
-The model demonstrates a significant performance gain over concurrent ViTs in video recognition tasks without relying on large-scale external pre-training, and it’s particularly efficient in computation and parameters.
-For instance, on the Kinetics-400 dataset, MViT achieves 80.5% top-1 accuracy with approximately 30% fewer frames per video and 40% faster inference than the current leading method.
-The approach’s novelty lies in its ability to leverage the Transformer architecture for dense visual signals, offering an efficient and scalable solution for video understanding tasks.
-The paper’s findings suggest that MViT can serve as a robust framework for various visual recognition tasks, setting new benchmarks in performance and efficiency.
-In 2022, a significant shift in video recognition modeling is presented by Ze liu et al.
-[104] , moving from conventional CNNs to Transformer-based architectures.
-the research presents the Video Swin Transformer, a novel adaptation of the Swin Transformer architecture tailored for video recognition tasks, leveraging the inherent spatiotemporal locality of videos.
-Unlike traditional Transformer models that compute self-attention globally, the Video Swin Transformer employs local self-attention within non-overlapping windows across both spatial and temporal dimensions, significantly reducing computational overhead and model size.
-This approach not only capitalizes on the local nature of video data but also maintains the hierarchical structure of the original Swin Transformer, extending local attention to the spatiotemporal domain.
-The model demonstrates exceptional performance on various video recognition benchmarks, including Kinetics-400, Kinetics-600, and Something-Something v2, achieving state-of-the-art accuracy with a smaller model size and less pre-training data compared to its predecessors.
-The Video Swin Transformer’s innovative design, focusing on local spatiotemporal relationships and efficient computation, marks a substantial advancement in the field of video understanding, offering a robust and scalable solution for video-based recognition tasks.
-Jiawei Chen et al.
-[105] introduced a pioneering Transformer-based approach, MM-ViT, for video action recognition, distinctively operating within the compressed video domain.
-This approach capitalizes on the rich, readily available multi-modal data in videos, including I-frames, motion vectors, residuals, and audio waveforms, diverging from traditional methods that primarily depend on decoded RGB frames.
-MM-ViT’s architecture is meticulously designed to manage the extensive spatiotemporal tokens from these modalities, employing scalable model variants and innovative cross-modal attention mechanisms to factorize self-attention across space, time, and modality dimensions efficiently.
-The model’s efficacy is rigorously validated on prominent action recognition benchmarks like UCF-101, Something-Something-v2, and Kinetics-600, where it demonstrates superior performance.
-Notably, MM-ViT not only surpasses existing video Transformers in efficiency and accuracy but also competes favorably with state-of-the-art CNN models that rely on computationally intensive optical flow, marking a significant advancement in the domain of video action recognition.
-Hashiguchi et al.
-[106] introduced a novel approach to enhance action recognition in videos through the innovative Multi-head Self/Cross-Attention (MSCA) mechanism, integrated within the ViT framework.
-This research addresses the limitations of previous methods like Temporal Shift Module (TSM) and TokenShift by fully leveraging the multi-head self-attention structure of ViT for temporal feature interaction.
-MSCA innovatively allows attention heads to focus not only on the current frame but also on adjacent frames, enriching the temporal context of the action recognition process.
-Evaluated on the Kinetics400 dataset, the MSCA mechanism, particularly the MSCA-KV variant, demonstrates superior performance, outperforming the baseline ViT and TokenShift models.
-This improvement underscores the model’s ability to effectively capture the dynamic temporal relationships in videos, marking a significant advancement in the field of action recognition.
-The paper’s contribution lies in its novel approach to temporal interaction within the Transformer block, achieving enhanced recognition performance without increasing computational complexity or altering the underlying model architecture.
-Momal Ijaz et al.
-[107] introduced an innovative multimodal Transformer-based network aimed at improving elderly patient care in healthcare facilities.
-This network adeptly fuses features from skeletal joints and acceleration data to recognize nursing activities, addressing the critical need for automated and accurate monitoring of nurse care.
-The model’s distinctiveness lies in its dual modality approach, combining acceleration and skeletal joints data, a method not previously explored in the Nurse Care Activity Recognition Challenge (NCRC) dataset.
-The proposed network not only outperforms traditional single modality networks but also sets a new benchmark by surpassing state-of-the-art methods like ST-GCNs, GRUs, and classifiers based on handcrafted features, achieving an impressive accuracy of 81.8%.
-The research underscores the potential of Transformer-based models in healthcare, particularly in creating a fair and efficient care environment by ensuring consistent patient care and alleviating the documentation burden on nursing staff.
-In 2023, Wensel et al.
-[108] introduced a groundbreaking ViT-ReT framework, integrating ViT and Recurrent Transformer (ReT) to address the challenges in HAR in videos.
-This innovative approach aims to overcome the limitations of traditional activity recognition systems that heavily rely on CNNs for feature extraction and RNNs for analyzing time-dependent data.
-The ViT-ReT framework replaces CNNs with ViT for efficient and salient feature extraction from images and supplants RNNs with ReT to effectively handle sequences of data, addressing the high model complexity and lack of parallelizability inherent in traditional models.
-The framework is designed to be lightweight and efficient, making it particularly suitable for resource-constrained and real-time environments like IoT and edge devices.
-Extensive comparative experiments across four publicly available human action datasets demonstrate that ViT-ReT not only achieves a significant speedup over the baseline ResNet50-LSTM approach but also outperforms state-of-the-art HAR methods in both accuracy and runtime, marking a significant advancement in the field of computer vision and activity recognition.
-Chen et al.
-[109] presented a novel approach to HAR from still images, a notably challenging task in computer vision.
-The Swin-Fusion model, at the core of this research, leverages the Swin-Transformer’s hierarchical structure and introduces a feature fusion module to effectively extract and integrate multi-scale behavioral information.
-This innovative model addresses the inherent limitations of CNNs, such as the lack of long-range perception and difficulty in understanding global structures, by utilizing the Transformer’s ability to focus on discriminative features and calculate relationships between distant pixels.
-Evaluated on five distinct still image-based HAR datasets, the Swin-Fusion model demonstrates superior performance, outperforming advanced CNN-based models.
-It achieves this by enhancing behavior recognition through the strategic sharing and reuse of feature maps at various stages, without significantly increasing the computational resources required for training.
-The research marks a significant advancement in still image-based HAR, offering a robust solution that combines the strengths of Swin-Transformer architecture and feature fusion techniques.
-Xing et al.
-[110] introduces a novel approach to semi-supervised action recognition by leveraging Transformer models, a departure from the conventional use of CNNs in this domain.
-Addressing the challenge of efficiently utilizing the abundance of unlabeled video data for action recognition, the authors propose SVFormer, a model that incorporates a steady pseudo-labeling mechanism (EMA-Teacher) and introduces innovative data augmentation strategies tailored for video data: Tube TokenMix and Temporal Warping Augmentation.
-These augmentations are designed to capture the complex temporal dynamics inherent in video content, a critical aspect that has been challenging for existing semi-supervised learning (SSL) methods.
-Evaluated on three benchmark datasets—Kinetics-400, UCF-101, and HMDB-51—SVFormer demonstrates a significant improvement in action recognition performance, outperforming state-of-the-art methods by a notable margin, especially under conditions of limited labeled data.
-This breakthrough underscores the potential of Transformers in video understanding tasks and sets a new precedent for future research in semi-supervised learning for video action recognition, highlighting the effectiveness of SVFormer’s methodological innovations in leveraging unlabeled video data.
-Mondal et al.
-[111] introduced an approach to actor-agnostic, multi-modal, multi-label action recognition that significantly diverges from traditional actor-specific models.
-By ingeniously integrating visual and textual modalities through a Transformer-based detection framework inspired by DETR, and leveraging the capabilities of a pretrained vision-language model like CLIP, MSQNet is designed to recognize a multitude of actions within a single video across different actors, including both humans and animals.
-This integration not only facilitates a richer representation of action classes but also obviates the necessity for actor-specific pose estimation, thereby streamlining the model’s complexity and reducing its maintenance demands.
-The novelty of MSQNet lies in its ability to address the concurrent occurrence of multiple actions and utilize additional information sources such as class name text, which have been largely overlooked in existing methods.
-Demonstrated through extensive experimentation across five publicly available benchmarks, MSQNet’s superior performance, achieving up to a 50% improvement over prior arts, underscores its potential to set a new standard in the field of action recognition by offering a versatile, efficient, and more generalized solution.
-Table 2: Transformer-based Approaches.
-Paper Year Model Dataset Novelty
-Arnab et al.
-[30] 2021 Pure-Transformer architecture Kinetics 400 and 600, Epic Kitchens 100, Something-Something v2 and Moments in Time The novelty lies in its introduction of a Transformer based model for video classification, efficiently handling spatio-temporal tokens through factorized spatial and temporal dimensions, and achieving state-of-the-art results on multiple benchmarks
-Wang et al.
-[101] 2021 A ViT-based “ActionClip” architecture Kinetics-400 The novelty lies in its multimodal learning framework that transforms video HAR into a video-text matching problem, enabling robust zero-shot recognition and introducing the ”pre-train, prompt, and fine-tune” paradigm for leveraging large-scale web data
-Bertasius et al.
-[29] 2021 ViT-based “TimeSformer” architecture Kinetics-400, Kinetics-600, Something-Something-V2 and Diving-48 The novelty lies in its innovative application of divided space-time attention, enabling efficient and effective video understanding by separately processing spatial and temporal components within the Transformer architecture
-Sharir et al.
-[102] 2021 Transformer architecture Kinetics-400, Kinetics-600, Something-Something-V2 and Diving-48 The novelty lies in introducing STAM, a Transformer-based model for video HAR that leverages global attention over video frames, significantly reducing the required number of frames and computational load while maintaining or surpassing state-of-the-art accuracy
-Fan et al.
-[103] 2021 ViT-based “MViT” architecture Kinetics-400, Kinetics-600, Something-Something-V2, Charadesand AVA The novelty lies in its integration of multiscale feature hierarchies with Transformer models, creating a scalable and efficient architecture for advanced video and image recognition tasks
-architecture Ze liu et al.
-[104] 2022 ViT-based architecture “Video Swin Transformer” Kinetics-400, Kinetics-600 and Something-Something v.
-The novelty lies in its hierarchical, local self-attention mechanism within non-overlapping windows across both spatial and temporal dimensions
-Jiawei Chen et al.
-[105] 2022 Transformer-based “MM-ViT” model UCF-101, Something-Something-v2 and Kinetics-600 The novelty lies in its unique approach of operating exclusively in the compressed video domain, leveraging a pure Transformer-based model to exploit all available modalities (I-frames, motion vectors, residuals, and audio) for enhanced video action recognition
-Hashiguchi et al.
-[106] 2022 ViT-based “MSCA” architecture Kinetics400 The novelty is a mechanism within the ViT, enabling enhanced temporal interaction across frames for HAR without increasing computational complexity
-Momal Ijaz et al.
-[107] 2022 Pure Transformer-based architecture Nurse Care Activity Recognition Challenge (NCRC) dataset The novelty of the study lies in its development of a multimodal Transformer network that uniquely fuses skeletal joints and acceleration data for enhanced nursing activity recognition
-Wensel et al.
-[108] 2023 Integrated ViT and Recurrent Transformer (ReT) - The novelty lies in the ViT-ReT framework, which innovatively combines ViT and ReT to enhance speed and accuracy in HAR in videos, particularly in resource-constrained and real-time environments
-Chen et al.
-[109] 2023 ViT-based architecture Li’s action, Stanford-40, PPMI-24, AUC-V1, and AUC-V2 datasets The novelty lies in the Swin-Fusion model, which innovatively combines the Swin-Transformer with a feature fusion module for enhanced HAR from still images, surpassing the performance of traditional CNN-based models
-Xing et al.
-[110] 2023 ViT architecture Kinetics-400, UCF-101, and HMDB-51 The novelty lies in its pioneering use of Transformers for semi-supervised video recognition, enhanced by data augmentation strategies specifically designed for video data: Tube TokenMix, Temporal Warping Augmentation
-3.3 CNN-Transformer Hybrid Architectures
-In 2021, Neimark et al.
-[112] presented a transformative approach to video recognition by introducing a Transformer-based framework that diverges from the traditional reliance on 3D ConvNets.
-Video Transformer Networks (VTN) capitalizes on the strengths of Transformers to process entire video sequences, employing a modular architecture that includes a 2D spatial backbone for feature extraction, a temporal attention-based encoder leveraging Longformer for integrating temporal dynamics, and a classification MLP head for action prediction.
-This innovative framework demonstrates remarkable efficiency, achieving training speeds 16.1 times faster and inference speeds 5.1 times faster than existing state-of-the-art methods, while requiring significantly fewer GFLOPs.
-Evaluated on the Kinetics-400 and Moments in Time benchmarks, VTN not only maintains competitive accuracy but also offers a scalable solution for analyzing long video sequences in an end-to-end manner.
-The paper’s findings suggest that VTN could serve as a new baseline for video recognition research, highlighting the potential of Transformer models in efficiently handling complex video data and paving the way for future advancements in the field.
-In 2022, Liu et al.
-[113] presented TadTR, an innovative end-to-end framework for Temporal Action Detection (TAD) utilizing Transformer technology, addressing the complexity and rigidity of previous TAD methods that rely on multi-stage pipelines and hand-designed components.
-TadTR simplifies the TAD process by introducing action queries, learnable embeddings that adaptively extract relevant temporal context from videos for direct action instance prediction.
-This approach is enhanced with three key innovations: a temporal deformable attention module for focused context extraction, a segment refinement mechanism for accurate boundary detection, and an actionness regression head for confident instance prediction.
-Demonstrating its effectiveness, TadTR achieves state-of-the-art performance on benchmark datasets like THUMOS14 and HACS Segments, and competitive results on ActivityNet-1.3 when combined with an additional action classifier.
-This streamlined, efficient framework marks a significant advancement in video understanding, offering a flexible, powerful solution for TAD with lower computational costs and improved end-to-end learning capabilities.
-In 2023, Dasom Ahn et al.
-[114] presented an innovative approach to HAR by integrating spatio-temporal video and skeleton data into a cohesive model, addressing the complexity and inefficiency of handling cross-modal data with separate models.
-This method innovatively aggregates video frames and skeleton sequences into global grid tokens and joint map tokens, respectively, which are then processed by the STAR-Transformer.
-The core of the STAR-Transformer includes a unique encoder-decoder structure with full spatio-temporal attention, zigzag spatio-temporal attention, and binary spatio-temporal attention modules, designed to enhance the model’s ability to capture and represent multi-feature spatio-temporal data effectively.
-Evaluated on benchmark datasets such as Penn-Action, NTU-RGB+D 60, and NTU-RGB+D 120, the STAR-Transformer demonstrates superior performance over existing state-of-the-art methods, showcasing its potential to significantly advance the field of action recognition.
-This approach not only simplifies the pipeline for processing cross-modal data but also offers a more efficient and flexible solution for action recognition tasks, highlighting the effectiveness of integrating video and skeleton features through a Transformer-based model.
-Surek et al.
-[115] introduces an innovative approach by combining a residual network (ResNet) and a ViT architecture with a semi-supervised learning method, using DINO (self-DIstillation with NO labels).
-This novel methodology is tested using the Human Motion Database (HMDB51), chosen for its diverse range of human actions.
-The study’s results are particularly promising, with the bi-dimensional ViT combined with long short-term memory (LSTM) demonstrating high accuracy rates.
-These findings indicate a substantial improvement in the field of video-based HAR, showcasing the effectiveness of integrating advanced deep learning techniques.
-The research’s approach to HAR, utilizing a blend of supervised and semi-supervised learning models, is a noteworthy contribution to the domain.
-Wang et al.
-[116] focused on developing an efficient model capable of operating on edge devices with limited computing power.
-The core innovation is the 3D-ShuffleViT network, which ingeniously integrates the self-attention mechanism with convolution, incorporating the ACISA module to enhance performance.
-This model was rigorously evaluated on various datasets, including EgoGesture, Jester, and UCF-101, covering a broad spectrum of action recognition scenarios.
-The 3D-ShuffleViT network stands out for its computational efficiency, achieving a significant reduction in computational cost (only 6% of that of SlowFast-ResNet101) while maintaining high accuracy (98% of SlowFast-ResNet101’s Top1 accuracy on the EgoGesture dataset).
-Additionally, it offers a remarkable speed advantage, being 2.5 times faster than SlowFast-ResNet101 on the same CPU.
-The 3D-ShuffleViT network’s ability to balance accuracy and computational efficiency makes it an ideal solution for deployment on edge devices.
-Its performance on various datasets demonstrates its versatility and effectiveness in different action recognition contexts.
-The model’s innovative design, combining the strengths of self-attention and convolution, represents a significant step forward in the development of lightweight yet powerful video action recognition systems.
-Djenouri et al.
-[117] introduced the CVTN (Convolution Visual Transformer Network), a novel hybrid model that synergizes CNNs and Visual Transformers (VTs) to enhance the analysis and comprehension of human activities.
-The model is designed to leverage sensor data in both spatial and temporal dimensions, with the CNN component learning spatial visual features from diverse sensor data, followed by the VT segment capturing temporal insights.
-This integration aims to provide a more robust and comprehensive understanding of human activities.
-The efficacy of the CVTN model is evaluated using the Kinetics dataset, which simulates real-world HAR scenarios.
-The results of this evaluation are significant, with the CVTN model demonstrating a remarkable performance, achieving a recognition rate of 95%, thereby outperforming recent baseline HAR solutions.
-The CVTN model’s approach to combining spatial and temporal data through a hybrid of CNNs and VTs represents a significant advancement in HAR.
-Ye et al.
-[118] introduced a transformative approach to video prediction tasks such as future frame prediction and frame interpolation using a single, unified model.
-This model employs a transformer-based neural process combined with implicit neural representations (INRs), enabling continuous and task-agnostic prediction across various datasets including KTH, BAIR, KITTI, and Cityscapes.
-Distinctively, the model operates beyond traditional discrete prediction frameworks by leveraging INRs to handle spatio-temporal coordinates, which allows it to predict video frames at any desired frame rate, a significant advancement over existing task-specific models.
-The findings highlight the model’s capability to generate high-quality, temporally consistent videos across different scenarios, achieving state-of-the-art performance on standard metrics and demonstrating substantial improvements in video prediction technology.
-This research suggests a scalable and efficient pathway for future video processing applications, where a single model can adapt to multiple video prediction tasks without the need for task-specific adjustments.
-Ahmadabadi et al.
-[119] delves into the fusion of CNNs and ViT models to advance the field of HAR.
-This innovative research introduces a methodology where knowledge distillation is employed to transfer insights from a comprehensive, high-capacity CNN model (teacher) to a more compact ViT model (student), aiming to amalgamate the teacher’s adeptness at extracting local image features with the student’s capability to apprehend global features through attention mechanisms.
-Utilizing the Stanford 40 dataset for empirical validation, the study showcases a significant enhancement in the performance of the student models, evidenced by improved accuracy and mean Average Precision (mAP) metrics compared to conventional training approaches.
-The novelty lies in its strategic approach to integrating the distinct advantages of CNNs and ViTs, leveraging the depth and breadth of feature recognition capabilities across different models, thereby setting a new precedent for efficiency and effectiveness in the recognition of human actions from still images.
-This approach not only underscores the potential of hybrid model architectures in overcoming the limitations of individual models but also highlights the efficacy of knowledge distillation as a pivotal technique for optimizing model performance in complex vision tasks.
-Table 3: Hybrid Approaches.
-Paper Year Model Datase Novelty
-Neimark et al.
-[112] 2021 “VTN” a hybrid CNN-transformer Kinetics-400 and Moments in Time The novelty of the lies in its Transformer-based framework for video recognition that processes entire video sequences efficiently, moving beyond traditional 3D ConvNets and enabling faster training and inference with competitive accuracy
-Liu et al.
-[113] 2022 “TadTR” a hybrid CNN-transformer THUMOS14, HACS Segments and ActivityNet-1.3 TadTR introduces a novel end-to-end Transformer-based framework for Temporal Action Detection, simplifying the detection pipeline with action queries and incorporating unique enhancements like temporal deformable attention, segment refinement, and actionness regression for efficient and accurate action instance prediction
-Dasom Ahn et al.
-[114] 2023 “STAR-Transformer” a hybrid CNN-Transformer model Penn-Action, NTU-RGB+D 60, and NTU-RGB+D 120 datasets The novelty lies in the unique integration of spatio-temporal video and skeleton data into a cohesive Transformer-based model, employing specialized attention mechanisms to efficiently process cross-modal data for HAR
-Surek et al.
-[115] 2023 A hybrid CNN-Transformer model HMDB51 The novelty of the study lies in its integration of ResNet and ViT with DINO (self-DIstillation with NO labels) for semi-supervised learning in HAR from RGB videos
-Wang et al.
-[116] 2023 “3D-ShuffleViT” a hybrid CNN-transformer architecture EgoGesture, Jester and UCF-101 The novelty lies in the deep integration of self-attention and convolution within a 3D framework, achieving high video HAR accuracy with significantly reduced computational costs for deployment on edge devices
-Djenouri et al.
-[117] 2023 A hybrid CNN-transformer architecture Kinetics dataset The novelty of the study lies in the CVTN model’s unique hybrid integration of CNNs and ViTs to enhance HAR by leveraging both spatial and temporal sensor data
-Ye et al.
-[118] 2023 A hybrid CNN-transformer architecture KTH, BAIR, KITTI and Cityscapes The novelty of the paper lies in its development of a unified, transformer-based neural process model that enables continuous conditional video prediction across multiple tasks using implicit neural representations.
-Ahmadabadi et al.
-[119] 2023 A hybrid CNN-transformer architecture Stanford 40 dataset The novelty lies in enhancing HAR by distilling knowledge from high-capacity CNNs to compact ViTs, merging local and global image feature extraction for improved accuracy and efficiency
+Recent advancements in action recognition have seen a significant shift towards ViT, highlighting their efficacy in video understanding tasks.
+[4] introduced ViViT, extending the vision Transformer architecture to handle video sequences.
+They demonstrated its potential on datasets like Kinetics-400 and Something-Something-V2, marking a substantial improvement in video action recognition capabilities.
+Building on this, [10] proposed a space-time Transformer that models temporal information innovatively, and achieved competitive results on similar datasets.
+The efficiency of multiscale ViTs was further illustrated by [30], who showed that such architectures could effectively capture fine-grained video details and enhance classification performance on comprehensive video datasets.
+Moreover, [68] presented the Swin Transformer, utilizing a shifted window mechanism to model long-range dependencies more efficiently, and leading to significant improvements in action recognition accuracy.
+Together, these works underscore the transformative impact of ViTs in advancing the field of HAR.
+Additionally, [102] introduced ActionCLIP, leveraging the CLIP model for enhanced video action recognition on multiple standard video datasets, including Kinetics-400 and HMDB-51.
+This novel approach integrated visual and linguistic representations.
+[16] introduced Mm-ViT, a multi-modal video Transformer designed for compressed video action recognition, and demonstrated high performance by leveraging multi-modal inputs on compressed video datasets such as HACS and UCF101.
+[80] explored the extension of ViT to video data, showing its potential in capturing temporal dynamics effectively across several standard video datasets including Kinetics-400 and HMDB-51.
+Furthermore, [108] developed SVFormer, a semi-supervised video Transformer that leverages both labeled and unlabeled data to bridge the gap between supervised and unsupervised learning, and achieved significant improvements in action recognition tasks on various standard HAR datasets such as Kinetics-400 and UCF101.
+Together, these works underscore the transformative impact of ViTs in advancing the field of HAR.
+Table LABEL:table:TransformerbasedApproaches1 presents the works discussed in this section that utilized ViTs.
+Table 4: ViT-based approaches in HARs.
+Paper Model Dataset Novelty
+[4] ViViT Kinetics-400, Something-Something-V2 Extended ViT to video sequences.
+[10] Space-Time Transformer Kinetics-400 Innovative temporal information modeling.
+[30] Multiscale ViT Kinetics-400, Something-Something-V2 Efficient capture of fine-grained video details.
+[68] Swin Transformer Kinetics-400, Something-Something-V2 Shifted window mechanism for long-range dependency modeling.
+[102] ActionCLIP Kinetics-400, HMDB-51 Leveraged CLIP for enhanced video action recognition.
+[16] Mm-ViT HACS, UCF101 Multi-modal inputs for compressed video action recognition.
+[80] ViT Kinetics-400, HMDB-51 Applied ViT to video data.
+[108] SVFormer Kinetics-400, UCF101 Semi-supervised learning for action recognition.
+3.3 CNN-ViT Hybrid Architectures
+The integration of ViTs with CNNs has significantly advanced HAR tasks.
+[112] proposed a two-stream hybrid CNN-Transformer network (THCT-Net), which demonstrated enhanced generalization ability and convergence speed on the NTU RGB+D dataset by combining CNNs for low-level context sensitivity and Transformers for capturing global information.
+Following this, [48] applied a similar hybrid model to driver action recognition, leveraging multi-view data to achieve high accuracy through the integration of CNNs for spatial feature extraction and Transformers for temporal dependencies.
+[52] extended this approach by integrating 3D CNNs with Transformers for late temporal modeling, and achieved substantial improvements in action recognition accuracy on the HMDB-51 and UCF101 datasets.
+Moreover, [110] proposed Swin-Fusion, which combines Swin Transformers with CNN-based feature fusion to achieve state-of-the-art performance on datasets like Kinetics-400 and Something-Something-V2, demonstrating the robustness and superior performance of hybrid models in HAR tasks.
+[27] proposed a hybrid visual Transformer model that integrates CNNs and Transformers for efficient and accurate human activity recognition.
+They demonstrated its capability on datasets like Kinetics-400 and UCF101, and showed that the hybrid approach leverages the local feature extraction of CNNs with the global context modeling of Transformers.
+Following this, [86] provided a comprehensive review of deep learning approaches for video-based human activity recognition, emphasizing the potential of hybrid models.
+This review underscored the effectiveness of such hybrid models in capturing both spatial and temporal features from video data, and evaluated on various human activity datasets including NTU RGB+D and UTD-MHAD.
+[2] explored the use of knowledge distillation techniques to enhance the performance of hybrid CNN-Transformer models.
+Their approach was validated on datasets such as HMDB-51 and Kinetics-400, showing significant improvements in HAR by effectively transferring knowledge from complex models to more efficient ones.
+Together, these works highlight the evolving landscape of hybrid models in human activity recognition, showcasing their robustness and efficiency in handling complex video data.
+Table LABEL:table:3DCNNBased presents the works discussed in this section that utilized CNN-ViT approaches.
+Table 5: CNN-ViT hybrid approaches in HARs.
+Paper Model Datase Novelty
+[112] The two-stream hybrid CNN-Transformer network (THCT-Net) NTU RGB+D Combined CNNs and Transformers for improved generalization and convergence speed.
+[48] Multi-view vision Transformer Custom driver action datasets Leveraged multi-view data for spatial and temporal feature integration.
+[52] 3D CNN-Transformer HMDB-51, UCF101 Integrated 3D CNNs with Transformers for late temporal modeling.
+[110] Swin-Fusion Kinetics-400, Something-Something-V2 Combined Swin Transformers with CNN-based feature fusion for state-of-the-art performance
+[27] Hybrid visual Transformer Kinetics-400, UCF101 Efficient and accurate human activity recognition leveraging strengths of CNNs and Transformers
+[86] Various deep learning models including hybrid models NTU RGB+D, UTD-MHAD Comprehensive review highlighting the potential of hybrid models.
+[2] Hybrid CNN-Transformer HMDB-51, Kinetics-400 Knowledge distillation from CNN-Transformer models for enhanced performance.
 3.4 Discussion
-In the field of action recognition, the choice of model – whether CNN-based, ViT-based, or a hybrid of CNN and ViT – significantly influences the outcome and efficiency of the task.
+In the field of HAR, the choice of models – whether CNN-based, ViT-based, or a hybrid of CNN and ViT – significantly influences the outcome and efficiency of the task.
 CNN-based models are particularly adept at extracting local features due to their convolutional nature [60], making them highly effective in pattern recognition within images and videos.
@@
 For a summary of the advantages and disadvantages of these three architectural variations, see Table 6.
-Table 4: Capability Comparison of Transformer-based, CNN-based, and Hybrid Models in Action Recognition
+Table 6: Capability comparison between Transformer-based, CNN-based, and hybrid models in HARs.
 Criteria ViT-based CNN-based Hybrid Models
@@
 4 Proposed CNN-ViT Hybrid Architecture
-In this section, we will detail our proposed CNN-ViT architecture for human action recognition, which leverages the benefits of both approaches described above.
+In this section, we present our proposed CNN-ViT architecture for HAR, leveraging the benefits of both approaches described in previous sections, see Figure 7.
 The architecture incorporates a TimeDistributed layer with a CNN backbone, followed by a ViT model to classify actions in video sequences.
-Spatial Component (TimeDistributed CNN): The CNN backbone (i.e.
-Mobilenet [121]) in TimeDistributed layer processes each frame $\boldsymbol{X}\in\mathbb{R}^{K\times M\times N}$, and outputs spatial features vector $\boldsymbol{v}=p_{\theta}(\boldsymbol{X})\in\mathbb{R}^{L}$ where $p_{\theta}$ is the CNN model (e.g.
-MobileNet or VGG16) with parameters $\theta$ wrapped by TimeDistributed layer.
-Let $\mathcal{X}$ be a collection of $n$ frames $\mathcal{X}=\{\boldsymbol{X}^{i}\}_{i=1}^{n}$, which is fed into the Spatial component of our architecture in Figure 7.
-The trained CNN $p_{\theta}$ is then used $n$ times to process $\mathcal{X}$ frame by frame and create a set of $n$ spatial feature vectors $\mathcal{V}=\{\boldsymbol{v}^{i}\}_{i=1}^{n}$, where $\boldsymbol{v}^{i}=p_{\theta}(\boldsymbol{X}^{i})$.
-Temporal Component (ViT): In this proposed hybrid CNN-ViT model, the Vit is engineered to process sequences of $N$ vectors, see Figure 7, each representing a distinct frame of the video, but in the proposed ViT-only model, each vector is representing a distinct patch, see Figure 8.
+Spatial component.
+Let $\mathcal{X}$ be a collection of $N$ frames, i.e., $\mathcal{X}=\{\boldsymbol{X}_{i}\}_{i=1}^{N}$.
+The CNN backbone (i.e.
+MobileNet in Howard et al.
+45) in the TimeDistributed layer (see Figure 7) processes the indifivual frames $\boldsymbol{X}_{i}$ and outputs the spatial features vector $\boldsymbol{v}_{i}=p_{\theta}(\boldsymbol{X}_{i})\in\mathbb{R}^{L}$, where $p_{\theta}$ is the CNN model (e.g.
+MobileNet or VGG16) with parameters in $\theta$ wrapped by the TimeDistributed layer.
+Temporal component.
+In the proposed hybrid CNN-ViT model, ViT is engineered to process the sequence of the $N$ spatial features vectors, i.e., $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$, where each $\boldsymbol{v}_{i}$ represents a distinct frame of the input video clip, see Figure 7.
+Afterwards, the ViT block outputs a final representation $\boldsymbol{z}$, which is then fed into the softmax layer to classify the action in the video.
+In detail, the Transformer encoder is designed to process a sequence of vectors, each representing one frame, and aggregate information into a single vector for classification.
+In the proposed ViT-only model in Figure 8 for the purpose of comparison, each vector represents a distinct patch.
 These vectors are first linearly projected into a high-dimensional space, facilitating the model’s ability to learn complex patterns within the data.
@@
 The output vector associated with this [CLS] token, after propagation through the Transformer layers, serves as a comprehensive representation of the entire input, suitable for downstream classification tasks.
-The sequence of vectors $\mathcal{V}$ which is outputted by the TimeDistributed CNN is then fed into the ViT.
-The ViT processes this sequence and outputs a final representation $\boldsymbol{z}$ which is represented as
-$\boldsymbol{z}=\text{ViT}([\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{N}]).$ (16)
-The final representation $\boldsymbol{z}$ is then fed into the Softmax layer to classify the action in the video.
-The Transformer encoder is designed to process a sequence of vectors, each representing one frame, and aggregate information into a single vector for classification.
 […]
-Figure 7: The hybrid CNN-ViT architecture.
+Figure 7: The hybrid CNN-ViT architecture for HARs.
 […]
-Figure 8: The ViT-only architecture.
-5 Experiments
-The goal of the presented experiments is not necessarily to produce a model that outperforms the state-of-the-art models in HAR field.
-Rather, the aim is to conduct a comparison among the ViT-only, CNN, and hybrid models to give further insights.
-The KTH dataset was chosen for its balanced representation of spatial and temporal features.
-Six experiments were conducted, training each of the aforementioned models on three different lengths of frame sequences.
-Care was taken to avoid pretraining in order to ensure the neutrality of the results.
-The Transnet model by Alomar et al.
-[126], was used to represent the CNN model.
-The ViT model was depicted in Figure 8.
-For the Spatial Component of the hybrid model, we employed the spatial component of Transnet and for the Temporal Component, we employed the same ViT model that we used in ViT-only model.
-Experiment Settings.
-We constructed our model utilizing Python 3.6, incorporating the Keras deep learning framework, OpenCV for image processing, matplotlib, and the scikit-learn library.
-The training and testing were performed on a computer equipped with an Intel Core i7 processor, an NVidia RTX 2070 graphics card, and 64GB of RAM.
-Data.
-In 2004, the Royal Institute of Technology unveiled the KTH dataset, a significant and publicly accessible dataset for action recognition [127].
+Figure 8: The ViT-only architecture for HARs.
+4.1 Experiments
+The goal of the presented experiments is not necessarily to produce a model that outperforms the state-of-the-art models in the HAR field.
+Rather, the aim is to conduct a comparison among the CNN, ViT-only, and hybrid models to give further insights.
+The Royal Institute of Technology in 2004 unveiled the KTH dataset, a significant and publicly accessible dataset for action recognition [79].
+The KTH dataset was chosen here for its balanced representation of spatial and temporal features.
 Renowned as a benchmark dataset, it encompasses six types of actions: walking, jogging, running, boxing, hand-waving, and hand-clapping.
-The dataset features performances by twenty-five different individuals, introducing a diversity in execution.
+The dataset features performances by 25 different individuals, introducing a diversity in execution.
 Additionally, the environment for each participant’s actions was deliberately altered, including settings such as outdoors, outdoors with scale changes, outdoors with clothing variations, and indoors.
 The KTH dataset comprises 2,391 video sequences, all recorded at 25 frames per second using a stationary camera against uniform backgrounds.
-5.1 Results and Discussion
-In the exploration of HAR within video sequences, three distinct models were evaluated i.e., CNN, ViT-only, and a Hybrid model.
-These models were trained on the KTH dataset, focusing on three different context lengths i.e., short (12 frames), medium (18 frames) and long (24 frames).
-The results from these experiments provide insightful revelations into the efficacy of each model under different temporal contexts, see Table 5.
-Table 5: Experimental Results of three models on the KTH Dataset using three different context lengths.
-Hybrid-N refers to the model trained without pre-training whereas Hybrid-I is for the model pretrained on Imagenet[34].
-Context length CNN-based ViT-only Hybrid-N Hybrid-I
+Six experiments were conducted, with each of the aforementioned models trained on three different lengths of frame sequences.
+Care was taken to avoid pre-training in order to ensure the neutrality of the results.
+The TransNet model by [3] was adopted to represent the CNN model, and the ViT model was depicted in Figure 8.
+For the spatial component of the hybrid model, we employed the spatial component of TransNet; and for the temporal component, we employed the same ViT model that we used in the ViT-only model.
+We constructed our model utilizing Python 3.6, incorporating the Keras deep learning framework, OpenCV for image processing, matplotlib, and the scikit-learn library.
+The training and test were performed on a computer equipped with an Intel Core i7 processor, an NVidia RTX 2070 graphics card, and 64GB of RAM.
+4.1.1 Results and Discussion
+Table 7: Experimental results of different models on the KTH Dataset using three different context lengths.
+In particular, the hybrid model was trained without pre-training whereas Hybridpre is for the hybrid model pre-trained on ImageNet.
+Context length CNN-based ViT-only Hybrid Hybridpre
 12 frames 94.35 92.44 94.12 96.34
@@
 24 frames 93.49 93.69 95.78 97.89
+Table 7 presents the quantitative results of the three distinct models, i.e., CNN, ViT-only, and a hybrid model on the KTH dataset, focusing on three different context lengths, i.e., short (12 frames), medium (18 frames), and long (24 frames).
+The results from these experiments provide insightful revelations into the efficacy of each model under different temporal contexts.
+More details are given below.
 The CNN model exhibited a decrease in accuracy as the frame length increased, recording 94.35% for 12 frames, 93.91% for 18 frames, and 93.49% for 24 frames.
 This descending trend suggests that CNN may struggle with processing longer sequences where temporal dynamics become more complex, potentially leading to challenges such as overfitting or difficulties in temporal feature retention over extended durations.
-In contrast, the ViT model demonstrated an improvement in performance with longer sequences, achieving accuracies of 92.44% for 12 frames, 92.82% for 18 frames, and 93.69% for 24 frames.
+In contrast, the ViT model demonstrated an improvement in performance with longer sequences, achieving accuracy of 92.44% for 12 frames, 92.82% for 18 frames, and 93.69% for 24 frames.
 This ascending pattern supports the notion that ViT architectures, with their inherent self-attention mechanisms, are well-suited to managing longer sequences.
@@
 The hybrid CNN-ViT model showcased the highest and continuously improving accuracy rates across all frame lengths: 94.12% for 12 frames, 94.56% for 18 frames, and an impressive 95.78% for 24 frames.
-This model synergistically combines CNN’s robust spatial feature extraction capabilities with ViT’s efficient handling of temporal relationships via self-attention.
+Moreover, the pre-trained hybrid model showcased the same trend, with the best accuracy achieved.
+This type of model synergistically combines CNN’s robust spatial feature extraction capabilities with ViT’s efficient handling of temporal relationships via self-attention.
 The results from this model indicate that such a hybrid approach is particularly effective in capturing the complexities of action recognition tasks in video sequences, especially as the sequence length increases.
@@
 Additionally, further research might investigate the optimal balance of CNN and ViT components within hybrid models to maximize both performance and efficiency.
-Table 6: Comparison of proposed hybrid model with the state-of-the-art models on KTH dataset.
+Table 8: Comparison of the proposed hybrid model with the state-of-the-art models on the KTH dataset.
 Methods Venue Accuracy
-Geng et al. [59] ICCSAE ’15 92.49
-Arunnehru et al. [73] RoSMa ’18 94.90
-Abdelbakya et al. [86] ITCE ’20 87.52
-Jaouedi et al. [87] KSUCI journal ’20 96.30
-Liu X et al [83] JAIHC ’20 91.93
-Sahoo et al. [85] TETCI ’20 97.67
-Lee et al. [91] CVF ’21 89.40
-Basha et al. [96] MTA journal ’22 96.53
-Ye et al. [118] CVF ’23 90.90
+[35] ICCSAE ’16 92.49
+[5] RoSMa ’18 94.90
+[1] ITCE ’20 87.52
+[47] KSUCI journal ’20 96.30
+[67] JAIHC ’20 91.93
+[78] TETCI ’20 97.67
+[62] CVF ’21 89.40
+[8] MTA journal ’22 96.53
+[109] CVF ’23 90.90
 Ours - 97.89
-Table 6 Shows that the presented CNN-ViT hybrid model achieved an impressive 97.89% accuracy on the KTH dataset, placing it prominently among state-of-the-art models for HAR.
-This performance is notably superior when compared to earlier benchmarks reported in the literature such as Geng et al.
-[59], who achieved 92.49%, and Arunnehru et al.[73] with 94.90%.
+To complete the comparison, Table 8 shows that the impressive 97.89% accuracy achieved by the presented CNN-ViT hybrid model on the KTH dataset places it prominently among state-of-the-art models for HAR.
+This performance is notably superior when compared to earlier benchmarks reported in the literature such as [35] with 92.49% and [5] with 94.90%.
 Our model utilizes an ImageNet-pre-trained MobileNet [45] as the CNN backbone in the spatial component, which enhances its robust feature extraction capabilities.
-Combined with the dynamic attention mechanisms of ViT, it thereby enhances both the spatial and temporal processing of video sequences.
-Furthermore, the model not only surpasses other contemporary approaches like Liu X et al [83] (91.93%) and Lee et al.
-[91] (89.40%) but also shows competitive or superior performance against some of the highest accuracies in the field, such as Jaouedi et al.
-[87] (96.30%) and Basha et al.
-[96] (96.53%).
-Even in comparison to the high benchmark set by Sahoo et al.
-[85] (97.67%), our model demonstrates a marginal but significant improvement, underscoring the efficacy of integrating CNN with ViT.
+Combined with the dynamic attention mechanisms of ViT, it can thereby enhance both the spatial and temporal processing of video sequences.
+Furthermore, our hybrid model not only surpasses other contemporary approaches like [67] (91.93%) and [62] (89.40%), but also shows competitive/superior performance against some of the highest accuracy in the field, such as [47] (96.30%) and [8] (96.53%).
+Even in comparison to the high benchmark set by [78] (97.67%), our hybrid model demonstrates a marginal but significant improvement, underscoring the efficacy of integrating CNN with ViT.
 This integration not only facilitates more nuanced feature extraction across both spatial and sequential dimensions but also adapts more dynamically to the varied contexts inherent in video data, making it a potent solution for realistic action recognition scenarios.
-The integration of CNN with ViT is particularly advantageous for enhancing feature extraction capabilities and focusing on relevant segments dynamically through the attention mechanisms of ViTs.
+On the whole, the integration of CNN with ViT is particularly advantageous for enhancing feature extraction capabilities and focusing on relevant segments dynamically through the attention mechanisms of ViTs.
 This not only helps in improving accuracy but also in making the model more adaptable to varied video contexts, a key requirement for action recognition in realistic scenarios.
-This comparative advantage suggests that hybrid models are paving the way for future explorations in human action recognition, combining the best of convolutional and ViT-based architectures for improved performance and efficiency.
+This comparative advantage suggests that hybrid models are paving the way for future explorations in HAR, combining the best of convolutional and ViT-based architectures for improved performance and efficiency.
 5 Challenges and Future Directions
-Challenges The field of HAR faces several formidable challenges that stem from the inherent complexity of interpreting human movements within diverse and dynamic environments.
+The field of HAR faces several formidable challenges that stem from the inherent complexity of interpreting human movements within diverse and dynamic environments.
 One of the primary obstacles is the variability in human actions themselves, which can differ significantly in speed, scale, and execution from one individual to another [74].
@@
 Systems must be adept at isolating and recognizing human actions against a backdrop of potentially distracting or obstructive elements, which can vary from the bustling activity of a city street to the unpredictable conditions of outdoor settings [97, 43].
-Moreover, HAR systems must navigate the fine line between inter-class similarity and intra-class variability, where actions that are similar to each other (such as running versus jogging) require nuanced differentiation, while the same action can appear markedly different when performed by different individuals or under varying circumstances [130, 131].
+HAR systems furthermore must navigate the fine line between inter-class similarity and intra-class variability, where actions that are similar to each other (such as running versus jogging) require nuanced differentiation, while the same action can appear markedly different when performed by different individuals or under varying circumstances [38, 113].
 The challenge of temporal segmentation adds another layer of complexity, as accurately determining the start and end of an action within a continuous video stream is crucial for effective recognition [114].
@@
 Addressing these issues is critical for advancing the field and enhancing the practical applicability of HAR systems in real-world applications, from surveillance and security to healthcare and entertainment.
-Future Directions The motivation behind this work has been driven by the compelling need to bridge the existing gap between the spatial feature extraction capabilities inherent in CNNs and the dynamic temporal processing strengths found in ViTs [30].
+The motivation behind this work has been driven by the compelling need to bridge the existing gaps between the spatial feature extraction capabilities inherent in CNNs and the dynamic temporal processing strengths found in ViTs [4].
 Through the introduction of a novel hybrid model, an attempt has been made to leverage the synergistic potential of these technologies, thereby enhancing the accuracy and efficiency of HAR systems in capturing the complex spatial-temporal dynamics of human actions.
@@
 Moreover, addressing the temporal segmentation challenge requires the integration of specialized temporal modules and sequence-to-sequence models to accurately determine the start and end of an action within continuous video streams.
-Computational efficiency remains paramount for real-time processing, necessitating ongoing efforts in model optimization and the exploration of synthetic data generation to mitigate the difficulties associated with obtaining large, accurately annotated datasets.
+Computational efficiency remains paramount for real-time processing, necessitating ongoing efforts in model optimization and the exploration of synthetic data generation to mitigate the difficulties associated with obtaining large and accurately annotated datasets.
 Customizable hybrid CNN-ViT models that can be tailored for specific applications, from surveillance to healthcare, will ensure that these advancements not only push the boundaries of academic research but also enhance practical applicability in real-world scenarios.
@@
 6 Conclusions
-This survey provides a comprehensive overview of the current state of HAR by examining the roles and advancements of RNNs, CNNs, and ViTs.
+This survey provides a comprehensive overview of the current state of HAR by examining the roles and advancements of CNNs, RNNs, and ViTs.
 It delves into the evolution of these architectures, emphasizing their individual contributions to the field.
@@
 This work sets a foundation for future studies to build upon, aiming to push the boundaries of what is currently achievable in HAR and to explore new applications of these technologies in real-world scenarios.
+[…]
+Acknowledgements
+K.A. and H.I.A. are thankful for the support from The Ministry of Education in Saudi Arabia and the Republic of Turkiye Ministry of National Education, respectively.
+[…]
+Author contributions Conceptualisation, K.A.
+and X.C.; methodology, K.A.; software, K.A.; validation,, all authors.; investigation, all authors; resources, K.A.
+and H.I.A.; data curation, K.A.; writing—original draft preparation, all authors; writing—review and editing, all authors; visualisation, K.A.
+and H.I.A.; supervision, X.C.
+All authors have read and agreed to the published version of the manuscript.
+Declarations
+Competing interests The authors declare no competing interests.